### v1

# FIPER: Generalizable Factorized Fields for Joint Image Compression and Super-Resolution

Abstract
This paper introduces a framework that integrates image compression and super-resolution using a unified representation called Generalizable Factor Fields. Traditional approaches to image compression often lead to significant information loss, particularly at low bitrates, which blocks image usability for both human viewing and computer vision applications. Super-resolution techniques strive to reconstruct high-quality images from these degraded compressions but face challenges when starting with heavily compressed sources. Our approach leverages a coefficient-basis structure within the Factor Fields, optimizing the joint task of compressing multiple images and enhancing their resolution in a single cohesive process. By encoding and transmitting images using shared basis functions, our method significantly reduces information loss and boosts efficiency in handling large image datasets. We demonstrate superior performance in both super-resolution and compression tasks compared to existing methods, offering improved image reconstruction quality and more efficient low-bitrate operation.
1 Introduction
Single Image Super-Resolution (SISR) aims to reconstruct high-quality images from low-resolution counterparts. Specifically, the key lies in accurately restoring fine details and reconstructing the correct arrangement of visual features. Thus, geometric correspondences or repetitive patterns, such as stripes, grids, or textures, are commonly used for evaluation due to their rich details that are crucial to image fidelity. Early CNN-based approaches flickr2k, SRCNNlaid the foundation for SISR, which was later enhanced with GAN-based methods srgan for improved perceptual realism. Follow-up Transformer-based networks IPT address non-local dependencies, with subsequent Swin Transformer-based approaches sparking tremendous advancements swin2sr, DAT, RGT, which then inspire more and more complex, delicate designs of heavy network architectures in SISR. However, these prior works have primarily focused on network architecture design rather than addressing the capability of representation. The visual patterns and the inherent nature of image content structure that play an influential role in SISR have not been explicitly considered in the representation learning process. This raises a critical question: beyond a simple network output, can we derive a formulation that more effectively captures these patterns and aligns with the goals of SISR?
On the other hand, image compression serves as a fundamental task in low-level vision applications, where the traditional compression standards VVC-21.2, jpeg, jpeg2000 lay the groundwork. The emerging learned image compression models imc_cnn1, ballé2018variational, guo-hua2023evc, imc_lic_tcm, minnen2018joint, cheng2020learned, compression algorithms of which mostly follow the pixel-space transform coding  imc_cnn1, Goyal2001paradigm, then introduce neural networks to further optimize compression efficiency by learning more compact latent representations and improving reconstruction quality. Specifically, they convert pixels into compact representations through a transform module, which eliminates the redundancy and reduces the bit cost in the subsequent entropy coding process. However, the core challenge of image compression is to accurately reconstruct the information lost during compression and quantization. In other words, the models are to recover the impaired components in images, which can essentially be viewed as reconstructing a high-quality image from its ‘low-resolution’ version, much like Super-Resolution.
Based on the aforementioned analysis, it becomes clear that although Super-Resolution and Image Compression appear to be two distinct tasks, they share mutual similarities in two key aspects: (1) Both tasks require models to restore fine details from low-quality image content, as well as implicitly capture and reconstruct repetitive structural elements. (2) Both aim to conserve image quality, either by enhancing resolution or efficiently compressing data without significant loss of perceptual fidelity.
Hence, inspired by recent advances in decomposition fields and matrices factorization in 3D scene modeling tensorf, mueller2022instantngp, eg3d, kplanes, cao2023hexplane, factor-field, gao2023strivec, we propose a unified representation, Factorized Fields, with generalizable Coefficient Backbone and Basis Transformer. This approach explicitly captures multi-scale visual features and repetitive structural components in images through a basis-coefficient decomposition. The resulting representation strikes a balance between being compact and information-rich, enabling the resolution of structural ambiguities and the precise modeling of image details through a multi-frequency formulation. In the meantime, such a formulation imposes a factorization constraint during model training, which not only enhances the quality of single-image super-resolution (SISR) but also reduces distortion and improves compression efficiency by explicitly modeling structural elements. On top of these, to leverage the robust information-recovering capability of SR models, we integrate such priors with Image Compression models, whose knowledge of detail compensation further refines the lost key elements.
Finally, we propose merging the bases by introducing an additional compression branch, consolidating multiple bases into one alongside multi-image transmission. This approach leverages the mutual information across multiple images, reducing the need for redundant transmissions and refining the basis structure. The main contributions of this paper are summarized as follows:
  • We propose Factorized Fields, a unified representation that explicitly models multi-scale visual features and structural components for both super-resolution and image compression.
  • We integrate super-resolution with image compression by introducing SR prior during decompression to compensate for lost details and developing a merged-basis compression branch for multi-image compression.
  • We demonstrate state-of-the-art performance on benchmarks for both super-resolution and image compression through extensive experiments.
2 Related Works
Super-Resolution (SR).
Image super-resolution is critical in computer vision, focusing on recovering high-resolution (HR) images from low-resolution (LR) inputs. Following the foundational studies, CNN-based strategies SRCNN, flickr2k, zhou2020cross, shufflemixer, VDSR were initially introduced with modeling techniques such as residual learning ledig2017photo, liu2020residual, tong2017image, zhang2018residual, RCAN, drct, EDSR , or recursive learning RGT, tai2017recursive. Besides, subsequent research also sheds light on GAN-base methods srgan, ledig2017photo, esrgan, real_esrgan to enhance realisticity and detail quality. However, the inductive bias of CNN-based networks by restricting spatial locality hinders the capture of long-range dependency from images, which is alleviated by Transformer-based SISR networks IPT, EDT. Afterward, SwinIR swinir is proposed to combine spatial locality and non-local information by Swin Transformer liu2021swin with window attention and achieve breakthrough improvement in SISR. Following SwinIR’s success, several works have built upon its framework swin2sr, zhu2023attention, ATD to reach better image quality as well as solve information bottleneck drct. Hybrid approaches CRAFT (CRAFT) merge the benefits of convolutional and transformer structures to further elevate SR performance. For better feature aggregation, DAT (DAT) and HAT (chen2023hat) integrate spatial and channel information using attention mechanisms to enhance their representation capabilities. Moreover, RGT (RGT) introduces a unique recursive-generalization self-attention mechanism that efficiently captures comprehensive spatial details with a linear increase in computational complexity.
Image Compression (IC).
Deep learning has significantly advanced image compression, offering superior compression ratios and image quality compared to traditional methods like JPEG jpeg and JPEG-2000 jpeg2000. Early CNN-based approaches imc_cnn1, imc_cnn2, imc_cnn3 have been surpassed by transformer-based models imc_trans1, imc_trans2, imc_trans3, duan2023cross-component, which leverage spatio-channel attention for better performance. GAN-based methods imc_gan1, imc_gan2, imc_gan3, imc_gan4 have further contributed to real-time adaptive compression. Recently, ELIC imc_elic introduced efficient compression with unevenly grouped space-channel contextual adaptive coding, while LIC TCM imc_lic_tcm integrated transformers and CNNs to capture both global and local image features. The eContextformer koyuncu2024efficient introduced patch-wise, checkered, and channel-wise grouping techniques for parallel context modeling with a shifted window spatial-channel attention mechanism. GroupedMixer li2024groupedmixer proposed a transformer-based entropy model with token mixers for inner and cross-group context modeling. Meanwhile, Wavelet Conditional Diffusion song2024wavelet-compression introduced a wavelet-based model with uncertainty-aware loss, balancing high perceptual quality with low distortion.
In summary, while recent super-resolution and image compression advances focus on increasingly complex architectures, our work takes a different approach. We propose Factorized Fields, a unified framework that models visual and structural features, offering a more comprehensive solution to enhance performance in both tasks.
3 Methods
In this section, we first briefly introduce the background of the factor fields factor-field, their properties, and also the preliminary of Learned Image Compression in Sec. 3.1. Inspired by factor fields, we next explain the motivation and how we derive our Factorized Fields for enhanced image reconstruction quality in Sec. 3.2. We then describe how to adapt the formulation to Super-Resolution in Sec. 3.3. Finally, we show how to incorporate such representation with image(s) compression and how we integrate Super-Resolution and Image Compression in Sec. 3.4.
Figure 1: The correlation between coordinate transformation and downsampling. (a) The sawtooth transformation example with $k=2$. (b) The PixelUnShuffle downsample. (c) To explicitly model the information for sampling with a sawtooth, we rearrange the feature map in a dilation-like manner in the downsample layer of the Basis Swin Transformer. This way, the feature sampled would capture the information in the original layout correctly.
3.1 Preliminaries
Factor Fields.
The concept of decomposition fields or factorized matrices in the reconstruction of 2D images or 3D scenes has shown superior rendering quality and improved efficiency recently tensorf, gao2023strivec, mueller2022instantngp, kplanes, cao2023hexplane, eg3d, while factor-field first proposed a unified representation called factor fields. For a 1D signal $s(x)$, the factor fields formulation is:
$\hat{s}(x)=c(x)^{T}b(\gamma(x)),$ (1)
where $c(x)=(c_{1}(x),...,c_{N}(x))^{T}$ are spatial-varying coefficient fields, $b(x)=(b_{1}(x),...,b_{N}(x))^{T}$ are basis functions, and $\gamma(x)$ is a coordinate transformation. The coefficient and basis are queried by sampling with the coordinates $x$. This generalizes to Q-dimensional signals:
$\hat{s}(x)=\mathcal{P}(c(x)\odot b(\gamma(x))),$ (2)
where $\mathcal{P}$ is a projection function and $\odot$ denotes the element-wise product. In practice, for multi-dimensional coefficients:
$\hat{s}(x)=\mathcal{P}\Bigl{(}\textbf{Concat}_{i=1}^{N}\Bigl{\{}c_{i}(x)\odot b_{i}(\gamma_{i}(x))\Bigl{\}}\Bigl{)}.\end{aligned}$
This formulation allows the same basis to be applied at multiple spatial locations while varying the coefficients, especially when $\gamma$ is a periodic function.
Learned Image Compression.
Following (channel-autoregressive, ballé2018variational), a learned image compression model with a channel-wise entropy model can be formulated as:
$z$ $=h_{a}(y;\phi_{h}),\ y=g_{a}(x;\phi),$ (4)
$\{X_{\text{mean}},X_{\text{scale}}\}$ $=h_{s}(\hat{z};\theta_{h}),\ \hat{z}=Q(z)$
$\hat{y}$ =$=\{Q(y_{0}-\mu_{0})+\mu_{0},...,Q(y_{t}-\mu_{t})+\mu_{t}\},\ 0<=t<l,\ \mu_{t}=e_{i}(\hat{y}_{<i},X_{\text{mean}}))$
$\hat{x}$ $=g_{s}(\overline{y};\theta),\ \overline{y}=\textbf{Refine}_{\theta_{r}}({\mu_{0},...,\mu_{t}},\hat{y}).$
The encoder $g_{a}$ transforms the raw image $x$ into a latent representation $y$. A hyper-prior encoder $h_{a}$ further processes $y$ to output $z$, capturing spatial dependencies. $z$ is quantized to $\hat{z}$, which is decoded by $h_{s}$ to produce features $X_{\text{mean}}$ and $X_{\text{scale}}$, used to estimate the mean $\mu$ and variance $\sigma$ of $y$. The latent $y$ is divided into $l$ slices, and each quantized around computed means $\mu_{t}$. These means are derived from earlier quantized slices and $X_{\text{mean}}$ by a slice network $e_{i}$. The quantized slices form $\hat{y}$. For decompression, $\hat{y}$ is refined using $\textbf{Refine}_{\theta_{r}}$ based on $\mu_{t}$ and $\hat{y}$ to produce $\overline{y}$, approximating the original $y$. Finally, $g_{s}$ reconstructs the decompressed image $\hat{x}$ from $\overline{y}$. The model is trained using a Lagrangian multiplier-based rate-distortion optimization:
$L=R(\hat{y})+R(\hat{z})+\lambda\cdot D(x,\hat{x}),$ (5)
where $R(\hat{y})$ and $R(\hat{z})$ denote bit rates, $D(x,\hat{x})$ is the distortion term (calculated by MSE), and $\lambda$ balances compression efficiency and image fidelity. In our experiments, we follow imc_lic_tcm, modifying only $g_{s}$ to demonstrate our representation’s effectiveness.
Figure 2: The overall pipeline of image super-resolution with our Factorized Fields. Given a low-resolution image $I_{LR}$, we first extract coefficient feature map $X_{\text{coeff}}$ with the coefficient backbone, which is then decoded into coefficient and passed through the basis Swin Transformer for basis, separately. Finally, as in Eq. 8, the coefficient and basis are sampled, multiplied, and decoded for final high-resolution output $I_{HR}$, where $s$, $H$, $W$ denote the scale factor, height, and width respectively.
3.2 Formulation of Factorized Fields
As discussed in Sec. 1, the key to superior rendering quality in image regression tasks of both Image Compression and Super-Resolution lies in the capability of the representation to capture accurate structural distribution and fine visual details. Meanwhile, transformations such as Fourier Transform or Wavelet Transform have long been used to model multi-frequency information fuoli2021fourier, waveletsr to express different implicit functions in images; however, such methods often suffer from under-expression due to the pre-defined and limited frequency bands or restriction from its formulation. Thus, we seek a representation that explicitly incorporates and fits multi-scale and multi-frequency components and yet is highly flexible and learnable according to individual features in an image.
Inspired from recent success on such decomposition fields kplanes, factor-field, tensorf, mueller2022instantngp, cao2023hexplane, eg3d, we primarily built our Factorized Fields framework on factor fields factor-field from Eq. 3:
$\hat{I}(x)=\mathcal{P}\Bigl{(}\textbf{Concat}_{i=1}^{N}\Bigl{\{}c_{i}(x)\odot b_{i}(\gamma_{i}(x))\Bigl{\}}\Bigl{)}.\end{aligned}$
Note that $\hat{I}$ denotes the approximated images, and $x\in R^{2}$ are the pixel coordinates.
Such formulation has several key properties: First of all, by decomposing the images into basis frequencies, we can learn the implicit functions of an image and capture the mutual dependencies between pixels and across spatial composition; meanwhile, since the basis and coefficient are specific to every single image and both learnable in all spatial dimensions, the restriction on a limited number of basis (we use $N=6$ in all of our experiments) can be alleviated. Finally, as in factor-field, of all the tested coordinate transformation $\gamma$ in Eq. 3 sawtooth $\gamma(x)=x\bmod{k},k\in\mathbb{R}$ performs mostly the best in image regression tasks. We can easily observe that such transformation implicitly captures patch-like frequency information as shown in Fig. 1a, and thus, we propose that by leveraging the inter-patch information from the sawtooth coordinate transformation, the visual correspondence between spatial locations can be effectively represented.
However, in practice, we sample the basis via bilinear or bicubic sampling due to memory constraints, i.e., the feature size is less than image height and width, and this poses a severe problem: These nearby sampled pixel features are actually linear or based on cubic interpolation with respect to the basis, and such inductive bias hinders the representation of non-linearity in images. To resolve this, we propose a modified version, our Factorized Fields:
$\hat{I}(x)=\mathcal{P}\Bigl{(}{\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl{\{}c_{i}(x)\odot\psi(\alpha_{j}\cdot b_{i}(\gamma_{i}(x)))\Bigl{\}}\Bigl{)},\psi\in\{\text{sin},\text{cos}\},\alpha_{j}\in R.$ (7)
Here, inspired by the Fourier Series, with a scalar $\alpha$ which can be viewed as increasing to sampling frequency and a transformation function $\psi$ to add to implicit non-linearity functions and modulation of product, our complete Factorized Fields method stands out by the fact that such formulation significantly enhances the ability to represent complex non-linear structures in images and effectively composes high-frequency components between pixels.
Figure 3: The illustration of our joint image-compression and super-resolution framework compared with the traditional compression-only method. (a) Traditional learning-based compression methods. (b) Our approach surpasses (a) by incorporating our Super-Resolution (SR) Module from Sec.3.3 as information-recovery prior, detailed in Sec. 3.4.1. (c) Expanding on (b), Sec. 3.4.2 introduces a multi-image compression strategy that utilizes both our SR Module and a Basis Merging Transformer to capture shared structure.
3.3 Super-Resolution with Factorized Fields
We represent a super-resolved image using our Factorized Fields, where coefficients and basis are generated by networks $F_{\text{coeff}}$ and $F_{\text{basis}}$ from a low-resolution Image:
$\hat{I}_{\text{SR}}(x)=\mathcal{P}\Bigl{(}{\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl{\{}c_{i}^{\text{LR}}(x)\odot\psi(\alpha_{j}\cdot b_{i}^{\text{LR}}(\gamma_{i}(x)))\Bigl{\}}\Bigl{)},$ (8)
where $c^{\text{LR}}_{i}(x)=\text{Conv}(F_{\textbf{coeff}}(I_{\text{LR}}))_{i}(x)$ and $b^{\text{LR}}_{i}(x)=F_{\textbf{basis}}(F_{\textbf{coeff}}(I_{\text{LR}}))_{i}(\gamma_{i}(x))$. Note that we sample the outputs $\text{Conv}(F_{\textbf{coeff}}(I_{\text{LR}}))$ and $F_{\textbf{basis}}(F_{\textbf{coeff}}(I_{\text{LR}}))$ with coordinates $x$ and $\gamma(x)$, respectively.
Our model comprises three main components: Coefficient Backbone, Basis Swin Transformer, and Factorized Fields Reconstruction. As shown in Fig. 2, the process begins with $I_{LR}\in\mathbb{R}^{3\times H\times W}$. The Coefficient Backbone extracts features $X_{\text{coeff}}\in\mathbb{R}^{C_{c}\times H_{c}\times W_{c}}$, which are then used to generate coefficients $c$ through convolution and pixel shuffle operations, and fed into the Basis Swin Transformer to produce a multi-scale basis $b=\{b_{1},...,b_{N}\},\ b_{i}\in\mathbb{R}^{C_{b_{i}}\times H_{b_{i}}\times W_{b_{i}}}$. The coefficients and basis are combined to reconstruct $I_{SR}\in\mathbb{R}^{3\times sH\times sW}$ using Eq. 8, where $s$ is the scale factor. In our experiments, $\gamma_{i}(x)$ is a Sawtooth function: $\gamma_{i}(x)=x\mod k_{i}$, with $k_{i}\in R$. We optimize model parameters using an $L_{1}$ loss function. To demonstrate our method’s effectiveness, we use existing SR methods chen2023activating, DAT, ATD as the Coefficient Backbone. For the Basis Swin Transformer, we employ Swin Transformer Blocks liu2022swinv2 with a series of downsampling operations. We use a dilation-like downsampling technique (Fig. 1(c)) to accommodate the sawtooth sampling pattern. The final basis is refined using additional upsampling and convolution layers.
3.4 Image Compression with Factorized Fields and Super-Resolution
Image Compression, at its core, is to strike a balance between the amount of information contained in the latent bits and the final image quality. However, as discussed in Sec. 1,  2, and 3.2, most recent works draw emphasis on how to retrieve implicit elements through designing different architectures, such as analysis transforms and entropy models koyuncu2024efficient, li2024groupedmixer, or decompression modules song2024wavelet-compression, duan2023cross-component, while this paper aims at addressing the representation itself to better capture the structural correlations and thus achieve better image quality through explicit modeling of frequency components and Factorized Fields formulation as in Eq. 7. In addition, with our trained SR model described in Sec. 3.3, it intuitively serves as a strong prior for information recovery, i.e., it contains extensive knowledge of how to reconstruct missing details and enhance image quality by leveraging learned patterns from the training data. Thus, since Super-Resolution and Image Compression share the core principle of reconstructing and enhancing image details from low-quality sources, we can effectively integrate this prior into the compression pipeline. In the following, we respectively present the proposed joint framework for the single and multi-image compression tasks.
3.4.1 Single Image Compression
The overall pipeline is shown in Fig. 3(b). To demonstrate the robustness of our representation and the effectiveness of the SR prior, the compression and decompression networks greatly follow imc_lic_tcm, with only the synthesis transform replaced by our SR pipeline, where the details can be referenced in Supplementary Materials. In practice, the training is performed in two stages. After we obtain the trained SR prior, the model is fine-tuned with a lower learning rate alongside the compression module, which is then trained end-to-end with the loss function defined in Eq. 5.
3.4.2 Multi-Image Compression
For each basis $b_{i}\in R^{C_{b_{i}}\times H_{b_{i}}\times W_{b_{i}}}$ associated with any arbitrary image, we can consider it as encapsulating the inherent pixel structure. These bases can be combined into a unified generic basis that captures the structural distribution of images and potentially reduces noise. Given $M$ images and their respective bases $b^{n}_{i},n\in\{1,...,M\}$, we apply a Basis Merging Transformer $F_{\text{merge}}$ at each location to integrate the $M$ elements:
$b_{i}(h,w)=F_{\text{merge}}\left(\{b^{n}_{i}(h,w)\ |\ n\in{1,...,M}\}\right),\ 0\leq h<H_{b_{i}},\ 0\leq w<W_{b_{i}}.$ (9)
Here, $F_{\text{merge}}$ is a standard transformer, similar in architecture to that described in (dinov2). We treat the bases as tokens and prepend the sequence with a $CLS$ token, concluding the sequence to form the final merged basis.
Since compression induces error, according to Fig. 3b, the Coefficient feature map $X_{\text{coeff}}$ generated with Coefficient Backbone contains misinformation, and Basis Swin Transformer that uses such features would capture the information wrongfully for the basis and thus amplify error. To solve this, we exploit the mergeable property of basis, compress and transmit the merged basis independently along other quantized variables $\{Q(y,\mu),\ \hat{z}\}$, and finally reconstruct the $M$ images with the merged basis and their individual decoded coefficients. as illustrated in Fig. 3c. This way, we can enjoy less information loss with basis while maintaining low bit rates.
Table 1: Quantitative comparisons on 4×4$\times$4 × super-resolution with state-of-the-art methods. The best results are colored red. The models with †are those who use same-task pretraining chen2023activating. Please refer to quantitative results in Sec. 4.1 for details.
Params MACs Forward Pass Set5 Set14 B100 Urban100 Manga109
[…]
SwinIR (swinir)(ICCV2021W) 28.01 119.68 3,826 32.93 0.9043 29.15 0.7958 27.95 0.7484 27.56 0.8273 32.22 0.9273
ATD (zhang2024transcending)(CVPR2024) 20.26 77.10 6,572 33.14 0.9061 29.25 0.7976 28.02 0.7524 28.22 0.8414 32.65 0.9308
DAT (DAT))(ICCV2023) 14.80 61.66 4,192 33.15 0.9062 29.29 0.7983 28.03 0.7518 27.99 0.8365 32.67 0.9301
RGT (RGT)(ICLR2024) 13.37 834.25 3,404 33.16 0.9066 29.28 0.7979 28.03 0.7520 28.09 0.8388 32.68 0.9303
HAT (chen2023hat)(CVPR2023) 20.77 86.02 3,692 33.18 0.9073 29.38 0.8001 28.05 0.7534 28.37 0.8447 32.87 0.9319
HAT-L (chen2023hat)(CVPR2023) 40.84 167.27 6,804 33.30 0.9083 29.47 0.8015 28.09 0.7551 28.60 0.8498 33.09 0.9335
ATD-L* 49.42 184.83 15,582 33.12 0.9062 29.31 0.7985 28.02 0.7514 28.25 0.8422 32.78 0.9309
DAT-L* 43.01 175.42 11,326 33.33 0.9084 29.40 0.8009 28.04 0.7543 28.49 0.8473 33.02 0.9321
ATD-F (Ours) 45.46 149.87 8,674 33.29 0.9082 29.48 0.8017 28.03 0.7539 28.53 0.8487 33.11 0.9335
DAT-F (Ours) 40.00 134.42 6,206 33.45 0.9094 29.60 0.8039 28.13 0.7560 28.75 0.8520 33.23 0.9339
HAT-F (Ours) 45.97 158.79 5,750 33.53 0.9100 29.65 0.8050 28.18 0.7569 28.79 0.8527 33.33 0.9342
HAT-F-ImageNet (Ours) 45.97 158.79 5,750 33.55 0.9102 29.63 0.8049 28.18 0.7569 28.80 0.8529 33.33 0.9342
HAT-L-F (Ours) 66.04 240.03 8,888 33.75 0.9116 29.87 0.8091 28.31 0.7597 29.51 0.8637 33.36 0.9343
HAT-F-Basis-First (Ours) 46.67 161.66 5,696 33.33 0.9085 29.47 0.8015 28.10 0.7554 28.57 0.8494 33.14 0.9336
HAT-F-Concat (Ours) 45.52 129.05 4,826 33.46 0.9095 29.57 0.8035 28.16 0.7566 28.73 0.8518 33.28 0.9341
[…]
4 Experiments
4.1 Image Super-Resolution
Experimental Setup.
We conduct extensive experiments to validate the effectiveness of our Factorized Fields representation for Super-Resolution tasks. Following the strategy outlined in (chen2023hat) and (drct), we adapt the same-task pretraining approach for all the Super-Resolution models. Unlike these previous works, we leverage the SA-1B dataset from (sam), which includes approximately 10 million images, where we randomly sample 4 million for training, which is much less that of ImageNet (imagenet), which is used for chen2023hat and drct pretraining. Note that we use SA-1B just for its content-rich and high-resolution images since SA-1B is designed for training segmentation models, whereas ImageNet focuses primarily on single-class prediction. We further conduct experiments to compare training performance with SA-1B and ImageNet in Tab. 1. While the performances are mostly equivalent, we witness faster convergence with SA-1B during training.
To be more focused on the representation itself, we utilize various pre-trained SR models (SwinIR (swinir), HAT (chen2023activating), DAT (DAT), and ATD (zhang2024transcending)) as the Coefficient Backbone, with a consistent Basis Transformer architecture. Training involves initializing the Coefficient Backbones from pre-trained SR models and randomly initializing the Basis Transformers. Models are pre-trained for 300k iterations on the SA-1B dataset. After pretraining, we use DF2K (DIV2K (div2k) + Flickr2K (flickr2k)) as the finetuning dataset following (swin2sr, chen2023hat) for 200k iterations.
For pretraining, we utilize AdamW optimizer with learning rate 1e-4, batch size 16, betas (0.9, 0.99), and other parameters set to PyTorch default, while we use learning rate 1e-5 during finetuning. Throughout training, the input is randomly cropped to $256\times 256$ and bicubicly resized to $64\times 64$ for Coefficient Backbone input. As for the hyperparameter of Factorized Fields, the number $N$ of coefficient and basis is set to 6, and the scalars $\alpha_{j}$ are set to {1, 4, 16, 64} since we want to capture both base frequency $\alpha_{j}=1$ and high frequency $\alpha_{j}=64$ information.
Quantitative Results.
Tab. 1 presents the quantitative comparison between our approach and state-of-the-art (SoTA) methods. We evaluate the methods using five benchmark datasets, including Set5 (set5), Set14 (set14), BSD100 (bsd100), Urban100 (urban100), and Manga109 (manga109). For quantitative metrics, PSNR and SSIM are reported. The average relative improvement of 204.4% in PSNR over the baseline across the five datasets is calculated using the formula $(c-b)/(a-b)$, where $b$ represents the PSNR of a representative baseline, SwinIR, $a$ represents the PSNR of SOTA, HAT-L, and $c$ represents the PSNR of our best model, HAT-L-F. This formula measures the relative performance gain of our model compared to the gap between HAT-L and SwinIR.
To validate the effectiveness of our framework, we employ three different SoTA SR models—ATD ATD, DAT DAT, and HAT chen2023activating—as the Coefficient Backbone in our pipeline, which we denote as ATD-F, DAT-F, and HAT-F, respectively. These models exhibit significant improvements when compared to their counterparts ATD-L, DAT-L, and HAT-L, which possess similar parameter counts. It is important to note that only chen2023activating provides a large-scale model. To maintain fairness, we scale up ATD and DAT to match this size and train them under the same configuration as our model, including both pretraining and fine-tuning stages. To ensure an equitable comparison with other methods, we further train another model, HAT-F-ImageNet, using ImageNet as the pretraining dataset, following the protocols outlined in IPT, chen2023activating. The results demonstrate that its performance remains consistent with only minor perturbations.
Furthermore, in traditional Fourier Series and other image processing methods jpeg, the basis is typically derived first and then used to compute the coefficients. In contrast, our method derives the coefficient features first, as illustrated in Fig. 2. To explore this difference, we develop another variant of our model, denoted HAT-F-Basis-First, where we reverse the order of operations. In this case, we first pass the image through the Basis Swin Transformer and then use the resulting basis features and the image input to derive the coefficients. This approach, however, leads to a gigantic performance drop, showing the importance of the order of the pipeline. Specifically, we argue that in our pipeline, the Coefficient Backbone functions more as a feature extraction module, where the refined features facilitate downstream basis extraction.
Lastly, to evaluate the effectiveness of our Factorized Fields, we trained a model named HAT-F-Concat, which does not apply the formulation in Eq. 7. Instead, it concatenates the basis and coefficient directly and decodes the resulting features to produce the output. Although this approach results in reduced performance, which indicates the representation does act as an imperative role in modeling image information, the Basis Swin Transformer with Sawtooth downsampling still contributes to improved reconstruction, even without Factorized-Fields decoding, highlighting its effectiveness.
Visual Comparison.
We provide the visual comparison in Fig. 4. The images are randomly sampled from the DIV2K dataset. Our method faithfully reconstructs the image details, whereas the other approaches suffer from over-smoothing or hallucinating details absent in the ground truth.
Table 2: Comprehensive evaluation for image compression. Using VTM as an anchor for calculating BD-Rate. Latencies are measured under an NVIDIA GTX 3090 GPU.
Method BD-Rate (%) $\downarrow$ Latency(s) Params(M)
Tot Enc $\downarrow$ Tot Dec $\downarrow$
VTM 0.00 129.21 0.14 -
Xie (MM 21’) -0.78 2.93 6.00 50.0
Cheng (CVPR 22’) 5.44 1.98 4.69 29.6
STF (CVPR 22’) -4.31 0.14 0.13 99.9
ELIC (CVPR 22’) -7.24 0.07 0.09 36.9
TCM (CVPR 23’) -11.74 0.16 0.15 76.7
TCM-HAT-L-F (Ours) -21.09 0.109 0.264 110.34
TCM-HAT-F-multi M=1 (Ours) 27.96 0.232 0.174 131.35
TCM-HAT-F-multi M=2 (Ours) 2.70 0.232 0.174 131.35
TCM-HAT-F-multi M=4 (Ours) -10.11 0.232 0.174 131.35
TCM-HAT-F-multi M=8 (Ours) -16.61 0.232 0.174 131.35
TCM-HAT-F-multi M=16 (Ours) -19.88 0.232 0.174 131.35
TCM-HAT-F-multi M=24 (Ours) -20.97 0.232 0.174 131.35
{subfigure}
[b]0.32 Refer to caption {subfigure}[b]0.32 Refer to caption {subfigure}[b]0.32 Refer to caption
Figure 5: Kodak
Figure 6: CLIC
Figure 7: Tecnick
Figure 8: Performance (RD-Curve) evaluation on image compression using different datasets.
4.2 Single- and Multi-Image Compressions
Experimental Setup.
We evaluate our Factorized Fields representation for image compression tasks, comparing it against state-of-the-art methods. Following our Super-Resolution setting in Experiment Setup from Sec. 4.1, we use the same set of SA-1B for training. To emphasize our representation, we initialize compression and decompression modules in Fig.3 from pre-trained imc_lic_tcm and the SR Module from those in Sec. 4.1, we then train the pipeline end to end on 256x256 patches for 200k iterations, with AdamW optimizer (loshchilov2017decoupled) with learning rate 1e-5, batch size 16, betas (0.9, 0.99) and other parameters set to PyTorch default.
We integrate our SR Module with the pre-trained compression module TCM, creating TCM-HAT-F and TCM-HAT-L-F models. TCM-HAT-F-multi represents the multi-image compression pipeline. For multi-image compression in Fig. 3(c), we set the Basis Swin Transformer and the Basis Merging Transformer to be trainable while the other parts remain frozen.
Rate-Distortion Performance Comparison.
We compare our model with State-of-the-Art learned end-to-end image compression algorithms, including (liu2023learned), (Chen2022), zou2022the, xie2021enhanced, cheng2020learned, ballé2018variational, li2024groupedmixer, jiang2023mlic, channel-autoregressive, bellard_bpg, qi2023hyperthumbnail, and imc_elic. The classical image compression codec, VVC (vtm), is also tested by using VTM12.1.The rate-distortion performance on various datasets, including Kodak, Tecnick’s old test set with resolution 1200$\times$1200, and CLIC Professional Validation, is shown in Fig. 8.
In Tab. 2, our TCM-HAT-L-F model achieved a significant BD-Rate improvement of -21.09% compared to VTM, outperforming previous state-of-the-art methods. The multi-image compression approach (TCM-HAT-F-multi) shows increasing performance gains with the number of images compressed simultaneously, reaching -20.97% BD-Rate improvement for $M=24$. The result shows that direct transmission of bases would reduce the error from Coefficient Backbone to Basis Swin Transformer and that the distortion increases with $M$, as the information contained in the merged basis is limited, and merging multiple bases into one would cause increasing information loss.
Our analysis reveals several significant advantages of the FIPER framework in image compression tasks. The approach demonstrates substantial improvements in compression efficiency across various benchmark datasets, including Kodak, CLIC, and Tecnick, indicating its broad applicability. Notably, the multi-image compression strategy shows particularly promising results for larger image sets, suggesting scalability benefits. Furthermore, our method maintains competitive latency while significantly improving compression performance and balancing efficiency and quality.
Table 3: Comparison of improvements of Factorized Fields. $\psi$ and $\alpha$ are the same in Eq. 8
Metric PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Baseline (factor-field) 22.04 0.505 0.5296
Ours 38.44 0.999 0.0385
No $\psi$ to control magnitude 13.46 0.147 0.766
No $\alpha$ for pixel-wise frequency information 21.25 0.537 0.527
4.3 Ablation Studies
Effectiveness of Factorized Fields Design.
We conduct experiments to verify our modification of Factorized Fields, modified from Eq. 3 to Eq. 7. The quantitative performance reported on single-image regression is shown in Tab. 3, where each result is measured after 256 iterations. Compared to baseline results, our refinements in modeling pixel-level frequency have significantly improved all performance metrics. Additionally, our results demonstrate that the modulation function $\psi$ and the scalar $\alpha$ are interdependent, each essential to the other’s function.
Influence of SR Priors in Image Compression.
We conduct experiments with various configurations to verify the proposed image compression pipeline’s effectiveness. As shown in Tab. 4, we present the quantitative performance of models trained with different values of $\lambda$ in Eq. 5. Specifically, TCM-HAT refers to substituting our SR Module with the original HAT chen2023activating in the pipeline illustrated in Fig. 3.b. TCM-HAT-F represents our complete pipeline, while TCM-HAT-F-Scratch denotes the same pipeline but with the SR Module initialized randomly. Our results demonstrate that integrating SR priors with image compression improves performance, and our proposed representation further enhances results. This highlights the robustness of our Factorized Fields in capturing fine details in image regression tasks.
Table 4: Validation of the effectiveness of SR prior. The best PSNRs are marked in red.
Method Kodak CLIC Tecnick
$\lambda=0.0025$ $\lambda=0.0067$ $\lambda=0.025$ $\lambda=0.0025$ $\lambda=0.0067$ $\lambda=0.025$ $\lambda=0.0025$ $\lambda=0.0067$ $\lambda=0.025$
[…]
TCMimc_lic_tcm 0.1533 30.0834 0.2983 32.5841 0.6253 36.1345 0.1214 31.8207 0.2235 34.2098 0.4503 37.1201 0.1268 32.0588 0.2193 34.3669 0.3981 36.9066
TCM-HAT-F-Scratch 0.1570 30.0857 0.2976 32.5893 0.6211 36.1389 0.1214 31.9421 0.2235 34.2894 0.4503 37.1434 0.1258 32.0632 0.2189 34.3781 0.4001 36.9223
TCM-HAT 0.1567 30.1843 0.2992 32.6454 0.6268 36.2267 0.1220 31.9737 0.2266 34.3319 0.4512 37.2486 0.1262 32.1423 0.2174 34.5124 0.3971 36.9934
TCM-HAT-F 0.1574 30.4012 0.2998 32.8910 0.6276 36.4461 0.1229 32.1917 0.2249 34.4109 0.4512 37.3135 0.1255 32.4591 0.2186 34.7656 0.3975 37.3244
5 Conclusion
We proposed Factorized Fields, a representation that models implicit structures and patterns by decomposing images into multi-frequency components. This approach addresses challenges in Super-Resolution and Image Compression by restoring details and preserving visual fidelity. We integrate SR priors with Image Compression for improved information recovery and introduce a basis merging technique for enhanced rendering quality across multiple images. Experiments demonstrate state-of-the-art performance in both SISR and Image Compression benchmarks, addressing limitations of previous methods.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
[…]
    means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
[…]
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS paper checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
[…]
  4. Justification: Our abstract and introduction accurately reflect the paper’s contributions and scope by highlighting our novel contribution: we employ the factor fields for enhance the performance in super-resolution and multi-image compression.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
[…]
  9. Justification: Our paper explicitly discusses the limitations of our approach, including the long training and inference time.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
[…]
  14. Justification: Our paper primarily focuses on empirical research and does not introduce new theoretical results. As such, there are no theoretical assumptions or proofs discussed in the study.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
[…]
  19. Justification: Our paper thoroughly details the methodologies, parameter settings, and experimental setups, providing all necessary information for replicating the main experimental results. This comprehensive documentation supports the validity of our primary claims and conclusions, ensuring clarity and reproducibility.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
[…]
  24. Justification: We will make a full open access to both the code and the data used in our experiments.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
[…]
  29. Justification: Our paper meticulously details the experimental setup, including data splits for training and testing, hyperparameter configurations, and the types of optimizers utilized. These elements are crucial for a thorough understanding of the experimental procedures and the interpretation of results.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
[…]
  34. Justification: Due to the limited scope of our experiments, which encompass only two trials yielding similar results, our paper does not include detailed error bars or extensive statistical analysis. We recognize the potential for variability in these findings and suggest that additional experiments be conducted under varied conditions to validate and expand upon our results.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
[…]
  39. Justification: We use GPU GTX 4090 as our computational resources.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
[…]
  44. Justification: We carefully follow the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
[…]
  49. Justification: In the conclusion of our paper, we discuss the societal impacts directly related to our research on super resolution and image compression.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
[…]
  54. Justification: Our paper does not include the release of pre-trained models, new datasets, or any other resources that typically necessitate specific safeguards to prevent misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
[…]
  59. Justification: We have diligently credited the creators and original owners of all assets used in our research, including code, data, and models, ensuring that each source is accurately cited within our paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
[…]
  64. Justification:: Yes, any new assets introduced in the paper will be well documented.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
[…]
  69. Justification: Our research does not involve crowdsourcing experiments.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
[…]
  74. Justification: Our research does not involve human subjects, and therefore, there were no study participants exposed to any risks.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v2

@@
 Abstract
-This paper introduces a framework that integrates image compression and super-resolution using a unified representation called Generalizable Factor Fields.
-Traditional approaches to image compression often lead to significant information loss, particularly at low bitrates, which blocks image usability for both human viewing and computer vision applications.
-Super-resolution techniques strive to reconstruct high-quality images from these degraded compressions but face challenges when starting with heavily compressed sources.
-Our approach leverages a coefficient-basis structure within the Factor Fields, optimizing the joint task of compressing multiple images and enhancing their resolution in a single cohesive process.
-By encoding and transmitting images using shared basis functions, our method significantly reduces information loss and boosts efficiency in handling large image datasets.
-We demonstrate superior performance in both super-resolution and compression tasks compared to existing methods, offering improved image reconstruction quality and more efficient low-bitrate operation.
+In this work, we propose a unified representation for Super-Resolution (SR) and Image Compression, termed Factorized Fields, motivated by the shared principles between these two tasks.
+Both SISR and Image Compression require recovering and preserving fine image details—whether by enhancing resolution or reconstructing compressed data.
+Unlike previous methods that mainly focus on network architecture, our proposed approach utilizes a basis-coefficient decomposition to explicitly capture multi-scale visual features and structural components in images, addressing the core challenges of both tasks.
+We first derive our SR model, which includes a Coefficient Backbone and Basis Swin Transformer for generalizable Factorized Fields.
+Then, to further unify these two tasks, we leverage the strong information-recovery capabilities of the trained SR modules as priors in the compression pipeline, improving both compression efficiency and detail reconstruction.
+Additionally, we introduce a merged-basis compression branch that consolidates shared structures, further optimizing the compression process.
+Extensive experiments show that our unified representation delivers state-of-the-art performance, achieving an average relative improvement of 204.4% in PSNR over the baseline in Super-Resolution (SR) and 9.35% BD-rate reduction in Image Compression compared to the previous SOTA.
 1 Introduction
+Figure 1: The Proposed Factorized Fields.
+With explicit formulation of repetitive patterns and high frequency components, images can be represented with fine-grained details, especially useful for both Super-Resolution and Image Compression.
+Note that in this figure $\alpha=\{1,16\}$ and $\psi=sin$.
 Single Image Super-Resolution (SISR) aims to reconstruct high-quality images from low-resolution counterparts.
@@
 Thus, geometric correspondences or repetitive patterns, such as stripes, grids, or textures, are commonly used for evaluation due to their rich details that are crucial to image fidelity.
-Early CNN-based approaches flickr2k, SRCNNlaid the foundation for SISR, which was later enhanced with GAN-based methods srgan for improved perceptual realism.
-Follow-up Transformer-based networks IPT address non-local dependencies, with subsequent Swin Transformer-based approaches sparking tremendous advancements swin2sr, DAT, RGT, which then inspire more and more complex, delicate designs of heavy network architectures in SISR.
+Early CNN-based approaches [46, 21]laid the foundation for SISR, which was later enhanced with GAN-based methods [39] for improved perceptual realism.
+Follow-up Transformer-based networks [12] address non-local dependencies, with subsequent Swin Transformer-based approaches sparking tremendous advancements [19, 15, 16], which then inspire more and more complex, delicate designs of heavy network architectures in SISR.
 However, these prior works have primarily focused on network architecture design rather than addressing the capability of representation.
@@
 This raises a critical question: beyond a simple network output, can we derive a formulation that more effectively captures these patterns and aligns with the goals of SISR?
-On the other hand, image compression serves as a fundamental task in low-level vision applications, where the traditional compression standards VVC-21.2, jpeg, jpeg2000 lay the groundwork.
-The emerging learned image compression models imc_cnn1, ballé2018variational, guo-hua2023evc, imc_lic_tcm, minnen2018joint, cheng2020learned, compression algorithms of which mostly follow the pixel-space transform coding  imc_cnn1, Goyal2001paradigm, then introduce neural networks to further optimize compression efficiency by learning more compact latent representations and improving reconstruction quality.
+On the other hand, image compression serves as a fundamental task in low-level vision applications, where the traditional compression standards [33, 69, 66] lay the groundwork.
+The emerging learned image compression models [7, 2, 29, 48, 58, 17], compression algorithms of which mostly follow the pixel-space transform coding  [7, 27]paradigm, then introduce neural networks to further optimize compression efficiency by learning more compact latent representations and improving reconstruction quality.
 Specifically, they convert pixels into compact representations through a transform module, which eliminates the redundancy and reduces the bit cost in the subsequent entropy coding process.
@@
 (2) Both aim to conserve image quality, either by enhancing resolution or efficiently compressing data without significant loss of perceptual fidelity.
-Hence, inspired by recent advances in decomposition fields and matrices factorization in 3D scene modeling tensorf, mueller2022instantngp, eg3d, kplanes, cao2023hexplane, factor-field, gao2023strivec, we propose a unified representation, Factorized Fields, with generalizable Coefficient Backbone and Basis Transformer.
+Hence, inspired by recent advances in decomposition fields and matrices factorization in 3D scene modeling [9, 59, 8, 24, 6, 10, 26], we propose a unified representation, Factorized Fields, with generalizable Coefficient Backbone and Basis Transformer.
 This approach explicitly captures multi-scale visual features and repetitive structural components in images through a basis-coefficient decomposition.
@@
 The main contributions of this paper are summarized as follows:
-  • We propose Factorized Fields, a unified representation that explicitly models multi-scale visual features and structural components for both super-resolution and image compression.
+  • We propose Generalized Factorized Fields, a unified representation that explicitly models multi-scale visual features and structural components for both super-resolution and image compression.
   • We integrate super-resolution with image compression by introducing SR prior during decompression to compensate for lost details and developing a merged-basis compression branch for multi-image compression.
@@
 Image super-resolution is critical in computer vision, focusing on recovering high-resolution (HR) images from low-resolution (LR) inputs.
-Following the foundational studies, CNN-based strategies SRCNN, flickr2k, zhou2020cross, shufflemixer, VDSR were initially introduced with modeling techniques such as residual learning ledig2017photo, liu2020residual, tong2017image, zhang2018residual, RCAN, drct, EDSR , or recursive learning RGT, tai2017recursive.
-Besides, subsequent research also sheds light on GAN-base methods srgan, ledig2017photo, esrgan, real_esrgan to enhance realisticity and detail quality.
-However, the inductive bias of CNN-based networks by restricting spatial locality hinders the capture of long-range dependency from images, which is alleviated by Transformer-based SISR networks IPT, EDT.
-Afterward, SwinIR swinir is proposed to combine spatial locality and non-local information by Swin Transformer liu2021swin with window attention and achieve breakthrough improvement in SISR.
-Following SwinIR’s success, several works have built upon its framework swin2sr, zhu2023attention, ATD to reach better image quality as well as solve information bottleneck drct.
-Hybrid approaches CRAFT (CRAFT) merge the benefits of convolutional and transformer structures to further elevate SR performance.
-For better feature aggregation, DAT (DAT) and HAT (chen2023hat) integrate spatial and channel information using attention mechanisms to enhance their representation capabilities.
-Moreover, RGT (RGT) introduces a unique recursive-generalization self-attention mechanism that efficiently captures comprehensive spatial details with a linear increase in computational complexity.
+Following the foundational studies, CNN-based strategies [21, 46, 79, 64, 34] were initially introduced with modeling techniques such as residual learning [40, 47, 68, 78, 77, 18, 45] , or recursive learning [16, 65].
+Besides, subsequent research also sheds light on GAN-base methods [39, 40, 70, 71] to enhance realisticity and detail quality.
+However, the inductive bias of CNN-based networks by restricting spatial locality hinders the capture of long-range dependency from images, which is alleviated by Transformer-based SISR networks [12, 43].
+Afterward, SwinIR [44] is proposed to combine spatial locality and non-local information by Swin Transformer [50] with window attention and achieve breakthrough improvement in SISR.
+Following SwinIR’s success, several works have built upon its framework [19, 80, 75] to reach better image quality as well as solve information bottleneck [18].
+Hybrid approaches CRAFT [41] merge the benefits of convolutional and transformer structures to further elevate SR performance.
+For better feature aggregation, DAT [15] and HAT [13] integrate spatial and channel information using attention mechanisms to enhance their representation capabilities.
+Moreover, RGT [16] introduces a unique recursive-generalization self-attention mechanism that efficiently captures comprehensive spatial details with a linear increase in computational complexity.
 Image Compression (IC).
-Deep learning has significantly advanced image compression, offering superior compression ratios and image quality compared to traditional methods like JPEG jpeg and JPEG-2000 jpeg2000.
-Early CNN-based approaches imc_cnn1, imc_cnn2, imc_cnn3 have been surpassed by transformer-based models imc_trans1, imc_trans2, imc_trans3, duan2023cross-component, which leverage spatio-channel attention for better performance.
-GAN-based methods imc_gan1, imc_gan2, imc_gan3, imc_gan4 have further contributed to real-time adaptive compression.
-Recently, ELIC imc_elic introduced efficient compression with unevenly grouped space-channel contextual adaptive coding, while LIC TCM imc_lic_tcm integrated transformers and CNNs to capture both global and local image features.
-The eContextformer koyuncu2024efficient introduced patch-wise, checkered, and channel-wise grouping techniques for parallel context modeling with a shifted window spatial-channel attention mechanism.
-GroupedMixer li2024groupedmixer proposed a transformer-based entropy model with token mixers for inner and cross-group context modeling.
-Meanwhile, Wavelet Conditional Diffusion song2024wavelet-compression introduced a wavelet-based model with uncertainty-aware loss, balancing high perceptual quality with low distortion.
+Deep learning has significantly advanced image compression, offering superior compression ratios and image quality compared to traditional methods like JPEG [69] and JPEG-2000 [66].
+Early CNN-based approaches [7, 3, 28] have been surpassed by transformer-based models [37, 81, 72, 22], which leverage spatio-channel attention for better performance.
+GAN-based methods [53, 54, 23, 62] have further contributed to real-time adaptive compression.
+Recently, ELIC [30] introduced efficient compression with unevenly grouped space-channel contextual adaptive coding, while LIC TCM [48] integrated transformers and CNNs to capture both global and local image features.
+The eContextformer [38] introduced patch-wise, checkered, and channel-wise grouping techniques for parallel context modeling with a shifted window spatial-channel attention mechanism.
+GroupedMixer [42] proposed a transformer-based entropy model with token mixers for inner and cross-group context modeling.
+Meanwhile, Wavelet Conditional Diffusion [63] introduced a wavelet-based model with uncertainty-aware loss, balancing high perceptual quality with low distortion.
 In summary, while recent super-resolution and image compression advances focus on increasingly complex architectures, our work takes a different approach.
@@
 3 Methods
-In this section, we first briefly introduce the background of the factor fields factor-field, their properties, and also the preliminary of Learned Image Compression in Sec.
+In this section, we present the details of our proposed methods. Inspired by factor fields, which can be formulated by
+$\hat{s}(x)=\mathcal{P}\Bigl{(}\textbf{Concat}_{i=1}^{N}\Bigl{\{}c_{i}(x)\odot b_{i}(\gamma_{i}(x))\Bigl{\}}\Bigl{)}.\end{aligned}$
+and whose details can be found in supplementary materials , we first explain the motivation and how we derive our Factorized Fields for enhanced image reconstruction quality in  Sec.
 3.1.
-Inspired by factor fields, we next explain the motivation and how we derive our Factorized Fields for enhanced image reconstruction quality in Sec.
+We then describe how to adapt the formulation to Super-Resolution in  Sec.
 3.2.
-We then describe how to adapt the formulation to Super-Resolution in Sec.
+Finally, we show how to incorporate such representation with image(s) compression and how we integrate Super-Resolution and Image Compression in  Sec.
 3.3.
-Finally, we show how to incorporate such representation with image(s) compression and how we integrate Super-Resolution and Image Compression in Sec.
-3.4.
+Note that the background of learned image compression is described in supplementary materials.
 Figure 2: The correlation between coordinate transformation and downsampling.
@@
 (c) To explicitly model the information for sampling with a sawtooth, we rearrange the feature map in a dilation-like manner in the downsample layer of the Basis Swin Transformer.
-This way, the feature sampled would capture the information in the original layout correctly.
-3.1 Preliminaries
-Factor Fields.
-The concept of decomposition fields or factorized matrices in the reconstruction of 2D images or 3D scenes has shown superior rendering quality and improved efficiency recently tensorf, gao2023strivec, mueller2022instantngp, kplanes, cao2023hexplane, eg3d, while factor-field first proposed a unified representation called factor fields.
-For a 1D signal $s(x)$, the factor fields formulation is:
-$\hat{s}(x)=c(x)^{T}b(\gamma(x)),$ (1)
-where $c(x)=(c_{1}(x),...,c_{N}(x))^{T}$ are spatial-varying coefficient fields, $b(x)=(b_{1}(x),...,b_{N}(x))^{T}$ are basis functions, and $\gamma(x)$ is a coordinate transformation.
-The coefficient and basis are queried by sampling with the coordinates $x$.
-This generalizes to Q-dimensional signals:
-$\hat{s}(x)=\mathcal{P}(c(x)\odot b(\gamma(x))),$ (2)
-where $\mathcal{P}$ is a projection function and $\odot$ denotes the element-wise product. In practice, for multi-dimensional coefficients:
-$\hat{s}(x)=\mathcal{P}\Bigl{(}\textbf{Concat}_{i=1}^{N}\Bigl{\{}c_{i}(x)\odot b_{i}(\gamma_{i}(x))\Bigl{\}}\Bigl{)}.\end{aligned}$
-This formulation allows the same basis to be applied at multiple spatial locations while varying the coefficients, especially when $\gamma$ is a periodic function.
-Learned Image Compression.
-Following (channel-autoregressive, ballé2018variational), a learned image compression model with a channel-wise entropy model can be formulated as:
-$z$ $=h_{a}(y;\phi_{h}),\ y=g_{a}(x;\phi),$ (4)
-$\{X_{\text{mean}},X_{\text{scale}}\}$ $=h_{s}(\hat{z};\theta_{h}),\ \hat{z}=Q(z)$
-$\hat{y}$ =$=\{Q(y_{0}-\mu_{0})+\mu_{0},...,Q(y_{t}-\mu_{t})+\mu_{t}\},\ 0<=t<l,\ \mu_{t}=e_{i}(\hat{y}_{<i},X_{\text{mean}}))$
-$\hat{x}$ $=g_{s}(\overline{y};\theta),\ \overline{y}=\textbf{Refine}_{\theta_{r}}({\mu_{0},...,\mu_{t}},\hat{y}).$
-The encoder $g_{a}$ transforms the raw image $x$ into a latent representation $y$.
-A hyper-prior encoder $h_{a}$ further processes $y$ to output $z$, capturing spatial dependencies.
-$z$ is quantized to $\hat{z}$, which is decoded by $h_{s}$ to produce features $X_{\text{mean}}$ and $X_{\text{scale}}$, used to estimate the mean $\mu$ and variance $\sigma$ of $y$.
-The latent $y$ is divided into $l$ slices, and each quantized around computed means $\mu_{t}$.
-These means are derived from earlier quantized slices and $X_{\text{mean}}$ by a slice network $e_{i}$.
-The quantized slices form $\hat{y}$.
-For decompression, $\hat{y}$ is refined using $\textbf{Refine}_{\theta_{r}}$ based on $\mu_{t}$ and $\hat{y}$ to produce $\overline{y}$, approximating the original $y$.
-Finally, $g_{s}$ reconstructs the decompressed image $\hat{x}$ from $\overline{y}$.
-The model is trained using a Lagrangian multiplier-based rate-distortion optimization:
-$L=R(\hat{y})+R(\hat{z})+\lambda\cdot D(x,\hat{x}),$ (5)
-where $R(\hat{y})$ and $R(\hat{z})$ denote bit rates, $D(x,\hat{x})$ is the distortion term (calculated by MSE), and $\lambda$ balances compression efficiency and image fidelity.
-In our experiments, we follow imc_lic_tcm, modifying only $g_{s}$ to demonstrate our representation’s effectiveness.
+This way, the feature sampled would capture the information in the original layout correctly; in other words, $I=\text{Sawtooth-Pixel-Sampling}(\text{Sawtooth-Aware-Downsample}(I))$
 Figure 3: The overall pipeline of image super-resolution with our Factorized Fields.
@@
 Finally, as in Eq.
-8, the coefficient and basis are sampled, multiplied, and decoded for final high-resolution output $I_{HR}$, where $s$, $H$, $W$ denote the scale factor, height, and width respectively.
+4, the coefficient and basis are sampled, multiplied, and decoded for final high-resolution output $I_{HR}$, where $s$, $H$, $W$ denote the scale factor, height, and width respectively.
 3.1 Formulation of Factorized Fields
@@
 1, the key to superior rendering quality in image regression tasks of both Image Compression and Super-Resolution lies in the capability of the representation to capture accurate structural distribution and fine visual details.
-Meanwhile, transformations such as Fourier Transform or Wavelet Transform have long been used to model multi-frequency information fuoli2021fourier, waveletsr to express different implicit functions in images; however, such methods often suffer from under-expression due to the pre-defined and limited frequency bands or restriction from its formulation.
+Meanwhile, transformations such as Fourier Transform or Wavelet Transform have long been used to model multi-frequency information [25, 36] to express different implicit functions in images; however, such methods often suffer from under-expression due to the pre-defined and limited frequency bands or restriction from its formulation.
 Thus, we seek a representation that explicitly incorporates and fits multi-scale and multi-frequency components and yet is highly flexible and learnable according to individual features in an image.
-Inspired from recent success on such decomposition fields kplanes, factor-field, tensorf, mueller2022instantngp, cao2023hexplane, eg3d, we primarily built our Factorized Fields framework on factor fields factor-field from Eq.
-3:
+Inspired from recent success on such decomposition fields [24, 10, 9, 59, 6, 8], we primarily built our Factorized Fields framework on factor fields [10] from  Eq. 1:
 $\hat{I}(x)=\mathcal{P}\Bigl{(}\textbf{Concat}_{i=1}^{N}\Bigl{\{}c_{i}(x)\odot b_{i}(\gamma_{i}(x))\Bigl{\}}\Bigl{)}.\end{aligned}$
@@
 Such formulation has several key properties: First of all, by decomposing the images into basis frequencies, we can learn the implicit functions of an image and capture the mutual dependencies between pixels and across spatial composition; meanwhile, since the basis and coefficient are specific to every single image and both learnable in all spatial dimensions, the restriction on a limited number of basis (we use $N=6$ in all of our experiments) can be alleviated.
-Finally, as in factor-field, of all the tested coordinate transformation $\gamma$ in Eq.
-3 sawtooth $\gamma(x)=x\bmod{k},k\in\mathbb{R}$ performs mostly the best in image regression tasks.
-We can easily observe that such transformation implicitly captures patch-like frequency information as shown in Fig.
-1a, and thus, we propose that by leveraging the inter-patch information from the sawtooth coordinate transformation, the visual correspondence between spatial locations can be effectively represented.
-However, in practice, we sample the basis via bilinear or bicubic sampling due to memory constraints, i.e., the feature size is less than image height and width, and this poses a severe problem: These nearby sampled pixel features are actually linear or based on cubic interpolation with respect to the basis, and such inductive bias hinders the representation of non-linearity in images.
+A key insight of our work is the use of sawtooth coordinate transformation $\gamma(x)=x\bmod{k},k\in\mathbb{R}$, which performs best among tested transformations in image regression tasks.
+The sawtooth transformation maps input coordinates to a periodic space, implicitly capturing patch-like frequency information as shown in Fig.
+2a.
+However, traditional downsampling in neural networks fails to preserve this periodic structure.
+To address this, we introduce sawtooth-aware downsampling as in Fig.
+2c with our Basis Swin Transformer.
+This rearrangement allows the network to learn correlations between the periodic features in Sawtooth sampling.
+Moreover, in practice, we sample the basis via bilinear or bicubic sampling due to memory constraints, i.e., the feature size is less than image height and width, and this poses a severe problem: These nearby sampled pixel features are actually linear or based on cubic interpolation with respect to the basis, and such inductive bias hinders the representation of non-linearity in images.
 To resolve this, we propose a modified version, our Factorized Fields:
-$\hat{I}(x)=\mathcal{P}\Bigl{(}{\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl{\{}c_{i}(x)\odot\psi(\alpha_{j}\cdot b_{i}(\gamma_{i}(x)))\Bigl{\}}\Bigl{)},\psi\in\{\text{sin},\text{cos}\},\alpha_{j}\in R.$ (7)
-Here, inspired by the Fourier Series, with a scalar $\alpha$ which can be viewed as increasing to sampling frequency and a transformation function $\psi$ to add to implicit non-linearity functions and modulation of product, our complete Factorized Fields method stands out by the fact that such formulation significantly enhances the ability to represent complex non-linear structures in images and effectively composes high-frequency components between pixels.
+$\hat{I}(x)=\mathcal{P}\Bigl{(}{\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl{\{}c_{i}(x)\odot\psi(\alpha_{j}\cdot b_{i}(\gamma_{i}(x)))\Bigl{\}}\Bigl{)},$ (3)
+where $\psi\in\{\text{sin},\text{cos}\},\alpha_{j}\in R$.
+Here, inspired by the Fourier Series, with a scalar $\alpha$ which corresponds to an increase in sampling frequency and a transformation function $\psi$ to add to implicit non-linearity functions and modulation of product, our complete Factorized Fields method stands out by the fact that such formulation significantly enhances the ability to represent complex non-linear structures in images and effectively composes high-frequency components between pixels.
+Refer to captionRefer to caption
+Figure 4: Comparison of Different Representation This figure illustrates the performance comparison on a single-image regression task using three different representations: vanilla factor fields in Eq.
+1, our Factorized Fields (Eq.
+3), and Plain Fields, which use simple channel-wise tensor concatenation instead of coefficient-basis multiplication.
+The left panel shows the MAE on the magnitude of different frequency components obtained from the Fourier Transform, while the right panel shows PSNR along training.
+Although the basis-coefficient decomposition method in [10] yields improvements, it still struggles with slow convergence and inconsistencies between low and high-frequency components.
+In contrast, by introducing explicit formulations for $\alpha$ and $\psi$, our method aligns different frequency components effectively and achieves superior convergence speed.
+This advantage becomes especially crucial when the task extends to generalizable applications, such as super-resolution or image compression.
+Our Factorized Fields enhance the original Factor Fields [10] in three key ways: (1) non-linear transformation through $\psi\in\{\text{sin},\text{cos}\}$ enables better modeling of complex image patterns, (2) multi-frequency modulation using $\alpha_{j}$ allows explicit control over frequency components, and (3) sawtooth-aware downsampling preserves periodic structure during feature processing.
+As shown in Fig.
+4, our refined representation contributes to large margin of improvement in convergence and performance, demonstrating the modification is the key to superior information modeling.
 Figure 5: The illustration of our joint image-compression and super-resolution framework compared with the traditional compression-only method.
 (a) Traditional learning-based compression methods.
-(b) Our approach surpasses (a) by incorporating our Super-Resolution (SR) Module from Sec.3.3 as information-recovery prior, detailed in Sec.
-3.4.1.
+(b) Our approach surpasses (a) by incorporating our Super-Resolution (SR) Module from Sec.
+3.2 as information-recovery prior, detailed in Sec.
+3.3.1.
 (c) Expanding on (b),  Sec.
@@
 As shown in  Fig.
-2, the process begins with $I_{LR}\in\mathbb{R}^{3\times H\times W}$.
+3, the process begins with $I_{LR}\in\mathbb{R}^{3\times H\times W}$.
 The Coefficient Backbone extracts features $X_{\text{coeff}}\in\mathbb{R}^{C_{c}\times H_{c}\times W_{c}}$, which are then used to generate coefficients $c$ through convolution and pixel shuffle operations, and fed into the Basis Swin Transformer to produce a multi-scale basis $b=\{b_{1},...,b_{N}\},\ b_{i}\in\mathbb{R}^{C_{b_{i}}\times H_{b_{i}}\times W_{b_{i}}}$.
 The coefficients and basis are combined to reconstruct $I_{SR}\in\mathbb{R}^{3\times sH\times sW}$ using  Eq.
-8, where $s$ is the scale factor.
+4, where $s$ is the scale factor.
 In our experiments, $\gamma_{i}(x)$ is a Sawtooth function: $\gamma_{i}(x)=x\mod k_{i}$, with $k_{i}\in R$.
 We optimize model parameters using an $L_{1}$ loss function.
-To demonstrate our method’s effectiveness, we use existing SR methods chen2023activating, DAT, ATD as the Coefficient Backbone.
-For the Basis Swin Transformer, we employ Swin Transformer Blocks liu2022swinv2 with a series of downsampling operations.
+To demonstrate our method’s effectiveness, we use existing SR methods [14, 15, 75] as the Coefficient Backbone.
+For the Basis Swin Transformer, we employ Swin Transformer Blocks [51] with a series of downsampling operations.
 We use a dilation-like downsampling technique( Fig.
-1(c)) to accommodate the sawtooth sampling pattern.
+2c) to accommodate the sawtooth sampling pattern.
 The final basis is refined using additional upsampling and convolution layers.
@@
 However, as discussed in  Sec.
-1,  2, and 3.2, most recent works draw emphasis on how to retrieve implicit elements through designing different architectures, such as analysis transforms and entropy models koyuncu2024efficient, li2024groupedmixer, or decompression modules song2024wavelet-compression, duan2023cross-component, while this paper aims at addressing the representation itself to better capture the structural correlations and thus achieve better image quality through explicit modeling of frequency components and Factorized Fields formulation as in Eq.
-7.
+1,  Sec.
+2, and Sec.
+3.1, most recent works draw emphasis on how to retrieve implicit elements through designing different architectures, such as analysis transforms and entropy models [38, 42], or decompression modules [63, 22], while this paper aims at addressing the representation itself to better capture the structural correlations and thus achieve better image quality through explicit modeling of frequency components and Factorized Fields formulation as in  Eq.
+3.
 In addition, with our trained SR model described in  Sec.
-3.3, it intuitively serves as a strong prior for information recovery, i.e., it contains extensive knowledge of how to reconstruct missing details and enhance image quality by leveraging learned patterns from the training data.
+3.2, it intuitively serves as a strong prior for information recovery, i.e., it contains extensive knowledge of how to reconstruct missing details and enhance image quality by leveraging learned patterns from the training data.
 Thus, since Super-Resolution and Image Compression share the core principle of reconstructing and enhancing image details from low-quality sources, we can effectively integrate this prior into the compression pipeline.
@@
 The overall pipeline is shown in  Fig.
-3(b).
-To demonstrate the robustness of our representation and the effectiveness of the SR prior, the compression and decompression networks greatly follow imc_lic_tcm, with only the synthesis transform replaced by our SR pipeline, where the details can be referenced in Supplementary Materials.
+5b.
+To demonstrate the robustness of our representation and the effectiveness of the SR prior, the compression and decompression networks greatly follow [48], with only the synthesis transform replaced by our SR pipeline, where the details can be referenced in Supplementary Materials.
 In practice, the training is performed in two stages.
-After we obtain the trained SR prior, the model is fine-tuned with a lower learning rate alongside the compression module, which is then trained end-to-end with the loss function defined in Eq.
-5.
+After we obtain the trained SR prior, the model is fine-tuned with a lower learning rate alongside the compression module, which is then trained end-to-end with the loss function defined as
+$L=R(\hat{y})+R(\hat{z})+\lambda\cdot D(x,\hat{x}),$ (5)
 3.3.2 Multi-Image Compression
@@
 Given $M$ images and their respective bases $b^{n}_{i},n\in\{1,...,M\}$, we apply a Basis Merging Transformer $F_{\text{merge}}$ at each location to integrate the $M$ elements:
-$b_{i}(h,w)=F_{\text{merge}}\left(\{b^{n}_{i}(h,w)\ |\ n\in{1,...,M}\}\right),\ 0\leq h<H_{b_{i}},\ 0\leq w<W_{b_{i}}.$ (9)
-Here, $F_{\text{merge}}$ is a standard transformer, similar in architecture to that described in (dinov2).
-We treat the bases as tokens and prepend the sequence with a $CLS$ token, concluding the sequence to form the final merged basis.
+$b_{i}(h,w)=F_{\text{merge}}\left(\{b^{n}_{i}(h,w)\ |\ n\in{1,...,M}\}\right),$ (6)
+[…]
 Since compression induces error, according to  Fig.
-3b, the Coefficient feature map $X_{\text{coeff}}$ generated with Coefficient Backbone contains misinformation, and Basis Swin Transformer that uses such features would capture the information wrongfully for the basis and thus amplify error.
+5b, the Coefficient feature map $X_{\text{coeff}}$ generated with Coefficient Backbone contains misinformation, and Basis Swin Transformer that uses such features would capture the information wrongfully for the basis and thus amplify error.
 To solve this, we exploit the mergeable property of basis, compress and transmit the merged basis independently along other quantized variables $\{Q(y,\mu),\ \hat{z}\}$, and finally reconstruct the $M$ images with the merged basis and their individual decoded coefficients.
 as illustrated in Fig.
-3c.
+5c.
 This way, we can enjoy less information loss with basis while maintaining low bit rates.
@@
 The best results are colored red.
-The models with †are those who use same-task pretraining chen2023activating.
+The models with †are those who use same-task pretraining [14].
 Please refer to quantitative results in  Sec.
@@
 […]
-SwinIR (swinir)(ICCV2021W) 28.01 119.68 3,826 32.93 0.9043 29.15 0.7958 27.95 0.7484 27.56 0.8273 32.22 0.9273
-ATD (zhang2024transcending)(CVPR2024) 20.26 77.10 6,572 33.14 0.9061 29.25 0.7976 28.02 0.7524 28.22 0.8414 32.65 0.9308
-DAT (DAT))(ICCV2023) 14.80 61.66 4,192 33.15 0.9062 29.29 0.7983 28.03 0.7518 27.99 0.8365 32.67 0.9301
-RGT (RGT)(ICLR2024) 13.37 834.25 3,404 33.16 0.9066 29.28 0.7979 28.03 0.7520 28.09 0.8388 32.68 0.9303
-HAT (chen2023hat)(CVPR2023) 20.77 86.02 3,692 33.18 0.9073 29.38 0.8001 28.05 0.7534 28.37 0.8447 32.87 0.9319
-HAT-L (chen2023hat)(CVPR2023) 40.84 167.27 6,804 33.30 0.9083 29.47 0.8015 28.09 0.7551 28.60 0.8498 33.09 0.9335
+SwinIR [44](ICCV2021W) 28.01 119.68 3,826 32.93 0.9043 29.15 0.7958 27.95 0.7484 27.56 0.8273 32.22 0.9273
+ATD [76](CVPR2024) 20.26 77.10 6,572 33.14 0.9061 29.25 0.7976 28.02 0.7524 28.22 0.8414 32.65 0.9308
+DAT [15])(ICCV2023) 14.80 61.66 4,192 33.15 0.9062 29.29 0.7983 28.03 0.7518 27.99 0.8365 32.67 0.9301
+RGT [16](ICLR2024) 13.37 834.25 3,404 33.16 0.9066 29.28 0.7979 28.03 0.7520 28.09 0.8388 32.68 0.9303
+HAT [13](CVPR2023) 20.77 86.02 3,692 33.18 0.9073 29.38 0.8001 28.05 0.7534 28.37 0.8447 32.87 0.9319
+HAT-L [13](CVPR2023) 40.84 167.27 6,804 33.30 0.9083 29.47 0.8015 28.09 0.7551 28.60 0.8498 33.09 0.9335
 ATD-L* 49.42 184.83 15,582 33.12 0.9062 29.31 0.7985 28.02 0.7514 28.25 0.8422 32.78 0.9309
@@
 We conduct extensive experiments to validate the effectiveness of our Factorized Fields representation for Super-Resolution tasks.
-Following the strategy outlined in (chen2023hat) and (drct), we adapt the same-task pretraining approach for all the Super-Resolution models.
-Unlike these previous works, we leverage the SA-1B dataset from (sam), which includes approximately 10 million images, where we randomly sample 4 million for training, which is much less that of ImageNet (imagenet), which is used for chen2023hat and drct pretraining.
+Following the strategy outlined in [13] and [18], we adapt the same-task pretraining approach for all the Super-Resolution models.
+Unlike these previous works, we leverage the SA-1B dataset from [35], which includes approximately 10 million images, where we randomly sample 4 million for training, which is much less that of ImageNet [20], which is used for [13] and [18] pretraining.
 Note that we use SA-1B just for its content-rich and high-resolution images since SA-1B is designed for training segmentation models, whereas ImageNet focuses primarily on single-class prediction.
@@
 While the performances are mostly equivalent, we witness faster convergence with SA-1B during training.
-To be more focused on the representation itself, we utilize various pre-trained SR models (SwinIR (swinir), HAT (chen2023activating), DAT (DAT), and ATD (zhang2024transcending)) as the Coefficient Backbone, with a consistent Basis Transformer architecture.
+To be more focused on the representation itself, we utilize various pre-trained SR models (SwinIR [44], HAT [14], DAT [15], and ATD [76]) as the Coefficient Backbone, with a consistent Basis Transformer architecture.
 Training involves initializing the Coefficient Backbones from pre-trained SR models and randomly initializing the Basis Transformers.
 Models are pre-trained for 300k iterations on the SA-1B dataset.
-After pretraining, we use DF2K (DIV2K (div2k) + Flickr2K (flickr2k)) as the finetuning dataset following (swin2sr, chen2023hat) for 200k iterations.
+After pretraining, we use DF2K (DIV2K [1] + Flickr2K [46]) as the finetuning dataset following [19, 13] for 200k iterations.
 For pretraining, we utilize AdamW optimizer with learning rate 1e-4, batch size 16, betas (0.9, 0.99), and other parameters set to PyTorch default, while we use learning rate 1e-5 during finetuning.
@@
 As for the hyperparameter of Factorized Fields, the number $N$ of coefficient and basis is set to 6, and the scalars $\alpha_{j}$ are set to {1, 4, 16, 64} since we want to capture both base frequency $\alpha_{j}=1$ and high frequency $\alpha_{j}=64$ information.
+More implementation details can be found in supplementary materials.
 Quantitative Results.
@@
 1 presents the quantitative comparison between our approach and state-of-the-art (SoTA) methods.
-We evaluate the methods using five benchmark datasets, including Set5 (set5), Set14 (set14), BSD100 (bsd100), Urban100 (urban100), and Manga109 (manga109).
+We evaluate the methods using five benchmark datasets, including Set5 [5], Set14 [74], BSD100 [55], Urban100 [31], and Manga109 [56].
 For quantitative metrics, PSNR and SSIM are reported.
@@
 This formula measures the relative performance gain of our model compared to the gap between HAT-L and SwinIR.
-To validate the effectiveness of our framework, we employ three different SoTA SR models—ATD ATD, DAT DAT, and HAT chen2023activating—as the Coefficient Backbone in our pipeline, which we denote as ATD-F, DAT-F, and HAT-F, respectively.
+To validate the effectiveness of our framework, we employ three different SoTA SR models—ATD [75], DAT [15], and HAT [14]—as the Coefficient Backbone in our pipeline, which we denote as ATD-F, DAT-F, and HAT-F, respectively.
 These models exhibit significant improvements when compared to their counterparts ATD-L, DAT-L, and HAT-L, which possess similar parameter counts.
-It is important to note that only chen2023activating provides a large-scale model.
+It is important to note that only [14] provides a large-scale model.
 To maintain fairness, we scale up ATD and DAT to match this size and train them under the same configuration as our model, including both pretraining and fine-tuning stages.
-To ensure an equitable comparison with other methods, we further train another model, HAT-F-ImageNet, using ImageNet as the pretraining dataset, following the protocols outlined in IPT, chen2023activating.
+To ensure an equitable comparison with other methods, we further train another model, HAT-F-ImageNet, using ImageNet as the pretraining dataset, following the protocols outlined in [12, 14].
 The results demonstrate that its performance remains consistent with only minor perturbations.
-Furthermore, in traditional Fourier Series and other image processing methods jpeg, the basis is typically derived first and then used to compute the coefficients.
+Furthermore, in traditional Fourier Series and other image processing methods [69], the basis is typically derived first and then used to compute the coefficients.
 In contrast, our method derives the coefficient features first, as illustrated in  Fig.
-2.
+3.
 To explore this difference, we develop another variant of our model, denoted HAT-F-Basis-First, where we reverse the order of operations.
@@
 Lastly, to evaluate the effectiveness of our Factorized Fields, we trained a model named HAT-F-Concat, which does not apply the formulation in Eq.
-7.
+3.
 Instead, it concatenates the basis and coefficient directly and decodes the resulting features to produce the output.
@@
 We provide the visual comparison in Fig.
-4.
+6.
 The images are randomly sampled from the DIV2K dataset.
 Our method faithfully reconstructs the image details, whereas the other approaches suffer from over-smoothing or hallucinating details absent in the ground truth.
-Table 2: Comprehensive evaluation for image compression. Using VTM as an anchor for calculating BD-Rate. Latencies are measured under an NVIDIA GTX 3090 GPU.
+Multi-frame SR.
+Due to the space limit, we provide additional comparisons on multi-frame SR in the supplementary materials.
+(a) Kodak
+(b) CLIC
+(c) Tecnick
+Figure 7: Performance (RD-Curve) evaluation on image compression using different datasets.
 Method BD-Rate (%) $\downarrow$ Latency(s) Params(M)
@@
 TCM-HAT-F-multi M=24 (Ours) -20.97 0.232 0.174 131.35
-{subfigure}
-[b]0.32 Refer to caption {subfigure}[b]0.32 Refer to caption {subfigure}[b]0.32 Refer to caption
-Figure 5: Kodak
-Figure 6: CLIC
-Figure 7: Tecnick
-Figure 8: Performance (RD-Curve) evaluation on image compression using different datasets.
+Table 2: Comprehensive evaluation for image compression. Using VTM as an anchor for calculating BD-Rate. Latencies are measured under an NVIDIA GTX 3090 GPU.
+Method Kodak CLIC Tecnick
+$\lambda=0.0025$ $\lambda=0.0067$ $\lambda=0.025$ $\lambda=0.0025$ $\lambda=0.0067$ $\lambda=0.025$ $\lambda=0.0025$ $\lambda=0.0067$ $\lambda=0.025$
+[…]
+TCM[48] 0.1533 30.0834 0.2983 32.5841 0.6253 36.1345 0.1214 31.8207 0.2235 34.2098 0.4503 37.1201 0.1268 32.0588 0.2193 34.3669 0.3981 36.9066
+TCM-HAT-F-Scratch 0.1570 30.0857 0.2976 32.5893 0.6211 36.1389 0.1214 31.9421 0.2235 34.2894 0.4503 37.1434 0.1258 32.0632 0.2189 34.3781 0.4001 36.9223
+TCM-HAT 0.1567 30.1843 0.2992 32.6454 0.6268 36.2267 0.1220 31.9737 0.2266 34.3319 0.4512 37.2486 0.1262 32.1423 0.2174 34.5124 0.3971 36.9934
+TCM-HAT-F 0.1574 30.4012 0.2998 32.8910 0.6276 36.4461 0.1229 32.1917 0.2249 34.4109 0.4512 37.3135 0.1255 32.4591 0.2186 34.7656 0.3975 37.3244
+Table 3: Validation of the effectiveness of SR prior. The best PSNRs are marked in red.
 4.2 Single- and Multi-Image Compressions
@@
 4.1, we use the same set of SA-1B for training.
-To emphasize our representation, we initialize compression and decompression modules in Fig.3 from pre-trained imc_lic_tcm and the SR Module from those in Sec.
-4.1, we then train the pipeline end to end on 256x256 patches for 200k iterations, with AdamW optimizer (loshchilov2017decoupled) with learning rate 1e-5, batch size 16, betas (0.9, 0.99) and other parameters set to PyTorch default.
+To emphasize our representation, we initialize compression and decompression modules in Fig.
+5 from pre-trained [48] and the SR Module from those in Sec.
+4.1, we then train the pipeline end to end on 256x256 patches for 200k iterations, with AdamW optimizer [52] with learning rate 1e-5, batch size 16, betas (0.9, 0.99) and other parameters set to PyTorch default.
 We integrate our SR Module with the pre-trained compression module TCM, creating TCM-HAT-F and TCM-HAT-L-F models.
@@
 For multi-image compression in Fig.
-3(c), we set the Basis Swin Transformer and the Basis Merging Transformer to be trainable while the other parts remain frozen.
+5.c, we set the Basis Swin Transformer and the Basis Merging Transformer to be trainable while the other parts remain frozen.
 Rate-Distortion Performance Comparison.
-We compare our model with State-of-the-Art learned end-to-end image compression algorithms, including (liu2023learned), (Chen2022), zou2022the, xie2021enhanced, cheng2020learned, ballé2018variational, li2024groupedmixer, jiang2023mlic, channel-autoregressive, bellard_bpg, qi2023hyperthumbnail, and imc_elic.
-The classical image compression codec, VVC (vtm), is also tested by using VTM12.1.The rate-distortion performance on various datasets, including Kodak, Tecnick’s old test set with resolution 1200$\times$1200, and CLIC Professional Validation, is shown in Fig.
-8.
+We compare our model with State-of-the-Art learned end-to-end image compression algorithms, including [49], [11], [82], [73], [17], [2], [42], [32], [57], [4], [61], and [30].
+The classical image compression codec, VVC [67], is also tested by using VTM12.1.The rate-distortion performance on various datasets, including Kodak, Tecnick’s old test set with resolution 1200$\times$1200, and CLIC Professional Validation, is shown in Fig.
+7.
 In  Tab.
@@
 Furthermore, our method maintains competitive latency while significantly improving compression performance and balancing efficiency and quality.
-Table 3: Comparison of improvements of Factorized Fields. $\psi$ and $\alpha$ are the same in Eq. 8
 Metric PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-Baseline (factor-field) 22.04 0.505 0.5296
+Baseline [10] 22.04 0.505 0.5296
 Ours 38.44 0.999 0.0385
@@
 No $\alpha$ for pixel-wise frequency information 21.25 0.537 0.527
+Table 4: Comparison of improvements of Factorized Fields. $\psi$ and $\alpha$ are the same in Eq. 4.
 4.3 Ablation Studies
+Influence of SR Priors in Image Compression.
+We conduct experiments with various configurations to verify the proposed image compression pipeline’s effectiveness.
+As shown in Tab.
+3, we present the quantitative performance of models trained with different values of $\lambda$ in Eq.
+5.
+Specifically, TCM-HAT refers to substituting our SR Module with the original HAT [14] in the pipeline illustrated in Fig.
+5b.
+TCM-HAT-F represents our complete pipeline, while TCM-HAT-F-Scratch denotes the same pipeline but with the SR Module initialized randomly.
+Our results demonstrate that integrating SR priors with image compression improves performance, and our proposed representation further enhances results.
+This highlights the robustness of our Factorized Fields in capturing fine details in image regression tasks.
 Effectiveness of Factorized Fields Design.
@@
 1 to Eq.
-7.
+3.
 The quantitative performance reported on single-image regression is shown in Tab.
-3, where each result is measured after 256 iterations.
+4, where each result is measured after 256 iterations.
 Compared to baseline results, our refinements in modeling pixel-level frequency have significantly improved all performance metrics.
 Additionally, our results demonstrate that the modulation function $\psi$ and the scalar $\alpha$ are interdependent, each essential to the other’s function.
-Influence of SR Priors in Image Compression.
-We conduct experiments with various configurations to verify the proposed image compression pipeline’s effectiveness.
-As shown in Tab.
-4, we present the quantitative performance of models trained with different values of $\lambda$ in Eq.
-5.
-Specifically, TCM-HAT refers to substituting our SR Module with the original HAT chen2023activating in the pipeline illustrated in Fig.
-3.b.
-TCM-HAT-F represents our complete pipeline, while TCM-HAT-F-Scratch denotes the same pipeline but with the SR Module initialized randomly.
-Our results demonstrate that integrating SR priors with image compression improves performance, and our proposed representation further enhances results.
-This highlights the robustness of our Factorized Fields in capturing fine details in image regression tasks.
-Table 4: Validation of the effectiveness of SR prior. The best PSNRs are marked in red.
-Method Kodak CLIC Tecnick
-$\lambda=0.0025$ $\lambda=0.0067$ $\lambda=0.025$ $\lambda=0.0025$ $\lambda=0.0067$ $\lambda=0.025$ $\lambda=0.0025$ $\lambda=0.0067$ $\lambda=0.025$
-[…]
-TCMimc_lic_tcm 0.1533 30.0834 0.2983 32.5841 0.6253 36.1345 0.1214 31.8207 0.2235 34.2098 0.4503 37.1201 0.1268 32.0588 0.2193 34.3669 0.3981 36.9066
-TCM-HAT-F-Scratch 0.1570 30.0857 0.2976 32.5893 0.6211 36.1389 0.1214 31.9421 0.2235 34.2894 0.4503 37.1434 0.1258 32.0632 0.2189 34.3781 0.4001 36.9223
-TCM-HAT 0.1567 30.1843 0.2992 32.6454 0.6268 36.2267 0.1220 31.9737 0.2266 34.3319 0.4512 37.2486 0.1262 32.1423 0.2174 34.5124 0.3971 36.9934
-TCM-HAT-F 0.1574 30.4012 0.2998 32.8910 0.6276 36.4461 0.1229 32.1917 0.2249 34.4109 0.4512 37.3135 0.1255 32.4591 0.2186 34.7656 0.3975 37.3244
 5 Conclusion
@@
 Experiments demonstrate state-of-the-art performance in both SISR and Image Compression benchmarks, addressing limitations of previous methods.
-NeurIPS Paper Checklist
-The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact.
-Do not remove the checklist: The papers not including the checklist will be desk rejected.
-The checklist should follow the references and follow the (optional) supplemental material.
-The checklist does NOT count towards the page limit.
-Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
-[…]
-[…]
-    means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
-  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
-The checklist answers are an integral part of your paper submission.
-They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers.
-You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
-[…]
-IMPORTANT, please:
-  • Delete this instruction block, but keep the section heading “NeurIPS paper checklist",
-  • Keep the checklist subsection headings, questions/answers and guidelines below.
-  • Do not modify the questions and only use the provided macros for your answers.
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-[…]
-  4.
-Justification: Our abstract and introduction accurately reflect the paper’s contributions and scope by highlighting our novel contribution: we employ the factor fields for enhance the performance in super-resolution and multi-image compression.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-[…]
-  9. Justification: Our paper explicitly discusses the limitations of our approach, including the long training and inference time.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory Assumptions and Proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-[…]
-  14. Justification: Our paper primarily focuses on empirical research and does not introduce new theoretical results. As such, there are no theoretical assumptions or proofs discussed in the study.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental Result Reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-[…]
-  19.
-Justification: Our paper thoroughly details the methodologies, parameter settings, and experimental setups, providing all necessary information for replicating the main experimental results.
-This comprehensive documentation supports the validity of our primary claims and conclusions, ensuring clarity and reproducibility.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-[…]
-  24. Justification: We will make a full open access to both the code and the data used in our experiments.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental Setting/Details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-[…]
-  29.
-Justification: Our paper meticulously details the experimental setup, including data splits for training and testing, hyperparameter configurations, and the types of optimizers utilized.
-These elements are crucial for a thorough understanding of the experimental procedures and the interpretation of results.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment Statistical Significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-[…]
-  34.
-Justification: Due to the limited scope of our experiments, which encompass only two trials yielding similar results, our paper does not include detailed error bars or extensive statistical analysis.
-We recognize the potential for variability in these findings and suggest that additional experiments be conducted under varied conditions to validate and expand upon our results.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments Compute Resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-[…]
-  39. Justification: We use GPU GTX 4090 as our computational resources.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code Of Ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-[…]
-  44. Justification: We carefully follow the NeurIPS Code of Ethics.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader Impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-[…]
-  49. Justification: In the conclusion of our paper, we discuss the societal impacts directly related to our research on super resolution and image compression.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-[…]
-  54. Justification: Our paper does not include the release of pre-trained models, new datasets, or any other resources that typically necessitate specific safeguards to prevent misuse.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-[…]
-  59.
-Justification: We have diligently credited the creators and original owners of all assets used in our research, including code, data, and models, ensuring that each source is accurately cited within our paper.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New Assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-[…]
-  64. Justification:: Yes, any new assets introduced in the paper will be well documented.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and Research with Human Subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-[…]
-  69. Justification: Our research does not involve crowdsourcing experiments.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-[…]
-  74. Justification: Our research does not involve human subjects, and therefore, there were no study participants exposed to any risks.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+Limitations.
+While the FIPER framework offers advances in image compression and super-resolution, it requires future work to accommodate computation-limited scenarios such as real-time decoding.
+Also, Factorized Fields don’t consider semantic information, and we will leave this to future research for more possibilities.

### v3

@@
-FIPER: Generalizable Factorized Fields for Joint Image Compression and Super-Resolution
+FIPER: Generalizable Factorized Features for Robust Low-Level Vision Models
 Abstract
-In this work, we propose a unified representation for Super-Resolution (SR) and Image Compression, termed Factorized Fields, motivated by the shared principles between these two tasks.
-Both SISR and Image Compression require recovering and preserving fine image details—whether by enhancing resolution or reconstructing compressed data.
-Unlike previous methods that mainly focus on network architecture, our proposed approach utilizes a basis-coefficient decomposition to explicitly capture multi-scale visual features and structural components in images, addressing the core challenges of both tasks.
-We first derive our SR model, which includes a Coefficient Backbone and Basis Swin Transformer for generalizable Factorized Fields.
-Then, to further unify these two tasks, we leverage the strong information-recovery capabilities of the trained SR modules as priors in the compression pipeline, improving both compression efficiency and detail reconstruction.
-Additionally, we introduce a merged-basis compression branch that consolidates shared structures, further optimizing the compression process.
+In this work, we propose using a unified representation, termed Factorized Features, for low-level vision tasks, where we test on Single Image Super-Resolution (SISR) and Image Compression.
+Motivated by the shared principles between these tasks, they require recovering and preserving fine image details—whether by enhancing resolution for SISR or reconstructing compressed data for Image Compression.
+Unlike previous methods that mainly focus on network architecture, our proposed approach utilizes a basis-coefficient decomposition as well as an explicit formulation of frequencies to capture structural components and multi-scale visual features in images, which addresses the core challenges of both tasks.
+We replace the representation of prior models from simple feature maps with Factorized Features to validate the potential for broad generalizability.
+In addition, we further optimize the pipelines by leveraging the mergeable-basis property of our Factorized Features, which consolidates shared structures on multi-frame compression and super-resolution.
 Extensive experiments show that our unified representation delivers state-of-the-art performance, achieving an average relative improvement of 204.4% in PSNR over the baseline in Super-Resolution (SR) and 9.35% BD-rate reduction in Image Compression compared to the previous SOTA.
+Please see our project page at jayisaking.github.io/FIPER.
 1 Introduction
-Figure 1: The Proposed Factorized Fields.
-With explicit formulation of repetitive patterns and high frequency components, images can be represented with fine-grained details, especially useful for both Super-Resolution and Image Compression.
-Note that in this figure $\alpha=\{1,16\}$ and $\psi=sin$.
+Figure 1: The Proposed Factorized Features.
+With explicit formulation of repetitive patterns and high frequency components, images can be represented with fine-grained details, which is especially useful for low-level vision tasks such as Super-Resolution and Image Compression.
+Note that in this figure $\alpha=\{1,16\}$ and $\psi=\text{sin}(\cdot)$.
 Single Image Super-Resolution (SISR) aims to reconstruct high-quality images from low-resolution counterparts.
@@
 On the other hand, image compression serves as a fundamental task in low-level vision applications, where the traditional compression standards [35, 82, 79] lay the groundwork.
-The emerging learned image compression models [7, 2, 29, 48, 58, 17], compression algorithms of which mostly follow the pixel-space transform coding  [7, 27]paradigm, then introduce neural networks to further optimize compression efficiency by learning more compact latent representations and improving reconstruction quality.
+The emerging learned image compression models [8, 3, 30, 55, 67, 19], compression algorithms of which mostly follow the pixel-space transform coding  [8, 28] paradigm, then introduce a variety of networks  [83, 55] or difference losses [31] to further optimize compression efficiency by learning more compact latent representations and improving reconstruction quality.
 Specifically, they convert pixels into compact representations through a transform module, which eliminates the redundancy and reduces the bit cost in the subsequent entropy coding process.
 However, the core challenge of image compression is to accurately reconstruct the information lost during compression and quantization.
-In other words, the models are to recover the impaired components in images, which can essentially be viewed as reconstructing a high-quality image from its ‘low-resolution’ version, much like Super-Resolution.
-Based on the aforementioned analysis, it becomes clear that although Super-Resolution and Image Compression appear to be two distinct tasks, they share mutual similarities in two key aspects: (1) Both tasks require models to restore fine details from low-quality image content, as well as implicitly capture and reconstruct repetitive structural elements.
-(2) Both aim to conserve image quality, either by enhancing resolution or efficiently compressing data without significant loss of perceptual fidelity.
-Hence, inspired by recent advances in decomposition fields and matrices factorization in 3D scene modeling [9, 59, 8, 24, 6, 10, 26], we propose a unified representation, Factorized Fields, with generalizable Coefficient Backbone and Basis Transformer.
-This approach explicitly captures multi-scale visual features and repetitive structural components in images through a basis-coefficient decomposition.
+In other words, the models are to recover the impaired components in images, which can essentially be viewed as reconstructing a high-quality image from its ‘low-resolution’ version, much like Super-Resolution or all other common low-level vision tasks such as image restoration, deblurring, denoising, etc.
+Based on the aforementioned analysis, it becomes clear that although these tasks appear to be distinct, they share mutual similarities in two key aspects: (1) The tasks require models to restore fine details from low-quality image content, as well as implicitly capture and reconstruct repetitive structural elements.
+(2) They aim to conserve image quality, either by enhancing resolution or efficiently compressing data without significant loss of perceptual fidelity.
+Hence, inspired by recent advances in decomposition fields and matrices factorization in 3D scene modeling [11, 68, 9, 26, 7, 12, 27], we propose a unified representation, Factorized Features, with generalizable Coefficient Backbone and Basis Transformer for learned coefficient and basis, respectively.
+This approach explicitly captures multi-scale visual features and repetitive structral components in images through a basis-coefficient decomposition.
 The resulting representation strikes a balance between being compact and information-rich, enabling the resolution of structural ambiguities and the precise modeling of image details through a multi-frequency formulation.
-In the meantime, such a formulation imposes a factorization constraint during model training, which not only enhances the quality of single-image super-resolution (SISR) but also reduces distortion and improves compression efficiency by explicitly modeling structural elements.
-On top of these, to leverage the robust information-recovering capability of SR models, we integrate such priors with Image Compression models, whose knowledge of detail compensation further refines the lost key elements.
-Finally, we propose merging the bases by introducing an additional compression branch, consolidating multiple bases into one alongside multi-image transmission.
-This approach leverages the mutual information across multiple images, reducing the need for redundant transmissions and refining the basis structure.
+Finally, we propose merging the bases for multi-image processing by consolidating multiple bases into one through a network, as in traditional Discrete Fourier Transform, all signals of same size are reconstruct by the same set of basis functions, and we argue that learned shared basis can leverage the mutual information across multiple images to capture common structures.
 The main contributions of this paper are summarized as follows:
-  • We propose Generalized Factorized Fields, a unified representation that explicitly models multi-scale visual features and structural components for both super-resolution and image compression.
-  • We integrate super-resolution with image compression by introducing SR prior during decompression to compensate for lost details and developing a merged-basis compression branch for multi-image compression.
+  • We identify the weakness of previous similar representations and propose a refined representation, Factorized Features, based on thorough analysis as in Sec.
+3.1 for coarse and fine details of images via explicitly modeling multi-scale visual features and structural components.
+  • We generalize such image descriptor to learned settings on super-resolution and image compression, with potential applicability to broader low-level vision tasks.
   • We demonstrate state-of-the-art performance on benchmarks for both super-resolution and image compression through extensive experiments.
@@
 Super-Resolution (SR).
-Image super-resolution is critical in computer vision, focusing on recovering high-resolution (HR) images from low-resolution (LR) inputs.
-Following the foundational studies, CNN-based strategies [21, 46, 79, 64, 34] were initially introduced with modeling techniques such as residual learning [40, 47, 68, 78, 77, 18, 45] , or recursive learning [16, 65].
-Besides, subsequent research also sheds light on GAN-base methods [39, 40, 70, 71] to enhance realisticity and detail quality.
-However, the inductive bias of CNN-based networks by restricting spatial locality hinders the capture of long-range dependency from images, which is alleviated by Transformer-based SISR networks [12, 43].
-Afterward, SwinIR [44] is proposed to combine spatial locality and non-local information by Swin Transformer [50] with window attention and achieve breakthrough improvement in SISR.
-Following SwinIR’s success, several works have built upon its framework [19, 80, 75] to reach better image quality as well as solve information bottleneck [18].
-Hybrid approaches CRAFT [41] merge the benefits of convolutional and transformer structures to further elevate SR performance.
-For better feature aggregation, DAT [15] and HAT [13] integrate spatial and channel information using attention mechanisms to enhance their representation capabilities.
-Moreover, RGT [16] introduces a unique recursive-generalization self-attention mechanism that efficiently captures comprehensive spatial details with a linear increase in computational complexity.
+Image super-resolution (SISR) aims to reconstruct high-resolution (HR) images from low-resolution (LR) inputs, which is crucial for computer vision.
+Early CNN-based methods [23, 52, 95, 76, 36] introduced residual [41, 54, 81, 94, 93, 20, 51] and recursive learning [18, 77].
+GAN-based approaches [40, 41, 84, 85] improved perceptual quality but faced spatial locality constraints, leading to Transformer-based SISR models [14, 46].
+SwinIR [48] integrated window attention with Swin Transformer [58], inspiring further advancements [21, 96, 91, 20].
+Hybrid models like CRAFT [43], DAT [17], and HAT [15] optimize feature aggregation, while RGT [18] enhances spatial details efficiently.
 Image Compression (IC).
-Deep learning has significantly advanced image compression, offering superior compression ratios and image quality compared to traditional methods like JPEG [69] and JPEG-2000 [66].
-Early CNN-based approaches [7, 3, 28] have been surpassed by transformer-based models [37, 81, 72, 22], which leverage spatio-channel attention for better performance.
-GAN-based methods [53, 54, 23, 62] have further contributed to real-time adaptive compression.
-Recently, ELIC [30] introduced efficient compression with unevenly grouped space-channel contextual adaptive coding, while LIC TCM [48] integrated transformers and CNNs to capture both global and local image features.
-The eContextformer [38] introduced patch-wise, checkered, and channel-wise grouping techniques for parallel context modeling with a shifted window spatial-channel attention mechanism.
-GroupedMixer [42] proposed a transformer-based entropy model with token mixers for inner and cross-group context modeling.
-Meanwhile, Wavelet Conditional Diffusion [63] introduced a wavelet-based model with uncertainty-aware loss, balancing high perceptual quality with low distortion.
-In summary, while recent super-resolution and image compression advances focus on increasingly complex architectures, our work takes a different approach.
-We propose Factorized Fields, a unified framework that models visual and structural features, offering a more comprehensive solution to enhance performance in both tasks.
+Deep learning surpasses traditional compression methods like JPEG [82] and JPEG-2000 [79], with CNN-based [8, 4, 29], Transformer-based [38, 97, 86, 24], and GAN-based [61, 62, 25, 73] approaches improving performance.
+ELIC [32] introduced adaptive coding, LIC TCM [55] combined CNNs and Transformers, and eContextformer [39] leveraged spatial-channel attention.
+GroupedMixer [44] introduced token mixers, while Wavelet Conditional Diffusion [75] balanced perceptual quality and distortion.
+Our work diverges from these trends by introducing Factorized Features, a unified framework modeling both visual and structural features to enhance performance across super-resolution and image compression.
 3 Methods
-In this section, we present the details of our proposed methods. Inspired by factor fields, which can be formulated by
-$\hat{s}(x)=\mathcal{P}\Bigl{(}\textbf{Concat}_{i=1}^{N}\Bigl{\{}c_{i}(x)\odot b_{i}(\gamma_{i}(x))\Bigl{\}}\Bigl{)}.\end{aligned}$
-and whose details can be found in supplementary materials , we first explain the motivation and how we derive our Factorized Fields for enhanced image reconstruction quality in  Sec.
+In this section, we start by revisiting several previous renowned representations for images, e.g., 2D Fourier Transform, Learnable Fourier Series [57], and Factor Fields [12], and delve into the motivation and how we resolve the weaknesses of these formulations by our Factorized Features in Sec.
 3.1.
-We then describe how to adapt the formulation to Super-Resolution in  Sec.
-3.2.
-Finally, we show how to incorporate such representation with image(s) compression and how we integrate Super-Resolution and Image Compression in  Sec.
+Next, we discuss how to integrate such strategy into learned settings of Super-Resolution in Sec.
+3.2 and Image Compression Sec.
 3.3.
-Note that the background of learned image compression is described in supplementary materials.
+Note that the background of learned image compression is provided in supplementary materials.
 Figure 2: The correlation between coordinate transformation and downsampling.
@@
 This way, the feature sampled would capture the information in the original layout correctly; in other words, $I=\text{Sawtooth-Pixel-Sampling}(\text{Sawtooth-Aware-Downsample}(I))$
-Figure 3: The overall pipeline of image super-resolution with our Factorized Fields.
-Given a low-resolution image $I_{LR}$, we first extract coefficient feature map $X_{\text{coeff}}$ with the coefficient backbone, which is then decoded into coefficient and passed through the basis Swin Transformer for basis, separately.
+Figure 3: The overall pipeline of image super-resolution with our Factorized Features.
+Given a low-resolution image $I_{LR}$, we first extract the coefficient feature map $X_{\text{coeff}}$ with the coefficient backbone, which is then decoded into coefficients and passed through the basis Swin Transformer for basis, separately.
 Finally, as in Eq.
-4, the coefficient and basis are sampled, multiplied, and decoded for final high-resolution output $I_{HR}$, where $s$, $H$, $W$ denote the scale factor, height, and width respectively.
-3.1 Formulation of Factorized Fields
-As discussed in  Sec.
-1, the key to superior rendering quality in image regression tasks of both Image Compression and Super-Resolution lies in the capability of the representation to capture accurate structural distribution and fine visual details.
-Meanwhile, transformations such as Fourier Transform or Wavelet Transform have long been used to model multi-frequency information [25, 36] to express different implicit functions in images; however, such methods often suffer from under-expression due to the pre-defined and limited frequency bands or restriction from its formulation.
-Thus, we seek a representation that explicitly incorporates and fits multi-scale and multi-frequency components and yet is highly flexible and learnable according to individual features in an image.
-Inspired from recent success on such decomposition fields [24, 10, 9, 59, 6, 8], we primarily built our Factorized Fields framework on factor fields [10] from  Eq. 1:
-$\hat{I}(x)=\mathcal{P}\Bigl{(}\textbf{Concat}_{i=1}^{N}\Bigl{\{}c_{i}(x)\odot b_{i}(\gamma_{i}(x))\Bigl{\}}\Bigl{)}.\end{aligned}$
-Note that $\hat{I}$ denotes the approximated images, and $x\in R^{2}$ are the pixel coordinates.
-Such formulation has several key properties: First of all, by decomposing the images into basis frequencies, we can learn the implicit functions of an image and capture the mutual dependencies between pixels and across spatial composition; meanwhile, since the basis and coefficient are specific to every single image and both learnable in all spatial dimensions, the restriction on a limited number of basis (we use $N=6$ in all of our experiments) can be alleviated.
-A key insight of our work is the use of sawtooth coordinate transformation $\gamma(x)=x\bmod{k},k\in\mathbb{R}$, which performs best among tested transformations in image regression tasks.
-The sawtooth transformation maps input coordinates to a periodic space, implicitly capturing patch-like frequency information as shown in Fig.
-2a.
-However, traditional downsampling in neural networks fails to preserve this periodic structure.
-To address this, we introduce sawtooth-aware downsampling as in Fig.
-2c with our Basis Swin Transformer.
-This rearrangement allows the network to learn correlations between the periodic features in Sawtooth sampling.
-Moreover, in practice, we sample the basis via bilinear or bicubic sampling due to memory constraints, i.e., the feature size is less than image height and width, and this poses a severe problem: These nearby sampled pixel features are actually linear or based on cubic interpolation with respect to the basis, and such inductive bias hinders the representation of non-linearity in images.
-To resolve this, we propose a modified version, our Factorized Fields:
-$\hat{I}(x)=\mathcal{P}\Bigl{(}{\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl{\{}c_{i}(x)\odot\psi(\alpha_{j}\cdot b_{i}(\gamma_{i}(x)))\Bigl{\}}\Bigl{)},$ (3)
-where $\psi\in\{\text{sin},\text{cos}\},\alpha_{j}\in R$.
-Here, inspired by the Fourier Series, with a scalar $\alpha$ which corresponds to an increase in sampling frequency and a transformation function $\psi$ to add to implicit non-linearity functions and modulation of product, our complete Factorized Fields method stands out by the fact that such formulation significantly enhances the ability to represent complex non-linear structures in images and effectively composes high-frequency components between pixels.
-Refer to captionRefer to caption
-Figure 4: Comparison of Different Representation This figure illustrates the performance comparison on a single-image regression task using three different representations: vanilla factor fields in Eq.
-1, our Factorized Fields (Eq.
-3), and Plain Fields, which use simple channel-wise tensor concatenation instead of coefficient-basis multiplication.
-The left panel shows the MAE on the magnitude of different frequency components obtained from the Fourier Transform, while the right panel shows PSNR along training.
-Although the basis-coefficient decomposition method in [10] yields improvements, it still struggles with slow convergence and inconsistencies between low and high-frequency components.
-In contrast, by introducing explicit formulations for $\alpha$ and $\psi$, our method aligns different frequency components effectively and achieves superior convergence speed.
-This advantage becomes especially crucial when the task extends to generalizable applications, such as super-resolution or image compression.
-Our Factorized Fields enhance the original Factor Fields [10] in three key ways: (1) non-linear transformation through $\psi\in\{\text{sin},\text{cos}\}$ enables better modeling of complex image patterns, (2) multi-frequency modulation using $\alpha_{j}$ allows explicit control over frequency components, and (3) sawtooth-aware downsampling preserves periodic structure during feature processing.
+8, the coefficient and basis are sampled, multiplied, and decoded for final high-resolution output $I_{HR}$, where $s$, $H$, and $W$ denote the scale factor, height, and width, respectively.
+3.1 Formulation of Factorized Features
+Fourier Transform has been long used in many scenarios to represent periodic or finite signals with a group of sinusoidal functions.
+In practice, it often takes the form of Discrete Cosine Transform for real-valued 2D discrete signal $f$ with height $H$ and width $W$:
+$f[x,y]=\sum_{u=0}^{H-1}\sum_{v=0}^{W-1}S_{u,v}\cos\left[\frac{\pi u}{H}\left(x+\frac{1}{2}\right)\right]\cos\left[\frac{\pi v}{W}\left(y+\frac{1}{2}\right)\right]$ (1)
+where $S_{u,v}$ is the amplitude.
+Without the loss of generality, we consider square images in this work, with edge length (and also the period) denoted as $T$:
+$f[x,y]=\sum_{u=0}^{T-1}\sum_{v=0}^{T-1}S_{u,v}\cos\left[\frac{2\pi}{T}(ux+vy)-\xi_{u,v}\right]$ (2)
+where $x<T$, $y<T$, and $\xi_{u,v}$ is phase shift.
+In image processing tasks, the goal is to learn $S$ and $\xi$, then use the formula above to reconstruct high-quality images from degraded inputs.
+To achieve this, [57] treats $S$ and $\xi$ as learnable parameters, allowing the model to optimize them for better reconstruction.
+Furthermore, the frequencies $u$ and $v$ are also designed to be promptable by the model, enabling dynamic adaptation to spatial-varying high- and low-frequency components.
+More generally, we can formulate it as the multiplication of coefficient $c$ and basis $b$ with simplified notation, and in this work, we use $\mathbf{x}$ to represent the pixel coordinates $(x,y)$:
+$f(\mathbf{x})=\sum_{i=1}^{N}c_{i}\,\cdot\,b_{i}(\mathbf{x}),$ (3)
+where $b_{i}(\mathbf{x})=\cos\left[(u_{i}x+v_{i}y)-\xi_{i}\right]$ and $N$ denotes the number of frequency components.
+Spatially Variant Coefficient.
+Intuitively, the local frequency spectrum of a signal is spatially variant, meaning that different regions exhibit distinct spectral characteristics, e.g., certain regions of an image do not require high-frequency components.
+Therefore, in this work, we argue that coefficient $c_{i}$ should be spatial-varying, which is different from previous works, i.e.,
+$f(\mathbf{x})=\sum_{i=1}^{N}c_{i}(\mathbf{x})\,\cdot\,b_{i}(\mathbf{x}).$ (4)
+Learned Non-uniform Basis.
+Previous works learn basis through frequencies ${u,v}$ and phase shift $\xi$[57, 47, 42]; however, constructing signals with definitive sets of uniform sinusoidal harmonic functions does not facilitate the utmost details; in other words, since low-frequency components typically require more energy [45], during the learning process with training losses of MAE or MSE, models would attend more to low-frequency parts and sacrifice local details, and the predefined nature of these sinusoidal functions further restricts local flexibility, making fine-grained adjustments more challenging.
 As shown in Fig.
-4, our refined representation contributes to large margin of improvement in convergence and performance, demonstrating the modification is the key to superior information modeling.
+5, higher frequency components converge slower than their counterpart.
+Hence, in this work, we generate the entire basis map $b_{i}\in R^{T\times T}$ for better fitting and local reconstruction.
+To this end, we have introduced our rough idea: Under the generalizable setting, we use networks to generate a multi-channel coefficient map and a multi-channel basis map and combine them together.
+In this work, the composition process happens in the end with a projection function $\mathcal{P}$ to form RGB images $\hat{I}$:
+$\hat{I}(\mathbf{x})=\mathcal{P}\Bigl{(}\textbf{Concat}_{i=1}^{N}\Bigl{\{}c_{i}(\mathbf{x})\cdot b_{i}(\mathbf{x})\Bigl{\}}\Bigl{)}.$ (5)
+In contrast to previous works which employ fourier bottlenecks[57], or frequency decomposition blocks[47, 42, 45] among model blocks, here we seek to address a fundamental question: Can a well-learned representation effectively replace the simple feature map output ?
+That is, our solution is meant to be applied directly on networks to represent any images.
+Refer to caption Refer to caption Refer to caption Refer to caption
+(a) Eq. 4 (b) Eq. 6 (c) Eq. 7 with only $\psi$ (d) Eq. 7 with both $\psi$ and $\alpha$
+Figure 4: Visualization of different factorization methods.
+Coordinate Transformation.
+The success of Fourier-based representation originates from its sinusoidal decomposition.
+By using such periodic functions, one can effectively analyze and reconstruct signals with the enforcement of high- and low-frequency components.
+However, the formulation in Eq.
+5 does not have such explicit harmonic feature (coefficient and basis are learned maps), and thus as in Fig.
+4 we witness performance loss due to lack of such guidance.
+Inspired by position encoding [65, 11, 68, 7, 9] and Fourier feature mapping [78], we introduce coordinate transformation function $\gamma_{i}$ on $\mathbf{x}$ before sampling basis $b_{i}$:
+$\hat{I}(x)=\mathcal{P}\Bigl{(}\textbf{Concat}_{i=1}^{N}\Bigl{\{}c_{i}(x)\cdot b_{i}(\gamma_{i}(x))\Bigl{\}}\Bigl{)}.\end{aligned}$
+In practice, we empirically[12] use sawtooth transformation $\gamma(x)=x\bmod{k},k\in\mathbb{R}$, visualized in Fig.
+2.a.
+We argue that the function explicitly imposes patch-like periodic constraints and, therefore, enforces models to learn such repetitive patterns.
+Multi-frequency Modulation.
+[72] has shown that neural networks are biased towards learning low-frequency content.
+Hence, we propose multi-frequency modulation to compel the model to fit high-frequency components, visualized in Fig.
+4:
+$\hat{I}(x)=\mathcal{P}\Bigl{(}{\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl{\{}c_{ij}(x)\odot\psi(\alpha_{j}\cdot b_{i}(\gamma_{i}(x)))\Bigl{\}}\Bigl{)},$ (7)
+where $\psi\in{\sin,\cos}$ and $\alpha\in R$.
+In this way, each basis $b_{i}$ contributes to both high- and low-frequency components by different values of $\alpha$.
+Consequently, if a basis is only accurate in the low-frequency domain, the output will contain undesired high-frequency noise.
+Alternatively, Eq.
+7 can be interpreted as making the spatial coordinates $x,y$ in Eq.
+1 learnable instead of the frequency components $u,v$.
+To validate this, we conduct experiments (Fig.
+5) showing that learning coordinates mitigates the rigid constraint of uniform distribution, aligning better with the non-uniform nature of visual patterns across spatial layouts; that is, some pixels share similar RGB values, indicating smooth color transitions, while others exhibit abrupt changes, suggesting high-frequency variations.
+To this end, our full Factorized Features representation is formulated in Eq. 7.
+Figure 5: In this figure, we validate the claim that fixed sinusoidal functions attend more to low-frequency parts.
+Learnable basis denotes the Learned Non-uniform basis in Sec.
+3.1.
+And we can see that Learnable Coordinate $x$ has the same effect, verifying our conclusion.
 Figure 6: The illustration of our joint image-compression and super-resolution framework compared with the traditional compression-only method.
@@
 3.2 as information-recovery prior, detailed in Sec.
-3.3.1.
-(c) Expanding on (b),  Sec.
+3.3 (c) Expanding on (b),  Sec.
 3.3.1 introduces a multi-image compression strategy that utilizes both our SR Module and a Basis Merging Transformer to capture shared structure.
-3.2 Super-Resolution with Factorized Fields
-We represent a super-resolved image using our Factorized Fields, where coefficients and basis are generated by networks $F_{\text{coeff}}$ and $F_{\text{basis}}$ from a low-resolution Image:
-$\hat{I}_{\text{SR}}(x)=\mathcal{P}\Bigl{(}{\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl{\{}c_{i}^{\text{LR}}(x)\odot\psi(\alpha_{j}\cdot b_{i}^{\text{LR}}(\gamma_{i}(x)))\Bigl{\}}\Bigl{)},$ (4)
-where $c^{\text{LR}}_{i}(x)=\text{Conv}(F_{\textbf{coeff}}(I_{\text{LR}}))_{i}(x)$ and $b^{\text{LR}}_{i}(x)=F_{\textbf{basis}}(F_{\textbf{coeff}}(I_{\text{LR}}))_{i}(\gamma_{i}(x))$.
+3.2 Super-Resolution with Factorized Features
+We represent a super-resolved image using our Factorized Features, where coefficients and basis are generated by networks $F_{\text{coeff}}$ and $F_{\text{basis}}$ from a low-resolution Image:
+$\hat{I}_{\text{SR}}(x)=\mathcal{P}\Bigl{(}{\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl{\{}c_{ij}^{\text{LR}}(x)\odot\psi(\alpha_{j}\cdot b_{i}^{\text{LR}}(\gamma_{i}(x)))\Bigl{\}}\Bigl{)},$ (8)
+where $c^{\text{LR}}(x)=\text{Conv}(F_{\textbf{coeff}}(I_{\text{LR}}))(x)$ and $b^{\text{LR}}(x)=F_{\textbf{basis}}(F_{\textbf{coeff}}(I_{\text{LR}}))(\gamma(x))$.
 Note that we sample the outputs $\text{Conv}(F_{\textbf{coeff}}(I_{\text{LR}}))$ and $F_{\textbf{basis}}(F_{\textbf{coeff}}(I_{\text{LR}}))$ with coordinates $x$ and $\gamma(x)$, respectively.
-Our model comprises three main components: Coefficient Backbone, Basis Swin Transformer, and Factorized Fields Reconstruction.
+Our model comprises three main components: Coefficient Backbone $F_{\textbf{coeff}}$, Basis Swin Transformer $F_{\textbf{basis}}$, and Factorized Features Reconstruction.
 As shown in  Fig.
 3, the process begins with $I_{LR}\in\mathbb{R}^{3\times H\times W}$.
-The Coefficient Backbone extracts features $X_{\text{coeff}}\in\mathbb{R}^{C_{c}\times H_{c}\times W_{c}}$, which are then used to generate coefficients $c$ through convolution and pixel shuffle operations, and fed into the Basis Swin Transformer to produce a multi-scale basis $b=\{b_{1},...,b_{N}\},\ b_{i}\in\mathbb{R}^{C_{b_{i}}\times H_{b_{i}}\times W_{b_{i}}}$.
+The Coefficient Backbone extracts features $X_{\text{coeff}}\in\mathbb{R}^{C_{c}\times H_{c}\times W_{c}}$, which are then used to generate coefficients $c$ through convolution and pixel shuffle operations.
+Then, $X_{\text{coeff}}$ is fed into the Basis Swin Transformer to produce a multi-scale basis $b=\{b_{1},...,b_{N}\},\ b_{i}\in\mathbb{R}^{C_{b_{i}}\times H_{b_{i}}\times W_{b_{i}}}$.
 The coefficients and basis are combined to reconstruct $I_{SR}\in\mathbb{R}^{3\times sH\times sW}$ using  Eq.
-4, where $s$ is the scale factor.
-In our experiments, $\gamma_{i}(x)$ is a Sawtooth function: $\gamma_{i}(x)=x\mod k_{i}$, with $k_{i}\in R$.
-We optimize model parameters using an $L_{1}$ loss function.
-To demonstrate our method’s effectiveness, we use existing SR methods [14, 15, 75] as the Coefficient Backbone.
-For the Basis Swin Transformer, we employ Swin Transformer Blocks [51] with a series of downsampling operations.
-We use a dilation-like downsampling technique( Fig.
-2c) to accommodate the sawtooth sampling pattern.
+8, where $s$ is the scale factor.
+We optimize model parameters using $L_{1}$ loss.
+To demonstrate the effectiveness of our method, we use existing SR methods [16, 17, 91] as the Coefficient Backbone.
+For the Basis Swin Transformer, we employ simple Swin Transformer Blocks [59] with a series of special downsampling operations.
+Specifically, we use a dilation-like downsampling technique( Fig.
+2.c) to accommodate the sawtooth sampling pattern.
 The final basis is refined using additional upsampling and convolution layers.
-3.3 Image Compression with Factorized Fields and Super-Resolution
+3.3 Image Compression with Factorized Features
 Image Compression, at its core, is to strike a balance between the amount of information contained in the latent bits and the final image quality.
 However, as discussed in  Sec.
-1,  Sec.
-2, and Sec.
-3.1, most recent works draw emphasis on how to retrieve implicit elements through designing different architectures, such as analysis transforms and entropy models [38, 42], or decompression modules [63, 22], while this paper aims at addressing the representation itself to better capture the structural correlations and thus achieve better image quality through explicit modeling of frequency components and Factorized Fields formulation as in  Eq.
-3.
+1 and  Sec.
+2, most recent works draw emphasis on how to retrieve implicit elements through designing different architectures, such as analysis transforms and entropy models [39, 44], or decompression modules [75, 24], while this paper aims at addressing the representation itself to better capture the structural correlations and thus achieve better image quality through Factorized Features formulation as in  Eq.
+7
+Priors from SR.
 In addition, with our trained SR model described in  Sec.
@@
 Thus, since Super-Resolution and Image Compression share the core principle of reconstructing and enhancing image details from low-quality sources, we can effectively integrate this prior into the compression pipeline.
-In the following, we respectively present the proposed joint framework for the single and multi-image compression tasks.
-3.3.1 Single Image Compression
 The overall pipeline is shown in  Fig.
-5b.
+6b.
 To demonstrate the robustness of our representation and the effectiveness of the SR prior, the compression and decompression networks greatly follow [55], with only the synthesis transform replaced by our SR pipeline, where the details can be referenced in Supplementary Materials.
@@
 Since compression induces error, according to  Fig.
-5b, the Coefficient feature map $X_{\text{coeff}}$ generated with Coefficient Backbone contains misinformation, and Basis Swin Transformer that uses such features would capture the information wrongfully for the basis and thus amplify error.
+6b, the Coefficient feature map $X_{\text{coeff}}$ generated with Coefficient Backbone contains misinformation, and Basis Swin Transformer that uses such features would capture the information wrongfully for the basis and thus amplify error.
 To solve this, we exploit the mergeable property of basis, compress and transmit the merged basis independently along other quantized variables $\{Q(y,\mu),\ \hat{z}\}$, and finally reconstruct the $M$ images with the merged basis and their individual decoded coefficients.
 as illustrated in Fig.
-5c.
+6c.
 This way, we can enjoy less information loss with basis while maintaining low bit rates.
@@
 Experimental Setup.
-We conduct extensive experiments to validate the effectiveness of our Factorized Fields representation for Super-Resolution tasks.
+We conduct extensive experiments to validate the effectiveness of our Factorized Features representation for Super-Resolution tasks.
 Following the strategy outlined in [15] and [20], we adapt the same-task pretraining approach for all the Super-Resolution models.
@@
 After pretraining, we use DF2K (DIV2K [1] + Flickr2K [52]) as the finetuning dataset following [21, 15] for 200k iterations.
-For pretraining, we utilize AdamW optimizer with learning rate 1e-4, batch size 16, betas (0.9, 0.99), and other parameters set to PyTorch default, while we use learning rate 1e-5 during finetuning.
+For pretraining, we utilize AdamW optimizer with learning rate 2e-4, batch size 32, betas (0.9, 0.99), and other parameters set to PyTorch default, while we use learning rate 1e-5 during finetuning.
 Throughout training, the input is randomly cropped to $256\times 256$ and bicubicly resized to $64\times 64$ for Coefficient Backbone input.
-As for the hyperparameter of Factorized Fields, the number $N$ of coefficient and basis is set to 6, and the scalars $\alpha_{j}$ are set to {1, 4, 16, 64} since we want to capture both base frequency $\alpha_{j}=1$ and high frequency $\alpha_{j}=64$ information.
+As for the hyperparameter of Factorized Features, the number $N$ of coefficient and basis is set to 6, and the scalars $\alpha_{j}$ are set to {1, 4, 16, 64} since we want to capture both base frequency $\alpha_{j}=1$ and high frequency $\alpha_{j}=64$ information.
 More implementation details can be found in supplementary materials.
@@
 To validate the effectiveness of our framework, we employ three different SoTA SR models—ATD [91], DAT [17], and HAT [16]—as the Coefficient Backbone in our pipeline, which we denote as ATD-F, DAT-F, and HAT-F, respectively.
-These models exhibit significant improvements when compared to their counterparts ATD-L, DAT-L, and HAT-L, which possess similar parameter counts.
+These models exhibit significant improvements when compared to their counterparts, ATD-L, DAT-L, and HAT-L, which possess similar parameter counts.
 It is important to note that only [16] provides a large-scale model.
@@
 Specifically, we argue that in our pipeline, the Coefficient Backbone functions more as a feature extraction module, where the refined features facilitate downstream basis extraction.
-Lastly, to evaluate the effectiveness of our Factorized Fields, we trained a model named HAT-F-Concat, which does not apply the formulation in Eq.
-3.
+Lastly, to evaluate the effectiveness of our Factorized Features, we trained a model named HAT-F-Concat, which does not apply the formulation in Eq.
+7.
 Instead, it concatenates the basis and coefficient directly and decodes the resulting features to produce the output.
@@
 We provide the visual comparison in Fig.
-6.
+7.
 The images are randomly sampled from the DIV2K dataset.
 Our method faithfully reconstructs the image details, whereas the other approaches suffer from over-smoothing or hallucinating details absent in the ground truth.
-Multi-frame SR.
-Due to the space limit, we provide additional comparisons on multi-frame SR in the supplementary materials.
+Multi-frame SR. Due to space limitations, we provide additional comparisons on multi-frame SR in the supplementary materials.
 (a) Kodak
@@
 Figure 8: Performance (RD-Curve) evaluation on image compression using different datasets.
+Table 2: Comprehensive evaluation for image compression. Using VTM as an anchor for calculating BD-Rate. Latencies are measured under an NVIDIA GTX 3090 GPU.
 Method BD-Rate (%) $\downarrow$ Latency(s) Params(M)
@@
 VTM 0.00 129.21 0.14 -
-Xie (MM 21’) -0.78 2.93 6.00 50.0
-Cheng (CVPR 22’) 5.44 1.98 4.69 29.6
-STF (CVPR 22’) -4.31 0.14 0.13 99.9
-ELIC (CVPR 22’) -7.24 0.07 0.09 36.9
-TCM (CVPR 23’) -11.74 0.16 0.15 76.7
+Xie (MM’21) -0.78 2.93 6.00 50.0
+Cheng (CVPR’22) 5.44 1.98 4.69 29.6
+STF (CVPR’22) -4.31 0.14 0.13 99.9
+ELIC (CVPR’22) -7.24 0.07 0.09 36.9
+TCM (CVPR’23) -11.74 0.16 0.15 76.7
+CCA (NeurIps’24) -20.97 0.11 0.09 64.9
 TCM-HAT-L-F (Ours) -21.09 0.109 0.264 110.34
@@
 TCM-HAT-F-multi M=24 (Ours) -20.97 0.232 0.174 131.35
-Table 2: Comprehensive evaluation for image compression. Using VTM as an anchor for calculating BD-Rate. Latencies are measured under an NVIDIA GTX 3090 GPU.
+Table 3: Validation of the effectiveness of SR prior. The best PSNRs are marked in red.
 Method Kodak CLIC Tecnick
@@
 TCM-HAT-F 0.1574 30.4012 0.2998 32.8910 0.6276 36.4461 0.1229 32.1917 0.2249 34.4109 0.4512 37.3135 0.1255 32.4591 0.2186 34.7656 0.3975 37.3244
-Table 3: Validation of the effectiveness of SR prior. The best PSNRs are marked in red.
 4.2 Single- and Multi-Image Compressions
 Experimental Setup.
-We evaluate our Factorized Fields representation for image compression tasks, comparing it against state-of-the-art methods.
-Following our Super-Resolution setting in Experiment Setup from  Sec.
-4.1, we use the same set of SA-1B for training.
-To emphasize our representation, we initialize compression and decompression modules in Fig.
-5 from pre-trained [48] and the SR Module from those in Sec.
-4.1, we then train the pipeline end to end on 256x256 patches for 200k iterations, with AdamW optimizer [52] with learning rate 1e-5, batch size 16, betas (0.9, 0.99) and other parameters set to PyTorch default.
+We evaluate our Factorized Features for image compression tasks against state-of-the-art methods.
+Following the Super-Resolution setup in Sec.
+4.1, we train on the SA-1B dataset.
+To emphasize our representation, compression and decompression modules are initialized from pre-trained models [55], then fine-tuned end-to-end on 256×256 patches for 200k iterations using the AdamW optimizer [60] (learning rate=1e-5, batch size=16, betas=(0.9, 0.99)).
 We integrate our SR Module with the pre-trained compression module TCM, creating TCM-HAT-F and TCM-HAT-L-F models.
@@
 For multi-image compression in Fig.
-5.c, we set the Basis Swin Transformer and the Basis Merging Transformer to be trainable while the other parts remain frozen.
+6.c, we set the Basis Swin Transformer and the Basis Merging Transformer to be trainable while the other parts remain frozen.
 Rate-Distortion Performance Comparison.
@@
 The classical image compression codec, VVC [80], is also tested by using VTM12.1.The rate-distortion performance on various datasets, including Kodak, Tecnick’s old test set with resolution 1200$\times$1200, and CLIC Professional Validation, is shown in Fig.
-7.
+8.
 In  Tab.
@@
 The result shows that direct transmission of bases would reduce the error from Coefficient Backbone to Basis Swin Transformer and that the distortion increases with $M$, as the information contained in the merged basis is limited, and merging multiple bases into one would cause increasing information loss.
-Our analysis reveals several significant advantages of the FIPER framework in image compression tasks.
-The approach demonstrates substantial improvements in compression efficiency across various benchmark datasets, including Kodak, CLIC, and Tecnick, indicating its broad applicability.
-Notably, the multi-image compression strategy shows particularly promising results for larger image sets, suggesting scalability benefits.
-Furthermore, our method maintains competitive latency while significantly improving compression performance and balancing efficiency and quality.
+Our experiments validate the effectiveness of FIPER in image compression, demonstrating substantial improvements across benchmarks (Kodak, CLIC, Tecnick).
+The multi-image compression strategy particularly excels with larger image sets, highlighting strong scalability.
+Additionally, our method achieves competitive latency, significantly improving compression efficiency while balancing quality and speed.
+Table 4: Comparison of improvements of Factorized Features. $\psi$ and $\alpha$ are the same in Eq. 8.
 Metric PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
@@
 No $\alpha$ for pixel-wise frequency information 21.25 0.537 0.527
-Table 4: Comparison of improvements of Factorized Fields. $\psi$ and $\alpha$ are the same in Eq. 4.
 4.3 Ablation Studies
 Influence of SR Priors in Image Compression.
-We conduct experiments with various configurations to verify the proposed image compression pipeline’s effectiveness.
-As shown in Tab.
-3, we present the quantitative performance of models trained with different values of $\lambda$ in Eq.
-5.
-Specifically, TCM-HAT refers to substituting our SR Module with the original HAT [14] in the pipeline illustrated in Fig.
-5b.
-TCM-HAT-F represents our complete pipeline, while TCM-HAT-F-Scratch denotes the same pipeline but with the SR Module initialized randomly.
-Our results demonstrate that integrating SR priors with image compression improves performance, and our proposed representation further enhances results.
-This highlights the robustness of our Factorized Fields in capturing fine details in image regression tasks.
-Effectiveness of Factorized Fields Design.
-We conduct experiments to verify our modification of Factorized Fields, modified from Eq.
-1 to Eq.
-3.
+We conduct experiments with various configurations to verify our image compression pipeline’s effectiveness.
+TCM-HAT refers to using the original HAT [16] instead of our SR Module in Fig.
+6b.
+TCM-HAT-F represents our full pipeline, while TCM-HAT-F-Scratch initializes the SR Module randomly.
+Results demonstrate that integrating SR priors improves image compression performance, and our representation further enhances results, highlighting the robustness of Factorized Fields.
+Effectiveness of Factorized Features Design.
+We conduct experiments to verify our multi-frequency modulation.
 The quantitative performance reported on single-image regression is shown in Tab.
@@
 5 Conclusion
-We proposed Factorized Fields, a representation that models implicit structures and patterns by decomposing images into multi-frequency components.
-This approach addresses challenges in Super-Resolution and Image Compression by restoring details and preserving visual fidelity.
-We integrate SR priors with Image Compression for improved information recovery and introduce a basis merging technique for enhanced rendering quality across multiple images.
-Experiments demonstrate state-of-the-art performance in both SISR and Image Compression benchmarks, addressing limitations of previous methods.
+We proposed Factorized Features, a representation that decomposes images into multi-frequency components to model implicit structures and patterns.
+Our approach addresses challenges in Super-Resolution and Image Compression by effectively restoring details and preserving visual fidelity.
+We integrate SR priors with Image Compression for better information recovery and introduce a basis-merging technique to enhance multi-image processing performance.
 Limitations.
-While the FIPER framework offers advances in image compression and super-resolution, it requires future work to accommodate computation-limited scenarios such as real-time decoding.
-Also, Factorized Fields don’t consider semantic information, and we will leave this to future research for more possibilities.
+Although effective, our method requires further optimization for computation-limited scenarios such as real-time decoding.
+Additionally, incorporating semantic information remains a promising direction for future research.
+[…]
+Supplementary Material
+Appendix A Overview
+This supplementary material provides additional implementation details and experimental results to complement our main manuscript.
+Specifically, we first elaborate on the background of factor fields and learned image compression in Appendix B, which were omitted in the main paper due to space limitations.
+Next, in Appendix D, we present comprehensive details about the benchmarks, including the Video Super-Resolution task, which was not discussed in the main paper.
+Finally, we provide extensive visualizations to facilitate a more detailed comparison of the results in Appendix H.
+Appendix B Preliminaries
+In this section, we briefly introduce the background of the factor fields [12], and Learned Image Compression along with their properties.
+Factor Fields.
+The concept of decomposition fields or factorized matrices in the reconstruction of 2D images or 3D scenes has shown superior rendering quality and improved efficiency recently [11, 27, 68, 26, 7, 9], while [12] first proposed a unified representation called factor fields.
+For a 1D signal $s(x)$, the factor fields formulation is:
+$\hat{s}(x)=c(x)^{T}b(\gamma(x)),$ (11)
+where $c(x)=(c_{1}(x),...,c_{N}(x))^{T}$ are spatial-varying coefficient fields, $b(x)=(b_{1}(x),...,b_{N}(x))^{T}$ are basis functions, and $\gamma(x)$ is a coordinate transformation.
+The coefficient and basis are queried by sampling with the coordinates $x$.
+This generalizes to Q-dimensional signals:
+$\hat{s}(x)=\mathcal{P}(c(x)\odot b(\gamma(x))),$ (12)
+where $\mathcal{P}$ is a projection function and $\odot$ denotes the element-wise product. In practice, for multi-dimensional coefficients:
+$\hat{s}(x)=\mathcal{P}\Bigl{(}\textbf{Concat}_{i=1}^{N}\Bigl{\{}c_{i}(x)\odot b_{i}(\gamma_{i}(x))\Bigl{\}}\Bigl{)}.\end{aligned}$
+This formulation allows the same basis to be applied at multiple spatial locations while varying the coefficients, especially when $\gamma$ is a periodic function.
+Learned Image Compression.
+Following [66, 3], a learned image compression model with a channel-wise entropy model can be formulated as:
+$z$ $=h_{a}(y;\phi_{h}),\ y=g_{a}(x;\phi),$ (14)
+$\{X_{\text{mean}},X_{\text{scale}}\}$ $=h_{s}(\hat{z};\theta_{h}),\ \hat{z}=Q(z),$
+$\hat{y}$ =$=\{Q(y_{0}-\mu_{0})+\mu_{0},...,Q(y_{t}-\mu_{t})+\mu_{t}\},$
+$\hat{x}$ $=g_{s}(\overline{y};\theta),\ \overline{y}=\textbf{Refine}_{\theta_{r}}({\mu_{0},...,\mu_{t}},\hat{y}),$
+where $0<=t<l,\ \mu_{t}=e_{i}(\hat{y}_{<i},X_{\text{mean}})$.
+The encoder $g_{a}$ transforms the raw image $x$ into a latent representation $y$.
+A hyper-prior encoder $h_{a}$ further processes $y$ to output $z$, capturing spatial dependencies.
+$z$ is then quantized to $\hat{z}$, where a factorized density model $\varphi$ is used to encode quantized $\hat{z}$ as $p_{\hat{z}\mid\varphi}(\hat{z}\mid\varphi)=\prod_{j}\left(p_{z_{j}\mid\varphi}(\varphi)*\mathcal{U}\left(-\frac{1}{2},\frac{1}{2}\right)\right)(\hat{z}_{j})$ where $j$ specifies the position or each element of each signal.
+The derivation of the equation can be found in [2].
+Next, $\hat{z}$ is decoded by $h_{s}$ to produce features $X_{\text{mean}}$ and $X_{\text{scale}}$, used to estimate the mean $\mu$ and variance $\sigma$ of $y$.
+The latent $y$ is divided into $l$ slices, and each quantized around computed means $\mu_{t}$.
+These means are derived from earlier quantized slices and $X_{\text{mean}}$ by a slice network $e_{i}$.
+The quantized slices form $\hat{y}$.
+For decompression, $\hat{y}$ is refined using $\textbf{Refine}_{\theta_{r}}$ based on $\mu_{t}$ and $\hat{y}$ to produce $\overline{y}$, approximating the original $y$.
+Finally, $g_{s}$ reconstructs the decompressed image $\hat{x}$ from $\overline{y}$.
+The model is trained using a Lagrangian multiplier-based rate-distortion optimization:
+$L$ $=R(\hat{y})+R(\hat{z})+\lambda\cdot D(x,\hat{x})$ (15)
+$=\mathbb{E}\left[-\log_{2}\left(p_{\hat{{y}}|\hat{{z}}}\left(\hat{{y}}\mid\hat{{z}}\right)\right)\right]+\mathbb{E}\left[-\log_{2}\left(p_{\hat{{z}}|\varphi}\left(\hat{{z}}\mid\varphi\right)\right)\right]$
+$+\lambda\cdot\mathcal{D}\left({x},\hat{{x}}\right)$
+where $R(\hat{y})$ and $R(\hat{z})$ denote bit rates, $D(x,\hat{x})$ is the distortion term (calculated by MSE), and $\lambda$ balances compression efficiency and image fidelity.
+In our experiments, we follow [55], modifying only $g_{s}$ to demonstrate our representation’s effectiveness.
+Appendix C Differences from Factor Fields.
+We explicitly highlight how sawtooth-based sampling and multi-frequency modulation extend the original factor fields.
+In our 1D and 2D toy examples as shown in Figs.
+9 and 10, traditional SR/IC representations can be viewed as a “vanilla” scenario with $N=1$ and $c(x)=1$.
+Factor fields already improve accuracy over such baselines, but our introduction of $\psi$ and $\alpha$ further captures both high- and low-frequency details.
+Specifically, forcing $\alpha=4$ compels the network to learn broader frequency ranges and avoid noise from the fixed modulation.
+To confirm these advantages, we trained smaller SR models and compared them with factor fields in Tab.
+5.
+The results illustrate that our multi-frequency constraint yields significant gains even in simpler configurations, underscoring its importance for high-frequency detail preservation and overall reconstruction quality.
+Refer to caption Refer to caption Refer to caption Refer to caption
+(a) Vanilla fields (b) Factor fields [10] (c) Ours with $\psi$ only (d) Ours with both $\psi$ and $\alpha$
+Figure 9: 1D toy example.
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+(a) Convergence curve (b) Target image (c) Vanilla fields (d) Factor fields (e) Ours with $\psi$ only (f) Ours with $\psi$ and $\alpha$
+Figure 10: 2D toy example.
+Table 5: Comparison with Factor Fields.
+Method Sawtooth $\psi$ $\alpha$ Params MACs Forward Pass Set5 Set14 B100 Urban100 Manga109
+[…]
+HAT-S 9.62 40.38 3,214 32.92 0.9047 29.15 0.7958 27.97 0.7505 27.87 0.8346 32.35 0.9283
+HAT 20.77 86.02 3,692 33.04 0.9056 29.23 0.7973 28.00 0.7517 27.97 0.8368 32.48 0.9292
+HAT-FA-Small 15.99 57.61 3,678 33.06 0.9065 29.26 0.7976 28.04 0.7524 28.12 0.8413 32.52 0.9301
+HAT-F-Small-only-sin $\{\text{sin}\}$$\{1,4,16,64\}$ 16.13 66.34 4,176 33.16 0.9067 29.33 0.7989 28.05 0.7530 28.26 0.8429 32.64 0.9312
+HAT-F-Small-half-freq $\{\text{sin},\text{cos}\}$$\{1,16\}$ 16.13 66.34 4,148 33.18 0.9073 29.35 0.8003 28.02 0.7521 28.31 0.8431 32.70 0.9316
+HAT-F-Small $\{\text{sin},\text{cos}\}$$\{1,4,16,64\}$ 16.26 75.06 4,640 33.24 0.9079 29.39 0.8009 28.06 0.7543 28.41 0.8453 32.82 0.9324
+HAT-F (Our FIPER) $\{\text{sin},\text{cos}\}$$\{1,4,16,64\}$ 45.97 158.79 5,750 33.31 0.9085 29.45 0.8011 28.09 0.7553 28.55 0.8463 32.99 0.9330
+HAT-L (ImageNet pre-trained) 40.84 167.27 6,804 33.30 0.9083 29.47 0.8015 28.09 0.7551 28.60 0.8498 33.09 0.9335
+Appendix D Implementation Details
+Single Image Super-Resolution
+Figure 11: Detailed architecture of Super-Resolution Modules
+The architecture of our SR networks is shown in Fig.
+11.
+Firstly, for Coefficient Backbones from various pre-trained models like [15, 91, 17], we replace the output upsampling layer (mostly pixel shuffle) with convolution of channel $D=256$, i.e., for input image with resolution $64\times 64$, the original pre-trained models process hidden states in $64\times 64$, and then upsample the output to $256\times 256$ for final result, where we replace the upsampling with convolution and output the final hidden state $X_{coeff}\in R^{256\times 64\times 64}$.
+Next, the Basis Swin Transformer begins with feature embedding, extending the channels $D$ from $256$ to $384$, where the features are then passed through a series of Swinv2[59] and Sawtooth-Aware Downsampling Blocks.
+Specifically, each block contains two layers, one of which uses shifted window attention.
+After each block, the hidden state goes through upsampling by 4x and convolution to convert the output basis channel to $24$.
+Then, the Sawtooth-Aware Downsample block rearranges the hidden state $h$ as discussed in the main paper, which can be viewed as:
+$h=\mathbf{rearrange}(h,\text{'b c (s h) (s w) $\rightarrow$ b (s s c) h w'}),\ s=2$ (16)
+, and reduces the feature channel by 4. In all experiments we use $N=6$, i.e. there are 6 bases, where $Basis_{i}\in R^{24\times\frac{256}{2^{i}}\times\frac{256}{2^{i}}}\text{for i in }\{0\ ...\ 5\}$.
+Finally, we derive our Coefficient by a series of Convolution and Pixelshuffle layers and decoded the Bases and Coefficient to a predicted high-resolution image by Eq. 17:
+$\hat{I}(x)=\mathcal{P}\Bigl{(}{\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl{\{}c_{i}(x)\odot\psi(\alpha_{j}\cdot b_{i}(\gamma_{i}(x)))\Bigl{\}}\Bigl{)}.$ (17)
+In practice, the coefficient is generated with channels equal to the total number of base channels, and it is split accordingly by channels.
+Method Frames REDS4 Vimeo-90K-T Vid4
+REDS/Vimeo PSNR SSIM PSNR SSIM PSNR SSIM
+BasicVSR++[10] 30/14 32.39 0.9069 37.79 0.9500 27.79 0.8400
+VRT[50] 16/7 32.19 0.9006 38.20 0.9530 27.93 0.8425
+RVRT[49] 30/14 32.75 0.9113 38.12 0.9527 27.99 0.8462
+PSRT[74] 16/14 32.72 0.9106 38.27 0.9536 28.07 0.8485
+IA-RT[88] 16/7 32.90 0.9138 38.14 0.9528 28.26 0.8517
+Ours-HAT-F 16/7 32.67 0.9096 38.09 0.9512 27.95 0.8441
+Table 6: Performance comparison among different methods on REDS4, Vimeo-90K-T, and Vid4 datasets.
+Note that we don’t use any special design on the network, such as temporal correspondence or motion sampling like previous works, demonstrating our effectiveness.
+Multi-Image (Video) Super-Resolution
+Building upon our prior work in Single-Image Super-Resolution (SISR), we extend the applicability of Factorized Features to tackle the challenges in Multi-Image Super-Resolution (MISR).
+In the main paper, we introduced the Basis Merging Transformer for Multi-Image Compression.
+Here, we further expand its functionality to generate per-timestamp Coefficients and Bases.
+To evaluate the effectiveness of our approach, we use Video Super-Resolution (VSR) as the benchmark task.
+Specifically, we simply leverage transformer blocks from [70] to model temporal correlations.
+For a given video, we first derive the Coefficient and Basis for each frame and rearrange each level $i$ of the Coefficients or Basis such that the spatial dimensions $H$ and $W$ are treated as the batch dimension, while the temporal dimension is processed as a sequence of tokens for the Transformer.
+Next, we apply positional embedding to the tokens, process the tokens with the transformer blocks, and finally decode the Coefficient and Basis token from the respective position with convolution layers for the output super-resolved frames.
+Tab.
+6 shows quantitative comparison with state-of-the-art methods.
+All VSR experiments were conducted using bicubic 4X downsampling.
+The training dataset includes the REDS[69] and Vimeo-90K[89] datasets, while the testing dataset comprices REDS4[69], Vid4[53], and Vimeo-90K-T[89].
+For the REDS dataset, we train for 300k iterations using 16 input frames, with a learning rate of 1e-4 and a cosine learning rate that gradually decays to 1e-7.
+The batch size is set to 8, and the Coefficient Backbone and the Basis Swin Transformer are initialized from SISR.
+When training on the Vimeo-90K dataset, we first conduct 300k iterations with 14 input frames with flip sequence, we then train model on 7 input frames with flip sequence, using a learning rate of 1e-4 and a cosine learning rate decay to 1e-7.
+We initialize the weights using the model trained on the REDS dataset.
+The batch size remains 8.
+Note that the training is only conducted on the newly added transformer blocks, i.e., the Coefficient Backbone and Basis Swin Transformer are frozen.
+Test results for the REDS model are reported on the REDS4 dataset, while test results for the Vimeo-90K model are reported on Vimeo-90K-T and Vid4.
+We calculate PSNR and SSIM on the RGB channel for REDS4 and Y channel for Vimeo-90K-T and Vid4, following previous work[88, 74, 49].
+Single Image Compression
+Figure 12: Detailed architecture of Compression Modules
+As discussed in the main paper, to make a fair comparison with the state-of-the-art, our entropy model and compression backbone greatly follow that of [55], as shown in Fig.
+12(more details can be bound in [55].
+What we do here is that we use the output from the last main layer and convert it to match the hidden channel of the SR Module.
+Multi-Image Compression
+Following VSR, we use transformer blocks to merge the bases.
+The difference here is that we additionally prepend a CLS token so that we can use it as the final merged basis.
+For the merged basis compression, we utilize the same architecture of the image branch, i.e., we randomly initialize a module from [55] with input and out channels change.
+After the Basis Merging Transformer is trained, we follow the same training setting for Basis compression while randomly sampling 1 to 24 images for a batch and setting only the Basis compression module trainable.
+Appendix E Merged-Basis Strategy & SR Priors.
+The SR prior accelerates detail recovery by leveraging high-resolution knowledge while basis merging exploits shared structures for efficiency.
+Since merging is introduced after, the compression model is pre-trained, and improvements stem not only from the SR prior but also from the well-initialized basis merging.
+When trained from scratch (see Tab.
+7), the merging transformer tends to find a universal solution rather than image-specific structures.
+Additionally, merging inevitably causes information loss, prompting the coefficient branch to encode more image-specific details.
+This mutual offset between SR prior and basis merging ultimately strengthens our approach.
+TCM-HAT-F-multi $M=24$ TCM-HAT-F-multi $M=24$ Coeff-Frozen TCM-HAT-F-multi $M=24$ Scratch
+BD-rate % -20.97 -19.18 -18.34
+Table 7: Comparison of Merged-Basis Strategies.
+Appendix F Encoding/Decoding Times.
+Fig.
+5(c) processes each of the $M$ images through the Coefficient Backbone, Basis Swin Transformer, and Basis Merging Transformer, then compresses and transmits the merged basis.
+In Tab.
+2, different $M$ values primarily affect the run time of the Basis Merging Transformer.
+Although merging larger $M$ increases computation, per-image overhead remains small once averaged by $M$, as shown in Tab.
+8.
+Notably, Tab.
+2 assumes each image’s basis is compressed independently, which can be confusing.
+In practice, a merged basis is processed only once, so averaged encoding/decoding time is more realistic.
+$M=1$ $M=2$ $M=4$ $M=8$ $M=16$ $M=24$
+Total Basis Merging Time (sec) 0.00328 0.00628 0.01140 0.02312 0.04784 0.08064
+Averaged Basis Merging Time (sec) by $M$ 0.00328 0.00314 0.00285 0.00289 0.00299 0.00336
+Averaged Merged Basis Compression Time (sec) by $M$ 0.0645 0.0323 0.0161 0.0081 0.0040 0.0027
+Averaged Merged Basis Decompression Time (sec) by $M$ 0.0935 0.0467 0.0234 0.0117 0.0058 0.0039
+Adjusted Enc Time (sec) by $M$ 0.2320 0.1998 0.1836 0.1756 0.1715 0.1702
+Adjusted Dec Time (sec) by $M$ 0.1740 0.1272 0.1039 0.0922 0.0863 0.0844
+[…]
+Table 8: Comparison of Encoding/Decoding Times.
+Appendix G Inference Time Comparison for SR.
+We provide the inference time in Tab. 9, compared with extensive SR models. The inference time of different components of FIPER is also provided.
+HAT-S HAT HAT-L ATD DAT RGT HAT-F-S (Ours) HAT-F(Ours) HAT-L-F(Ours) Coefficient Backbone (HAT-F) Basis Swin Transformer (HAT-F) Factorized Features Reconstruction (HAT-F)
+Inference Time (s) 0.0387 0.0398 0.0778 0.0592 0.0594 0.0612 0.0671 0.0716 0.1084 0.0400 0.0132 0.0184
+Table 9: Inference time comparison for SR.
+Appendix H More Visualization
+H.1 Single Image Super-Resolution
+Below, we provide more visual comparisons of single-image super-resolution, where FIPER denotes HAT-F.
+[Uncaptioned image][Uncaptioned image]
+H.2 Image Compression
+Below, we provide more visual comparisons of image compression. The results are present with TCM-HAT-F and the metrics are PSNR/Bpp.
+[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
+[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
+[Uncaptioned image][Uncaptioned image][Uncaptioned image]

### v4

@@
-FIPER: Generalizable Factorized Features for Robust Low-Level Vision Models
+FIPER: Factorized Features for Robust Image Super-Resolution and Compression
 Abstract
 In this work, we propose using a unified representation, termed Factorized Features, for low-level vision tasks, where we test on Single Image Super-Resolution (SISR) and Image Compression.
-Motivated by the shared principles between these tasks, they require recovering and preserving fine image details—whether by enhancing resolution for SISR or reconstructing compressed data for Image Compression.
+Motivated by the shared principles between these tasks, they require recovering and preserving fine image details, whether by enhancing resolution for SISR or reconstructing compressed data for Image Compression.
 Unlike previous methods that mainly focus on network architecture, our proposed approach utilizes a basis-coefficient decomposition as well as an explicit formulation of frequencies to capture structural components and multi-scale visual features in images, which addresses the core challenges of both tasks.
 We replace the representation of prior models from simple feature maps with Factorized Features to validate the potential for broad generalizability.
-In addition, we further optimize the pipelines by leveraging the mergeable-basis property of our Factorized Features, which consolidates shared structures on multi-frame compression and super-resolution.
+In addition, we further optimize the compression pipeline by leveraging the mergeable-basis property of our Factorized Features, which consolidates shared structures on multi-frame compression.
 Extensive experiments show that our unified representation delivers state-of-the-art performance, achieving an average relative improvement of 204.4% in PSNR over the baseline in Super-Resolution (SR) and 9.35% BD-rate reduction in Image Compression compared to the previous SOTA.
-Please see our project page at jayisaking.github.io/FIPER.
+Project page: https://jayisaking.github.io/FIPER/
 1 Introduction
-Figure 1: The Proposed Factorized Features.
-With explicit formulation of repetitive patterns and high frequency components, images can be represented with fine-grained details, which is especially useful for low-level vision tasks such as Super-Resolution and Image Compression.
-Note that in this figure $\alpha=\{1,16\}$ and $\psi=\text{sin}(\cdot)$.
-Single Image Super-Resolution (SISR) aims to reconstruct high-quality images from low-resolution counterparts.
-Specifically, the key lies in accurately restoring fine details and reconstructing the correct arrangement of visual features.
-Thus, geometric correspondences or repetitive patterns, such as stripes, grids, or textures, are commonly used for evaluation due to their rich details that are crucial to image fidelity.
-Early CNN-based approaches [52, 23] laid the foundation for SISR, which was later enhanced with GAN-based methods [40] for improved perceptual realism.
-Follow-up Transformer-based networks [14] address non-local dependencies, with subsequent Swin Transformer-based approaches sparking tremendous advancements [21, 17, 18], which then inspire more and more complex, delicate designs of heavy network architectures in SISR.
-However, these prior works have primarily focused on network architecture design rather than addressing the capability of representation.
-The visual patterns and the inherent nature of image content structure that play an influential role in SISR have not been explicitly considered in the representation learning process.
+Figure 1: In this work, we propose to represent images by Factorized Features, contrasting with prior methods that either (a) rely on fixed-basis decomposition[66, 39] or (b) introduce various complex model blocks [15, 71] such as sparse attention, dense skip connection, etc.
+As highlighted in the orange boxes, (a) successfully captures periodic line structures but suffers from lower fidelity, while (b) cannot recover those repeating patterns without explicit frequency modeling.
+In contrast, our method (c) uses learned decomposition with generalizable bases to deliver high-fidelity reconstruction while preserving periodic structures.
+We also present LAM [32] and Diffusion Index (DI) [32] for reconstructing the red-boxed patch, with its pixel-importance map shown on the middle row; a higher DI indicates more pixel utilization.
+Single-image super-resolution (SISR) aims to recover high-quality images from low-resolution inputs, with accuracy depending on precise restoration of fine details and geometric correspondences (e.g., stripes, grids, textures).
+CNN pioneers [24, 59] were followed by GAN-based methods for perceptual realism [47] and Transformer networks for long-range context [14].
+Swin-based variants further boosted performance [22, 18, 19, 71], inspiring ever-heavier architectures.
+These prior works optimize network design while neglecting the underlying image-content representations themselves.
+On the other hand, decomposition-based methods[66, 44, 39] seek to model the frequency components directly through adding Fourier- or Wavelet-like bottlenecks in the middle.
+Although these designs can indeed better capture the recurring visual patterns in learning effective SR representations than aforementationed methods, as illustrated in fig.˜1, they still suffer from poor perceptual quality.
 This raises a critical question: beyond a simple network output, can we derive a formulation that more effectively captures these patterns and aligns with the goals of SISR?
+Conventional multi-scale pyramids aggregate features at several resolutions, but they store all activations densely, do not disentangle frequency content, and cannot be shared directly across tasks.
+Instead, our factorized descriptor decomposes images into generalizable basis $\times$ coefficient pairs, providing an explicit multi-frequency handle through different modulations.
 On the other hand, image compression serves as a fundamental task in low-level vision applications, where the traditional compression standards [41, 95, 92] lay the groundwork.
-The emerging learned image compression models [8, 3, 30, 55, 67, 19], compression algorithms of which mostly follow the pixel-space transform coding  [8, 28] paradigm, then introduce a variety of networks  [83, 55] or difference losses [31] to further optimize compression efficiency by learning more compact latent representations and improving reconstruction quality.
+The emerging learned image compression models [8, 3, 34, 62, 80, 20], compression algorithms of which mostly follow the pixel-space transform coding  [8, 31] paradigm, then introduce a variety of networks  [96, 62, 106, 67, 27] or difference losses [35] to further optimize compression efficiency by learning more compact latent representations and improving reconstruction quality.
 Specifically, they convert pixels into compact representations through a transform module, which eliminates the redundancy and reduces the bit cost in the subsequent entropy coding process.
 However, the core challenge of image compression is to accurately reconstruct the information lost during compression and quantization.
-In other words, the models are to recover the impaired components in images, which can essentially be viewed as reconstructing a high-quality image from its ‘low-resolution’ version, much like Super-Resolution or all other common low-level vision tasks such as image restoration, deblurring, denoising, etc.
+In other words, the models can essentially be viewed as reconstructing a high-quality image from its ‘low-resolution’ version, much like Super-Resolution or all other common low-level vision tasks such as image restoration, deblurring, denoising, etc.
 Based on the aforementioned analysis, it becomes clear that although these tasks appear to be distinct, they share mutual similarities in two key aspects: (1) The tasks require models to restore fine details from low-quality image content, as well as implicitly capture and reconstruct repetitive structural elements.
@@
 Hence, inspired by recent advances in decomposition fields and matrices factorization in 3D scene modeling [11, 81, 9, 28, 7, 12, 30], we propose a unified representation, Factorized Features, with generalizable Coefficient Backbone and Basis Transformer for learned coefficient and basis, respectively.
-This approach explicitly captures multi-scale visual features and repetitive structral components in images through a basis-coefficient decomposition.
+This approach explicitly captures multi-scale visual features and repetitive structural components in images through a basis-coefficient decomposition.
 The resulting representation strikes a balance between being compact and information-rich, enabling the resolution of structural ambiguities and the precise modeling of image details through a multi-frequency formulation.
-Finally, we propose merging the bases for multi-image processing by consolidating multiple bases into one through a network, as in traditional Discrete Fourier Transform, all signals of same size are reconstruct by the same set of basis functions, and we argue that learned shared basis can leverage the mutual information across multiple images to capture common structures.
+Finally, we also propose consolidating multiple bases into one through a network for multi-image processing, as in traditional Discrete Fourier Transform, all signals of same size are reconstruct by the same set of basis functions, and we argue that learned shared basis can leverage the mutual information across multiple images to capture common structures.
 The main contributions of this paper are summarized as follows:
-  • We identify the weakness of previous similar representations and propose a refined representation, Factorized Features, based on thorough analysis as in Sec.
-3.1 for coarse and fine details of images via explicitly modeling multi-scale visual features and structural components.
-  • We generalize such image descriptor to learned settings on super-resolution and image compression, with potential applicability to broader low-level vision tasks.
+  • We identify the weakness of previous similar representations and propose a refined representation, Factorized Features, based on thorough analysis as in section˜4.1 for coarse and fine details of images via explicitly modeling multi-scale visual features and structural components.
+  • We generalize such image descriptors to learned settings on super-resolution and image compression, with potential applicability to broader low-level vision tasks.
   • We demonstrate state-of-the-art performance on benchmarks for both super-resolution and image compression through extensive experiments.
@@
 Early CNN-based methods [24, 59, 113, 89, 43] introduced residual [48, 61, 94, 112, 111, 21, 58] and recursive learning [19, 90].
-GAN-based approaches [40, 41, 84, 85] improved perceptual quality but faced spatial locality constraints, leading to Transformer-based SISR models [14, 46].
-SwinIR [48] integrated window attention with Swin Transformer [58], inspiring further advancements [21, 96, 91, 20].
-Hybrid models like CRAFT [43], DAT [17], and HAT [15] optimize feature aggregation, while RGT [18] enhances spatial details efficiently.
+GAN-based approaches [47, 48, 97, 98] improved perceptual quality but faced spatial-locality constraints, leading to Transformer-based SISR models [14, 53].
+SwinIR [55] integrated window attention with Swin Transformer [69], inspiring further advancements [22, 114, 109, 21, 71].
+Hybrid models such as CRAFT [50], DAT [18], and HAT [15] optimize feature aggregation, while RGT [19] enhances spatial details efficiently.
+Recent efficient Transformer variants such as Restormer [105] and Uformer [99] demonstrate that lightweight attention or activation-free blocks can match or surpass heavier designs.
+Diffusion-based methods such as DDRM [42] provide an orthogonal line of plug-and-play restoration priors, with recent advances extending to video restoration [104], reference-based face restoration [37], and inpainting tasks [64].
+Additionally, learned pipeline approaches [68] reverse the image formation process for reconstruction tasks.
+Spectral Coordinate Transforms.
+A separate line of work seeks to encode coordinates rather than deepen the backbone.
+Fourier Features [91, 78] map input positions to a fixed harmonic basis, enabling MLPs to model high-frequency detail.
+CoordConv [65] expands convolutional layers with raw $(x,y)$ channels to facilitate spatial reasoning.
+On the other hand, sawtooth coordinate transformation $\gamma(\cdot)$, of our Factorized Features, folds space into a piece-wise linear triangle wave and produces evenly spaced spectral peaks (see section˜4.1) that favor sparse and low-rank Factorized Features.
+Unlike the above encoding, our basis-coefficient decomposition paired with $\gamma$ can be trained end-to-end and reused directly in both super-resolution and compression pipelines.
 Image Compression (IC).
-Deep learning surpasses traditional compression methods like JPEG [82] and JPEG-2000 [79], with CNN-based [8, 4, 29], Transformer-based [38, 97, 86, 24], and GAN-based [61, 62, 25, 73] approaches improving performance.
-ELIC [32] introduced adaptive coding, LIC TCM [55] combined CNNs and Transformers, and eContextformer [39] leveraged spatial-channel attention.
-GroupedMixer [44] introduced token mixers, while Wavelet Conditional Diffusion [75] balanced perceptual quality and distortion.
+Deep learning surpasses traditional codecs such as JPEG [95] and JPEG-2000 [92], with CNN-based [8, 4, 33], Transformer-based [45, 115, 100, 25, 27, 106, 67, 29], and GAN-based [73, 74, 26, 86] approaches improving performance.
+ELIC [36] introduced adaptive coding; LIC TCM [62] combines CNNs and Transformers; eContextformer [46] leverages spatial–channel attention.
+GroupedMixer [51] introduces token mixers, while Wavelet Conditional Diffusion [88] balances perceptual quality and distortion.
+Generative codecs such as HiFiC [77] pursue high-fidelity reconstructions at extremely low bitrates and offer a complementary direction to likelihood-based models.
 Our work diverges from these trends by introducing Factorized Features, a unified framework modeling both visual and structural features to enhance performance across super-resolution and image compression.
-3 Methods
-In this section, we start by revisiting several previous renowned representations for images, e.g., 2D Fourier Transform, Learnable Fourier Series [57], and Factor Fields [12], and delve into the motivation and how we resolve the weaknesses of these formulations by our Factorized Features in Sec.
-3.1.
-Next, we discuss how to integrate such strategy into learned settings of Super-Resolution in Sec.
-3.2 and Image Compression Sec.
-3.3.
-Note that the background of learned image compression is provided in supplementary materials.
-Figure 2: The correlation between coordinate transformation and downsampling.
-(a) The sawtooth transformation example with $k=2$.
-(b) The PixelUnShuffle downsample.
-(c) To explicitly model the information for sampling with a sawtooth, we rearrange the feature map in a dilation-like manner in the downsample layer of the Basis Swin Transformer.
-This way, the feature sampled would capture the information in the original layout correctly; in other words, $I=\text{Sawtooth-Pixel-Sampling}(\text{Sawtooth-Aware-Downsample}(I))$
-Figure 3: The overall pipeline of image super-resolution with our Factorized Features.
-Given a low-resolution image $I_{LR}$, we first extract the coefficient feature map $X_{\text{coeff}}$ with the coefficient backbone, which is then decoded into coefficients and passed through the basis Swin Transformer for basis, separately.
-Finally, as in Eq.
-8, the coefficient and basis are sampled, multiplied, and decoded for final high-resolution output $I_{HR}$, where $s$, $H$, and $W$ denote the scale factor, height, and width, respectively.
-3.1 Formulation of Factorized Features
-Fourier Transform has been long used in many scenarios to represent periodic or finite signals with a group of sinusoidal functions.
-In practice, it often takes the form of Discrete Cosine Transform for real-valued 2D discrete signal $f$ with height $H$ and width $W$:
-$f[x,y]=\sum_{u=0}^{H-1}\sum_{v=0}^{W-1}S_{u,v}\cos\left[\frac{\pi u}{H}\left(x+\frac{1}{2}\right)\right]\cos\left[\frac{\pi v}{W}\left(y+\frac{1}{2}\right)\right]$ (1)
+3 Frequency Decomposition Preliminary
+In this section, we start by revisiting several previous renowned representations for images, e.g., 2D Fourier Transform, Learnable Fourier Series [66], and Factor Fields [12], and then generalize them to a concise formulation of basis-coefficient decomposition as in eq.˜3.
+The Fourier Transform has long been used in many scenarios to represent periodic or finite signals with a group of sinusoidal functions.
+In practice, it often takes the form of the Discrete Cosine Transform for a real-valued 2D discrete signal $f$ with height $H$ and width $W$:
+$\small f[x,y]=\sum_{u=0}^{H-1}\sum_{v=0}^{W-1}S_{u,v}\cos\left[\frac{\pi u}{H}\left(x+\frac{1}{2}\right)\right]\cos\left[\frac{\pi v}{W}\left(y+\frac{1}{2}\right)\right],$ (1)
 where $S_{u,v}$ is the amplitude.
 Without the loss of generality, we consider square images in this work, with edge length (and also the period) denoted as $T$:
-$f[x,y]=\sum_{u=0}^{T-1}\sum_{v=0}^{T-1}S_{u,v}\cos\left[\frac{2\pi}{T}(ux+vy)-\xi_{u,v}\right]$ (2)
+$\small f[x,y]=\sum_{u=0}^{T-1}\sum_{v=0}^{T-1}S_{u,v}\cos\left[\frac{2\pi}{T}(ux+vy)-\xi_{u,v}\right],$ (2)
 where $x<T$, $y<T$, and $\xi_{u,v}$ is phase shift.
@@
 More generally, we can formulate it as the multiplication of coefficient $c$ and basis $b$ with simplified notation, and in this work, we use $\mathbf{x}$ to represent the pixel coordinates $(x,y)$:
-$f(\mathbf{x})=\sum_{i=1}^{N}c_{i}\,\cdot\,b_{i}(\mathbf{x}),$ (3)
-where $b_{i}(\mathbf{x})=\cos\left[(u_{i}x+v_{i}y)-\xi_{i}\right]$ and $N$ denotes the number of frequency components.
+$\small f(\mathbf{x})=\sum_{i=1}^{N}c_{i}\,\cdot\,b_{i}(\mathbf{x}),\ b_{i}(\mathbf{x})=\cos\left[(u_{i}x+v_{i}y)-\xi_{i}\right],$ (3)
+where $N$ denotes the number of frequency components.
+In the next section, starting from the concise formulation in eq.˜3, we introduce our Factorized Features framework and elaborate on its spatially variant coefficients and adaptive basis functions, while showing its application to super-resolution and image compression.
+4 Methods
+In this section, we dive into the motivation and how we resolve the weaknesses of previous decomposition-based methods by our Factorized Features in section˜4.1.
+Next, we discuss how to integrate such a strategy into learned settings of Super-Resolution in section˜4.2 and Image Compression section˜4.3.
+We provide the background of learned image compression in the supplementary materials.
+4.1 Formulation of Factorized Features
+In this subsection, we focus on the analysis of the properties from different factorization components, and derive our final Factorized Features through these insights, starting from the simplest expression eq.˜3.
 Spatially Variant Coefficient.
 Intuitively, the local frequency spectrum of a signal is spatially variant, meaning that different regions exhibit distinct spectral characteristics, e.g., certain regions of an image do not require high-frequency components.
-Therefore, in this work, we argue that coefficient $c_{i}$ should be spatial-varying, which is different from previous works, i.e.,
-$f(\mathbf{x})=\sum_{i=1}^{N}c_{i}(\mathbf{x})\,\cdot\,b_{i}(\mathbf{x}).$ (4)
+Therefore, in this work, we set coefficient $c_{i}$ to be spatial-varying, following [12], i.e.,
+$\small f(\mathbf{x})=\sum_{i=1}^{N}c_{i}(\mathbf{x})\,\cdot\,b_{i}(\mathbf{x}).$ (4)
 Learned Non-uniform Basis.
-Previous works learn basis through frequencies ${u,v}$ and phase shift $\xi$[57, 47, 42]; however, constructing signals with definitive sets of uniform sinusoidal harmonic functions does not facilitate the utmost details; in other words, since low-frequency components typically require more energy [45], during the learning process with training losses of MAE or MSE, models would attend more to low-frequency parts and sacrifice local details, and the predefined nature of these sinusoidal functions further restricts local flexibility, making fine-grained adjustments more challenging.
-As shown in Fig.
-5, higher frequency components converge slower than their counterpart.
-Hence, in this work, we generate the entire basis map $b_{i}\in R^{T\times T}$ for better fitting and local reconstruction.
-To this end, we have introduced our rough idea: Under the generalizable setting, we use networks to generate a multi-channel coefficient map and a multi-channel basis map and combine them together.
-In this work, the composition process happens in the end with a projection function $\mathcal{P}$ to form RGB images $\hat{I}$:
-$\hat{I}(\mathbf{x})=\mathcal{P}\Bigl{(}\textbf{Concat}_{i=1}^{N}\Bigl{\{}c_{i}(\mathbf{x})\cdot b_{i}(\mathbf{x})\Bigl{\}}\Bigl{)}.$ (5)
+Previous works[66, 54, 49] learn basis through frequencies u,v{u,v} and phase shift $\xi$; however, constructing signals with definitive sets of uniform sinusoidal harmonic functions does not facilitate the utmost details.
+The predefined nature of these sinusoidal functions restricts local flexibility, making fine-grained adjustments more challenging.
+Also, since low-frequency components typically require more energy [52], during the learning process with training losses of MAE or MSE, models would attend more to low-frequency parts and sacrifice local details.
+As shown in fig.˜2(a), higher frequency components converge after their counterpart.
+Hence, in this work, we generate the entire basis map $b_{i}=M_{i}\in R^{T\times T}$ for better fitting and local reconstruction.
+(a) $b_{i}(\mathbf{x})=\cos\left[({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}u_{i}}x+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}v_{i}}y)-{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\xi_{i}}\right]$
+(b) $b_{i}(\mathbf{x})=\cos\left[(u_{i}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x}+v_{i}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}y})-\xi_{i}\right]$
+(c) $b_{i}=\cos\left(\alpha_{1}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}M_{i}}\right)+\cos\left(\alpha_{2}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}M_{i}}\right)$
+Figure 2: Comparison of different trainable basis settings We test on single-image regression with $f(\mathbf{x})=\sum_{i=1}^{N}c_{i}(\mathbf{x})\,b_{i}(\mathbf{x})$ with equal parameter count.
+(a) fixed sinusoidal basis favoring low–frequency content, eq.˜3; (b) learnable (i.e.
+requires gradient) coordinate mapping with moderate gains; (c) Learned Non-uniform Basis $M_{i}\in\mathbb{R}^{T\times T}$ converges more synchronously.
+Note that $\alpha_{1},\alpha_{2}$ are the frequencies, functionally similar to $u,v$.
+To this end, we have introduced our rough idea: Under the generalizable setting, we use networks to generate a multi-channel coefficient map and a multi-channel basis map, and combine them.
+In this work, the composition process happens in the end with a projection function $\mathcal{P}$ for RGB images $\hat{I}$:
+$\small\hat{I}(\mathbf{x})=\mathcal{P}\Bigl(\textbf{Concat}_{i=1}^{N}\Bigl\{c_{i}(\mathbf{x})\cdot b_{i}(\mathbf{x})\Bigl\}\Bigl).$ (5)
 In contrast to previous works which employ Fourier bottlenecks[66], or frequency decomposition blocks[54, 49, 52] among model blocks, here we seek to address a fundamental question: Can a well-learned representation effectively replace the simple feature map output?
 That is, our solution is meant to be applied directly on networks to represent any images.
-Refer to caption Refer to caption Refer to caption Refer to caption
-(a) Eq. 4 (b) Eq. 6 (c) Eq. 7 with only $\psi$ (d) Eq. 7 with both $\psi$ and $\alpha$
-Figure 4: Visualization of different factorization methods.
+Next, we discuss more improvements on frequency decomposition to model image details.
+Refer to caption Refer to caption
+(a) eq.˜5 (b) eq.˜6
+Refer to caption Refer to caption
+(c) eq.˜7 with only $\psi$ (d) eq.˜7 with both $\psi$ and $\alpha$
+Figure 3: Visualization of different factorization methods This figure shows the difference between vanilla fields, coordinate transformation, and multi-frequency modulation.
+Specifically, (a) is a vanilla basis-coefficient field, and (b) adds sawtooth transformation.
+We can see in (b) that the coordinate transformation explicitly models a patch-like pattern, e.g., the second plot from the top is divided into two periods.
+By enforcing such frequency components, the models can decompose the signal effectively.
+Next, (d) is our full formulation, while (c) has no $\alpha$.
+Intuitively, from (b) to (c), applying $\psi$ (we use $\cos$ here) should not have much effect on the performance since we do not have any constraint on these learnable bases.
+With the introduction of $\alpha$, the models are forced to attend to signal components of different frequencies.
 Coordinate Transformation.
-The success of Fourier-based representation originates from its sinusoidal decomposition.
+The success of the Fourier-based representation originates from its sinusoidal decomposition.
 By using such periodic functions, one can effectively analyze and reconstruct signals with the enforcement of high- and low-frequency components.
-However, the formulation in Eq.
-5 does not have such explicit harmonic feature (coefficient and basis are learned maps), and thus as in Fig.
-4 we witness performance loss due to lack of such guidance.
+However, the formulation in eq.˜5 does not have such explicit harmonic feature (coefficient and basis are learned maps), and thus as in fig.˜3((a) vs (b)) we witness performance loss due to lack of such guidance.
 Inspired by position encoding [78, 11, 81, 7, 9] and Fourier feature mapping [91], we introduce coordinate transformation function $\gamma_{i}$ on $\mathbf{x}$ before sampling basis $b_{i}$:
-$\hat{I}(x)=\mathcal{P}\Bigl{(}\textbf{Concat}_{i=1}^{N}\Bigl{\{}c_{i}(x)\cdot b_{i}(\gamma_{i}(x))\Bigl{\}}\Bigl{)}.\end{aligned}$
-In practice, we empirically[12] use sawtooth transformation $\gamma(x)=x\bmod{k},k\in\mathbb{R}$, visualized in Fig.
-2.a.
+$\small\hat{I}(x)=\mathcal{P}\Bigl(\textbf{Concat}_{i=1}^{N}\Bigl\{c_{i}(x)\cdot b_{i}(\gamma_{i}(x))\Bigl\}\Bigl).$ (6)
+In practice, we empirically[12] use sawtooth transformation $\gamma(x)=x\bmod{k},k\in\mathbb{R}$; visualization is also provided in the supplementary material.
 We argue that the function explicitly imposes patch-like periodic constraints and, therefore, enforces models to learn such repetitive patterns.
@@
 [85] has shown that neural networks are biased towards learning low-frequency content.
-Hence, we propose multi-frequency modulation to compel the model to fit high-frequency components, visualized in Fig.
-4:
-$\hat{I}(x)=\mathcal{P}\Bigl{(}{\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl{\{}c_{ij}(x)\odot\psi(\alpha_{j}\cdot b_{i}(\gamma_{i}(x)))\Bigl{\}}\Bigl{)},$ (7)
+Hence, we propose multi-frequency modulation to compel the model to fit high-frequency components, visualized in fig.˜3c and fig.˜3d:
+$\small\hat{I}(x)=\mathcal{P}\Bigl({\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl\{c_{ij}(x)\odot\psi(\alpha_{j}\cdot b_{i}(\gamma_{i}(x)))\Bigl\}\Bigl),$ (7)
 where $\psi\in{\sin,\cos}$ and $\alpha\in R$.
@@
 Consequently, if a basis is only accurate in the low-frequency domain, the output will contain undesired high-frequency noise.
-Alternatively, Eq.
-7 can be interpreted as making the spatial coordinates $x,y$ in Eq.
-1 learnable instead of the frequency components $u,v$.
-To validate this, we conduct experiments (Fig.
-5) showing that learning coordinates mitigates the rigid constraint of uniform distribution, aligning better with the non-uniform nature of visual patterns across spatial layouts; that is, some pixels share similar RGB values, indicating smooth color transitions, while others exhibit abrupt changes, suggesting high-frequency variations.
-To this end, our full Factorized Features representation is formulated in Eq. 7.
-Figure 5: In this figure, we validate the claim that fixed sinusoidal functions attend more to low-frequency parts.
-Learnable basis denotes the Learned Non-uniform basis in Sec.
-3.1.
-And we can see that Learnable Coordinate $x$ has the same effect, verifying our conclusion.
-Figure 6: The illustration of our joint image-compression and super-resolution framework compared with the traditional compression-only method.
-(a) Traditional learning-based compression methods.
-(b) Our approach surpasses (a) by incorporating our Super-Resolution (SR) Module from Sec.
-3.2 as information-recovery prior, detailed in Sec.
-3.3 (c) Expanding on (b),  Sec.
-3.3.1 introduces a multi-image compression strategy that utilizes both our SR Module and a Basis Merging Transformer to capture shared structure.
+Alternatively, eq.˜7 can be interpreted as making the spatial coordinates $x,y$ of $b_{i}(\mathbf{x})=\cos\left[(u_{i}x+v_{i}y)-\xi_{i}\right]$ in eq.˜3 learnable instead of the frequency components $u,v$.
+To validate this, we conduct experiments (fig.˜2(b)) showing that learning coordinates mitigates the rigid constraint of uniform distribution, aligning better with the non-uniform nature of visual patterns across spatial layouts; that is, some pixels share similar RGB values, indicating smooth color transitions, while others exhibit abrupt changes, suggesting high-frequency variations.
+To this end, our full Factorized Features representation is formulated in eq.˜7, visualized in fig.˜4.
+Figure 4: The Proposed Factorized Features.
+In this figure $\alpha=\{1,16\}$, $\psi=\text{sin}(\cdot)$, and that $\gamma_{i}(x)$ is coordinate transformation.
+We can see in this figure with sawtooth transformation $\gamma$ the feature is prompted to learn pattern from different scale in a patch-like manner, e.g.
+$\gamma_{2}$ makes the same basis repeat four times in the grid.
+However, only $\gamma$ is not enough for very fine details; therefore, we introduce $\alpha$ and $\psi$.
+Each basis is forced to accommodate both high-frequency (large $\alpha$ ) and low-frequency (small $\alpha$) components together with periodic functions $\psi$, i.e., larger $\alpha$ will make $\psi$ oscillate sharper.With explicit formulation of repetitive patterns and high frequency components, images can be represented with fine-grained details, which is especially useful for low-level vision tasks such as Super-Resolution and Image Compression.
 4.2 Super-Resolution with Factorized Features
 We represent a super-resolved image using our Factorized Features, where coefficients and basis are generated by networks $F_{\text{coeff}}$ and $F_{\text{basis}}$ from a low-resolution Image:
-$\hat{I}_{\text{SR}}(x)=\mathcal{P}\Bigl{(}{\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl{\{}c_{ij}^{\text{LR}}(x)\odot\psi(\alpha_{j}\cdot b_{i}^{\text{LR}}(\gamma_{i}(x)))\Bigl{\}}\Bigl{)},$ (8)
-where $c^{\text{LR}}(x)=\text{Conv}(F_{\textbf{coeff}}(I_{\text{LR}}))(x)$ and $b^{\text{LR}}(x)=F_{\textbf{basis}}(F_{\textbf{coeff}}(I_{\text{LR}}))(\gamma(x))$.
-Note that we sample the outputs $\text{Conv}(F_{\textbf{coeff}}(I_{\text{LR}}))$ and $F_{\textbf{basis}}(F_{\textbf{coeff}}(I_{\text{LR}}))$ with coordinates $x$ and $\gamma(x)$, respectively.
+$\small\hat{I}_{\text{SR}}(x)=\mathcal{P}\Bigl({\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl\{c_{ij}^{\text{LR}}(x)\odot\psi(\alpha_{j}\cdot b_{i}^{\text{LR}}(\gamma_{i}(x)))\Bigl\}\Bigl),$ (8)
+where $c^{\text{LR}}(x)=F_{\textbf{coeff}}(I_{\text{LR}})(x)$ and $b^{\text{LR}}(x)=F_{\textbf{basis}}(F_{\textbf{coeff}}(I_{\text{LR}}))(\gamma(x))$.
+Note that we sample the outputs $F_{\textbf{coeff}}(I_{\text{LR}})$ and $F_{\textbf{basis}}(F_{\textbf{coeff}}(I_{\text{LR}}))$ with coordinates $x$ and $\gamma(x)$, respectively.
+We omit convolution layers here for simiplicity.
 Our model comprises three main components: Coefficient Backbone $F_{\textbf{coeff}}$, Basis Swin Transformer $F_{\textbf{basis}}$, and Factorized Features Reconstruction.
-As shown in  Fig.
-3, the process begins with $I_{LR}\in\mathbb{R}^{3\times H\times W}$.
+As shown in  fig.˜5a, the process begins with $I_{LR}\in\mathbb{R}^{3\times H\times W}$.
 The Coefficient Backbone extracts features $X_{\text{coeff}}\in\mathbb{R}^{C_{c}\times H_{c}\times W_{c}}$, which are then used to generate coefficients $c$ through convolution and pixel shuffle operations.
-Then, $X_{\text{coeff}}$ is fed into the Basis Swin Transformer to produce a multi-scale basis $b=\{b_{1},...,b_{N}\},\ b_{i}\in\mathbb{R}^{C_{b_{i}}\times H_{b_{i}}\times W_{b_{i}}}$.
-The coefficients and basis are combined to reconstruct $I_{SR}\in\mathbb{R}^{3\times sH\times sW}$ using  Eq.
-8, where $s$ is the scale factor.
+Also, $X_{\text{coeff}}$ is fed into the Basis Swin Transformer to produce a multi-scale basis $b=\{b_{1},...,b_{N}\},\ b_{i}\in\mathbb{R}^{C_{b_{i}}\times H_{b_{i}}\times W_{b_{i}}}$.
+The coefficients and basis are combined to reconstruct $I_{SR}\in\mathbb{R}^{3\times sH\times sW}$ using  eq.˜8, where $s$ is the scale factor.
 We optimize model parameters using $L_{1}$ loss.
+To be more clear, when the Basis Swin Transformer is optimized, the basis is optimized simultaneously, i.e.
+the models are trained to adapt to a content-aware basis, not calculated by heuristic rules.
+This design enables content-aware basis generation: smooth regions get low-frequency bases while textured areas receive high-frequency ones.
 To demonstrate the effectiveness of our method, we use existing SR methods [16, 18, 109] as the Coefficient Backbone.
 For the Basis Swin Transformer, we employ simple Swin Transformer Blocks [70] with a series of special downsampling operations.
-Specifically, we use a dilation-like downsampling technique( Fig.
-2.c) to accommodate the sawtooth sampling pattern.
+Specifically, we use a dilation-like downsampling technique to accommodate the $\gamma$ sawtooth sampling pattern, where the details and visualization can be found in Supplementary Materials.
 The final basis is refined using additional upsampling and convolution layers.
 4.3 Image Compression with Factorized Features
-Image Compression, at its core, is to strike a balance between the amount of information contained in the latent bits and the final image quality.
-However, as discussed in  Sec.
-1 and  Sec.
-2, most recent works draw emphasis on how to retrieve implicit elements through designing different architectures, such as analysis transforms and entropy models [39, 44], or decompression modules [75, 24], while this paper aims at addressing the representation itself to better capture the structural correlations and thus achieve better image quality through Factorized Features formulation as in  Eq.
-7
+Image compression balances bit-rate and visual fidelity.
+Recent work recovers hidden details mainly via architectural tweaks—analysis transforms and entropy models [46, 51] or sophisticated decoders [88, 25].
+We instead target the representation itself: our Factorized Features formulation (eq.˜7) explicitly models structural correlations, boosting quality.
+Figure 5: Super-Resolution and Image Compression with Factorized Features.
+This figure illustrates how Factorized Features are used in Super-Resolution and Image Compression.
+(a) Given a low-resolution input image, we first extract $X_{coeff}$ feature with Coefficient Backbone.
+Next, we generate the basis and coefficient with Basis Swin Transformer and convolution layers, respectively, from the same $X_{coeff}$.
+Finally, the prediction is reconstructed by Factorized Features Reconstruction.
+(b) To decrease distortion in image compression, we replace the synthesis transform of the traditional learned image compression pipeline with (a) by aligning spatial resolution and latent channels.
 Priors from SR.
-In addition, with our trained SR model described in  Sec.
-3.2, it intuitively serves as a strong prior for information recovery, i.e., it contains extensive knowledge of how to reconstruct missing details and enhance image quality by leveraging learned patterns from the training data.
+In addition, with our trained SR model described in  section˜4.2, it intuitively serves as a strong prior for information recovery, i.e., it contains extensive knowledge of how to reconstruct missing details and enhance image quality by leveraging learned patterns from the training data.
 Thus, since Super-Resolution and Image Compression share the core principle of reconstructing and enhancing image details from low-quality sources, we can effectively integrate this prior into the compression pipeline.
-The overall pipeline is shown in  Fig.
-6b.
-To demonstrate the robustness of our representation and the effectiveness of the SR prior, the compression and decompression networks greatly follow [55], with only the synthesis transform replaced by our SR pipeline, where the details can be referenced in Supplementary Materials.
+The overall pipeline is shown in  fig.˜5b.
+To demonstrate the robustness of our representation and the effectiveness of the SR prior, the compression and decompression networks greatly follow [62], with only the synthesis transform replaced by our SR pipeline, where the details can be referenced in the Supplementary Materials.
 In practice, the training is performed in two stages.
-After we obtain the trained SR prior, the model is fine-tuned with a lower learning rate alongside the compression module, which is then trained end-to-end with the loss function defined as
-$L=R(\hat{y})+R(\hat{z})+\lambda\cdot D(x,\hat{x}),$ (9)
+After we obtain the trained SR prior, the model is fine-tuned with a lower learning rate alongside the compression module, which is then trained end-to-end with common loss function used in learned image compression, defined as
+[…]
+where $\hat{z}$ is the hyperprior, $R(\cdot)$ refers to the bit-rate cost, $D(\cdot)$ is the distortion term, and $\lambda$ controls the trade-off.
 4.3.1 Multi-Image Compression
-For each basis $b_{i}\in R^{C_{b_{i}}\times H_{b_{i}}\times W_{b_{i}}}$ associated with any arbitrary image, we can consider it as encapsulating the inherent pixel structure.
-These bases can be combined into a unified generic basis that captures the structural distribution of images and potentially reduces noise.
-Given $M$ images and their respective bases $b^{n}_{i},n\in\{1,...,M\}$, we apply a Basis Merging Transformer $F_{\text{merge}}$ at each location to integrate the $M$ elements:
-$b_{i}(h,w)=F_{\text{merge}}\left(\{b^{n}_{i}(h,w)\ |\ n\in{1,...,M}\}\right),$ (10)
-[…]
-Since compression induces error, according to  Fig.
-6b, the Coefficient feature map $X_{\text{coeff}}$ generated with Coefficient Backbone contains misinformation, and Basis Swin Transformer that uses such features would capture the information wrongfully for the basis and thus amplify error.
-To solve this, we exploit the mergeable property of basis, compress and transmit the merged basis independently along other quantized variables $\{Q(y,\mu),\ \hat{z}\}$, and finally reconstruct the $M$ images with the merged basis and their individual decoded coefficients.
-as illustrated in Fig.
-6c.
-This way, we can enjoy less information loss with basis while maintaining low bit rates.
-Table 1: Quantitative comparisons on 4×4$\times$4 × super-resolution with state-of-the-art methods.
+Intuitively, different images often share common frequency components.
+By learning a single, shared set of basis functions, we can leverage the frequently appearing similar or repetitive patterns across images to greatly reduce coding redundancy in multi-image compression.
+In practice, this means fewer total parameters and lower computational cost required when compressing an image collection.
+Each per-image basis $b_{i}\in\mathbb{R}^{C_{b_{i}}\times H_{b_{i}}\times W_{b_{i}}}$ captures local pixel structure.
+We fuse the $M$ bases into a denoised, generic basis with a transformer $F_{\text{merge}}$ [83] applied at every position:
+$\small b_{i}(h,w)=F_{\text{merge}}\left(\{b^{n}_{i}(h,w)\ |\ n\in{1,...,M}\}\right),$ (10)
+Treating bases as tokens (prepended with a CLS token), the transformer’s output forms the merged basis.
+Because compression noise corrupts the coefficient map $X_{\text{coeff}}$ (fig.˜5b), feeding it directly into the Basis Swin Transformer would amplify errors.
+Instead, we transmit the merged basis separately, alongside the quantized variables $^{Q(y,\mu),\hat{z}}$, and reconstruct each image with this clean basis plus its own decoded coefficients, yielding higher quality at similar bit-rates.
+A single- vs.
+multi-image compression comparison is in the supplementary materials.
+Table 1: Quantitative comparisons on $4\times$ super-resolution with state-of-the-art methods.
 The best results are colored red.
-The models with †are those who use same-task pretraining [16].
-Please refer to quantitative results in  Sec.
-4.1 for details.
+The models with †are those who use the same-task pretraining [16], i.e., pretrained on ImageNet.
+Please refer to quantitative results in  section˜5.1 for details.
 Params MACs Forward Pass Set5 Set14 B100 Urban100 Manga109
 […]
-SwinIR [48](ICCV2021W) 28.01 119.68 3,826 32.93 0.9043 29.15 0.7958 27.95 0.7484 27.56 0.8273 32.22 0.9273
-ATD [92](CVPR2024) 20.26 77.10 6,572 33.14 0.9061 29.25 0.7976 28.02 0.7524 28.22 0.8414 32.65 0.9308
-DAT [17])(ICCV2023) 14.80 61.66 4,192 33.15 0.9062 29.29 0.7983 28.03 0.7518 27.99 0.8365 32.67 0.9301
-RGT [18](ICLR2024) 13.37 834.25 3,404 33.16 0.9066 29.28 0.7979 28.03 0.7520 28.09 0.8388 32.68 0.9303
-HAT [15](CVPR2023) 20.77 86.02 3,692 33.18 0.9073 29.38 0.8001 28.05 0.7534 28.37 0.8447 32.87 0.9319
-HAT-L [15](CVPR2023) 40.84 167.27 6,804 33.30 0.9083 29.47 0.8015 28.09 0.7551 28.60 0.8498 33.09 0.9335
-ATD-L* 49.42 184.83 15,582 33.12 0.9062 29.31 0.7985 28.02 0.7514 28.25 0.8422 32.78 0.9309
-DAT-L* 43.01 175.42 11,326 33.33 0.9084 29.40 0.8009 28.04 0.7543 28.49 0.8473 33.02 0.9321
+EDSR [58](CVPR’17) 43.09 207.07 1,182 32.46 0.8968 28.80 0.7876 27.71 0.7420 26.64 0.8033 31.02 0.9148
+RCAN  [111](ECCV’18) 15.59 65.52 1,176 32.63 0.9002 28.87 0.7889 27.77 0.7436 26.82 0.8087 31.22 0.9173
+SwinIR [55](ICCV’21) 28.01 119.68 3,826 32.93 0.9043 29.15 0.7958 27.95 0.7484 27.56 0.8273 32.22 0.9273
+CAT-A+ [17](NIPS’22 Spotlight) 16.60 70.29 3,508 33.14 0.9059 29.23 0.7968 28.01 0.7516 27.99 0.8356 32.52 0.9293
+ART  [108](ICLR’23 Spotlight) 16.56 69.94 3,010 33.04 0.9051 29.16 0.7958 27.97 0.7510 27.77 0.8321 32.31 0.9283
+ATD [110](CVPR’24) 20.26 77.10 6,572 33.14 0.9061 29.25 0.7976 28.02 0.7524 28.22 0.8414 32.65 0.9308
+DAT [18])(ICCV’23) 14.80 61.66 4,192 33.15 0.9062 29.29 0.7983 28.03 0.7518 27.99 0.8365 32.67 0.9301
+RGT [19](ICLR’24) 13.37 834.25 3,404 33.16 0.9066 29.28 0.7979 28.03 0.7520 28.09 0.8388 32.68 0.9303
+PFT [71](CVPR’25) 19.66 80.21 6,414 33.15 0.9065 29.29 0.7978 28.02 0.7527 28.20 0.8412 32.63 0.9306
+HAT [15](CVPR’23) 20.77 86.02 3,692 33.18 0.9073 29.38 0.8001 28.05 0.7534 28.37 0.8447 32.87 0.9319
+ATD-L 49.42 184.83 15,582 33.15 0.9062 29.31 0.7985 28.02 0.7514 28.25 0.8422 32.78 0.9309
 ATD-F (Ours) 45.46 149.87 8,674 33.29 0.9082 29.48 0.8017 28.03 0.7539 28.53 0.8487 33.11 0.9335
+DAT-L 43.01 175.42 11,326 33.33 0.9084 29.40 0.8009 28.04 0.7543 28.49 0.8473 33.02 0.9321
 DAT-F (Ours) 40.00 134.42 6,206 33.45 0.9094 29.60 0.8039 28.13 0.7560 28.75 0.8520 33.23 0.9339
+HAT-L 40.84 167.27 6,804 33.30 0.9083 29.47 0.8015 28.09 0.7551 28.60 0.8498 33.09 0.9335
 HAT-F (Ours) 45.97 158.79 5,750 33.53 0.9100 29.65 0.8050 28.18 0.7569 28.79 0.8527 33.33 0.9342
-HAT-F-ImageNet (Ours) 45.97 158.79 5,750 33.55 0.9102 29.63 0.8049 28.18 0.7569 28.80 0.8529 33.33 0.9342
 HAT-L-F (Ours) 66.04 240.03 8,888 33.75 0.9116 29.87 0.8091 28.31 0.7597 29.51 0.8637 33.36 0.9343
-HAT-F-Basis-First (Ours) 46.67 161.66 5,696 33.33 0.9085 29.47 0.8015 28.10 0.7554 28.57 0.8494 33.14 0.9336
-HAT-F-Concat (Ours) 45.52 129.05 4,826 33.46 0.9095 29.57 0.8035 28.16 0.7566 28.73 0.8518 33.28 0.9341
+HAT-F-Basis-First (Ours, ablation) 46.67 161.66 5,696 33.33 0.9085 29.47 0.8015 28.10 0.7554 28.57 0.8494 33.14 0.9336
+HAT-F-Concat (Ours, ablation) 45.52 129.05 4,826 33.46 0.9095 29.57 0.8035 28.16 0.7566 28.73 0.8518 33.28 0.9341
 […]
 5 Experiments
+In this section, we provide the experiment setup and performance of Super-Resolution and Image Compression tasks.
+Please refer to Supplementary Materials for more ablation studies, training, implementation details, and analysis.
 5.1 Image Super-Resolution
 Experimental Setup.
-We conduct extensive experiments to validate the effectiveness of our Factorized Features representation for Super-Resolution tasks.
-Following the strategy outlined in [15] and [20], we adapt the same-task pretraining approach for all the Super-Resolution models.
-Unlike these previous works, we leverage the SA-1B dataset from [37], which includes approximately 10 million images, where we randomly sample 4 million for training, which is much less that of ImageNet [22], which is used for [15] and [20] pretraining.
-Note that we use SA-1B just for its content-rich and high-resolution images since SA-1B is designed for training segmentation models, whereas ImageNet focuses primarily on single-class prediction.
-We further conduct experiments to compare training performance with SA-1B and ImageNet in  Tab.
-1.
-While the performances are mostly equivalent, we witness faster convergence with SA-1B during training.
-To be more focused on the representation itself, we utilize various pre-trained SR models (SwinIR [48], HAT [16], DAT [17], and ATD [92]) as the Coefficient Backbone, with a consistent Basis Transformer architecture.
-Training involves initializing the Coefficient Backbones from pre-trained SR models and randomly initializing the Basis Transformers.
-Models are pre-trained for 300k iterations on the SA-1B dataset.
-After pretraining, we use DF2K (DIV2K [1] + Flickr2K [52]) as the finetuning dataset following [21, 15] for 200k iterations.
-For pretraining, we utilize AdamW optimizer with learning rate 2e-4, batch size 32, betas (0.9, 0.99), and other parameters set to PyTorch default, while we use learning rate 1e-5 during finetuning.
-Throughout training, the input is randomly cropped to $256\times 256$ and bicubicly resized to $64\times 64$ for Coefficient Backbone input.
-As for the hyperparameter of Factorized Features, the number $N$ of coefficient and basis is set to 6, and the scalars $\alpha_{j}$ are set to {1, 4, 16, 64} since we want to capture both base frequency $\alpha_{j}=1$ and high frequency $\alpha_{j}=64$ information.
-More implementation details can be found in supplementary materials.
+Following the same-task pretraining strategy of [15, 21], we pretrain all Super-Resolution (SR) models on ImageNet[23].
+Our Factorized Features plug a shared Basis Transformer into four pretrained SR backbones—SwinIR [55], HAT [16], DAT [18], and ATD [110].
+Backbones keep their original weights, while Basis Transformers start randomly.
+Pretraining runs 300k iterations on ImageNet with AdamW (lr 2e-4, batch 32, $\beta$ = 0.9/0.99); finetuning runs 200k iterations on DF2K (DIV2K [1]+Flickr2K [59]) with lr 1e-5.
+Inputs are $256\times 256$ crops, bicubically down-sampled to $64\times 64$ for the backbone.
+We set the number of coefficient–basis pairs to $N=6$ and use frequency scalars $\alpha_{j}\in\{1,4,16,64\}$ to capture both low- and high-frequency details.
 Quantitative Results.
-Tab.
-1 presents the quantitative comparison between our approach and state-of-the-art (SoTA) methods.
+table˜1 presents the quantitative comparison between our approach and state-of-the-art (SoTA) methods.
 We evaluate the methods using five benchmark datasets, including Set5 [6], Set14 [107], BSD100 [75], Urban100 [38], and Manga109 [76].
 For quantitative metrics, PSNR and SSIM are reported.
-The average relative improvement of 204.4% in PSNR over the baseline across the five datasets is calculated using the formula $(c-b)/(a-b)$, where $b$ represents the PSNR of a representative baseline, SwinIR, $a$ represents the PSNR of SOTA, HAT-L, and $c$ represents the PSNR of our best model, HAT-L-F.
-This formula measures the relative performance gain of our model compared to the gap between HAT-L and SwinIR.
-To validate the effectiveness of our framework, we employ three different SoTA SR models—ATD [91], DAT [17], and HAT [16]—as the Coefficient Backbone in our pipeline, which we denote as ATD-F, DAT-F, and HAT-F, respectively.
-These models exhibit significant improvements when compared to their counterparts, ATD-L, DAT-L, and HAT-L, which possess similar parameter counts.
-It is important to note that only [16] provides a large-scale model.
-To maintain fairness, we scale up ATD and DAT to match this size and train them under the same configuration as our model, including both pretraining and fine-tuning stages.
-To ensure an equitable comparison with other methods, we further train another model, HAT-F-ImageNet, using ImageNet as the pretraining dataset, following the protocols outlined in [14, 16].
-The results demonstrate that its performance remains consistent with only minor perturbations.
+Our model delivers a 204.4% average PSNR gain, defined as (cb)/(ab)(c-b)/(a-b) with bb=SwinIR, aa=HAT-L, and cc=HAT-L-F.
+We embed three SoTA SR backbones: ATD [109], DAT [18], HAT [16], yielding ATD-F, DAT-F, and HAT-F.
+Each beats its parameter-matched baseline (ATD-L, DAT-L, HAT-L).
+As only HAT provides a large variant, we upscale ATD and DAT to the same size and train all models with identical pre-training and fine-tuning for fairness.
+The Order of Coefficient and Basis.
 Furthermore, in traditional Fourier Series and other image processing methods [95], the basis is typically derived first and then used to compute the coefficients.
-In contrast, our method derives the coefficient features first, as illustrated in  Fig.
-3.
+In contrast, our method derives the coefficient features first, as illustrated in  fig.˜5a.
 To explore this difference, we develop another variant of our model, denoted HAT-F-Basis-First, where we reverse the order of operations.
@@
 Specifically, we argue that in our pipeline, the Coefficient Backbone functions more as a feature extraction module, where the refined features facilitate downstream basis extraction.
-Lastly, to evaluate the effectiveness of our Factorized Features, we trained a model named HAT-F-Concat, which does not apply the formulation in Eq.
-7.
+The Importance of Factorized Features Reconstruction.
+Lastly, to evaluate the effectiveness of our Factorized Features, we trained a model named HAT-F-Concat, which does not apply the formulation in eq.˜7.
 Instead, it concatenates the basis and coefficient directly and decodes the resulting features to produce the output.
-Although this approach results in reduced performance, which indicates the representation does act as an imperative role in modeling image information, the Basis Swin Transformer with Sawtooth downsampling still contributes to improved reconstruction, even without Factorized-Fields decoding, highlighting its effectiveness.
+Although this approach results in reduced performance, the Basis Swin Transformer with Sawtooth downsampling still contributes to improved reconstruction, even without Factorized Features Reconstruction, highlighting its effectiveness.
 Visual Comparison.
-We provide the visual comparison in Fig.
-7.
+We provide the visual comparison in fig.˜6.
 The images are randomly sampled from the DIV2K dataset.
 Our method faithfully reconstructs the image details, whereas the other approaches suffer from over-smoothing or hallucinating details absent in the ground truth.
-Multi-frame SR. Due to space limitations, we provide additional comparisons on multi-frame SR in the supplementary materials.
 (a) Kodak
@@
 Table 2: Comprehensive evaluation for image compression. Using VTM as an anchor for calculating BD-Rate. Latencies are measured under an NVIDIA GTX 3090 GPU.
-Method BD-Rate (%) $\downarrow$ Latency(s) Params(M)
-Tot Enc $\downarrow$ Tot Dec $\downarrow$
+Method BD-Rate (%) $\downarrow$ Total Encoding Time (s) $\downarrow$ Total Decoding Time (s) $\downarrow$ Params(M)
 VTM 0.00 129.21 0.14 -
-Xie (MM’21) -0.78 2.93 6.00 50.0
-Cheng (CVPR’22) 5.44 1.98 4.69 29.6
-STF (CVPR’22) -4.31 0.14 0.13 99.9
-ELIC (CVPR’22) -7.24 0.07 0.09 36.9
-TCM (CVPR’23) -11.74 0.16 0.15 76.7
-CCA (NeurIps’24) -20.97 0.11 0.09 64.9
+Cheng (CVPR’20)[20] 5.44 1.98 4.69 29.6
+Xie (MM’21)[101] -0.78 2.93 6.00 50.0
+STF (CVPR’22)[116] -4.31 0.14 0.13 99.9
+ELIC (CVPR’22)[36] -7.24 0.07 0.09 36.9
+TCM (CVPR’23)[62] -11.74 0.16 0.15 76.7
+SegPIC (ECCV’24) [67] -8.26 0.14 0.13 83.5
+LALIC (CVPR’25) [27] -15.26 0.27 0.15 63.2
 TCM-HAT-L-F (Ours) -21.09 0.109 0.264 110.34
-TCM-HAT-F-multi M=1 (Ours) 27.96 0.232 0.174 131.35
-TCM-HAT-F-multi M=2 (Ours) 2.70 0.232 0.174 131.35
-TCM-HAT-F-multi M=4 (Ours) -10.11 0.232 0.174 131.35
-TCM-HAT-F-multi M=8 (Ours) -16.61 0.232 0.174 131.35
-TCM-HAT-F-multi M=16 (Ours) -19.88 0.232 0.174 131.35
-TCM-HAT-F-multi M=24 (Ours) -20.97 0.232 0.174 131.35
-Table 3: Validation of the effectiveness of SR prior. The best PSNRs are marked in red.
-Method Kodak CLIC Tecnick
-$\lambda=0.0025$ $\lambda=0.0067$ $\lambda=0.025$ $\lambda=0.0025$ $\lambda=0.0067$ $\lambda=0.025$ $\lambda=0.0025$ $\lambda=0.0067$ $\lambda=0.025$
-[…]
-TCM[55] 0.1533 30.0834 0.2983 32.5841 0.6253 36.1345 0.1214 31.8207 0.2235 34.2098 0.4503 37.1201 0.1268 32.0588 0.2193 34.3669 0.3981 36.9066
-TCM-HAT-F-Scratch 0.1570 30.0857 0.2976 32.5893 0.6211 36.1389 0.1214 31.9421 0.2235 34.2894 0.4503 37.1434 0.1258 32.0632 0.2189 34.3781 0.4001 36.9223
-TCM-HAT 0.1567 30.1843 0.2992 32.6454 0.6268 36.2267 0.1220 31.9737 0.2266 34.3319 0.4512 37.2486 0.1262 32.1423 0.2174 34.5124 0.3971 36.9934
-TCM-HAT-F 0.1574 30.4012 0.2998 32.8910 0.6276 36.4461 0.1229 32.1917 0.2249 34.4109 0.4512 37.3135 0.1255 32.4591 0.2186 34.7656 0.3975 37.3244
+TCM-HAT-F-multi M=1 (Ours) 27.96 0.2320 0.1742 131.35
+TCM-HAT-F-multi M=2 (Ours) 2.70 0.1998 0.1270 131.35
+TCM-HAT-F-multi M=4 (Ours) -10.11 0.1846 0.1039 131.35
+TCM-HAT-F-multi M=8 (Ours) -16.61 0.1756 0.0922 131.35
+TCM-HAT-F-multi M=16 (Ours) -19.88 0.1715 0.0863 131.35
+TCM-HAT-F-multi M=24 (Ours) -20.97 0.1702 0.0844 131.35
 5.2 Single- and Multi-Image Compressions
 Experimental Setup.
-We evaluate our Factorized Features for image compression tasks against state-of-the-art methods.
-Following the Super-Resolution setup in Sec.
-4.1, we train on the SA-1B dataset.
-To emphasize our representation, compression and decompression modules are initialized from pre-trained models [55], then fine-tuned end-to-end on 256×256 patches for 200k iterations using the AdamW optimizer [60] (learning rate=1e-5, batch size=16, betas=(0.9, 0.99)).
-We integrate our SR Module with the pre-trained compression module TCM, creating TCM-HAT-F and TCM-HAT-L-F models.
-TCM-HAT-F-multi represents the multi-image compression pipeline.
-For multi-image compression in Fig.
-6.c, we set the Basis Swin Transformer and the Basis Merging Transformer to be trainable while the other parts remain frozen.
+Image-compression experiments follow the SR protocol (section˜5.1).
+Models are pretrained on ImageNet and fine-tuned for 200k steps on $\times256$ crops.
+Compression/decompression networks start from TCM weights [62] and are optimized end-to-end with AdamW (lr 1e-5, batch 16, $\beta$ = 0.9/0.99) [72].
+Integrating our SR module into TCM yields TCM-HAT-F and TCM-HAT-L-F; TCM-HAT-F-multi is the multi-image variant.
 Rate-Distortion Performance Comparison.
 We compare our model with State-of-the-Art learned end-to-end image compression algorithms, including [63], [13], [116], [101], [20], [3], [51], [40], [79], [5], [84], and [36].
-The classical image compression codec, VVC [80], is also tested by using VTM12.1.The rate-distortion performance on various datasets, including Kodak, Tecnick’s old test set with resolution 1200$\times$1200, and CLIC Professional Validation, is shown in Fig.
-8.
-In  Tab.
-2, our TCM-HAT-L-F model achieved a significant BD-Rate improvement of -21.09% compared to VTM, outperforming previous state-of-the-art methods.
-The multi-image compression approach (TCM-HAT-F-multi) shows increasing performance gains with the number of images compressed simultaneously, reaching -20.97% BD-Rate improvement for $M=24$.
-The result shows that direct transmission of bases would reduce the error from Coefficient Backbone to Basis Swin Transformer and that the distortion increases with $M$, as the information contained in the merged basis is limited, and merging multiple bases into one would cause increasing information loss.
-Our experiments validate the effectiveness of FIPER in image compression, demonstrating substantial improvements across benchmarks (Kodak, CLIC, Tecnick).
-The multi-image compression strategy particularly excels with larger image sets, highlighting strong scalability.
-Additionally, our method achieves competitive latency, significantly improving compression efficiency while balancing quality and speed.
-Table 4: Comparison of improvements of Factorized Features. $\psi$ and $\alpha$ are the same in Eq. 8.
-Metric PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-Baseline [12] 22.04 0.505 0.5296
-Ours 38.44 0.999 0.0385
-No $\psi$ to control magnitude 13.46 0.147 0.766
-No $\alpha$ for pixel-wise frequency information 21.25 0.537 0.527
-4.3 Ablation Studies
-Influence of SR Priors in Image Compression.
-We conduct experiments with various configurations to verify our image compression pipeline’s effectiveness.
-TCM-HAT refers to using the original HAT [16] instead of our SR Module in Fig.
-6b.
-TCM-HAT-F represents our full pipeline, while TCM-HAT-F-Scratch initializes the SR Module randomly.
-Results demonstrate that integrating SR priors improves image compression performance, and our representation further enhances results, highlighting the robustness of Factorized Fields.
-Effectiveness of Factorized Features Design.
-We conduct experiments to verify our multi-frequency modulation.
-The quantitative performance reported on single-image regression is shown in Tab.
-4, where each result is measured after 256 iterations.
-Compared to baseline results, our refinements in modeling pixel-level frequency have significantly improved all performance metrics.
-Additionally, our results demonstrate that the modulation function $\psi$ and the scalar $\alpha$ are interdependent, each essential to the other’s function.
+The classical image compression codec, VVC [93], is also tested by using VTM12.1.The rate-distortion performance on various datasets, including Kodak, Tecnick’s old test set with resolution $\times1200$, and CLIC Professional Validation, is shown in fig.˜7.
+TCM-HAT-L-F achieves a -21.09% BD-Rate vs.
+VTM (Table 2), surpassing earlier work.
+Multi-image mode (TCM-HAT-F-multi) scales well, reaching -20.97% at M=24M{=}24: sending a shared basis curbs Coefficient-to-Basis error, though distortion grows slightly as basis capacity is stretched.
+Across Kodak, CLIC, and Tecnick, FIPER consistently boosts rate-distortion and maintains competitive latency.
 6 Conclusion
@@
 Additionally, incorporating semantic information remains a promising direction for future research.
-[…]
-Supplementary Material
+Acknowledgements.
+This research was funded by the National Science and Technology Council, Taiwan, under Grants NSTC 112-2222-E-A49-004-MY2, 113-2628-E-A49-023-, and Academia Sinica under Grant AS- CDA-110-M09.
+The authors are grateful to Google, NVIDIA, and MediaTek Inc.
+for their generous donations, and we would also like to express our sincere gratitude to Jia-Wei Liao for valuable assistance in paper writing and revision.
+Yu-Lun Liu acknowledges the Yushan Young Fellow Program by the MOE in Taiwan.
 Appendix A Overview
 This supplementary material provides additional implementation details and experimental results to complement our main manuscript.
-Specifically, we first elaborate on the background of factor fields and learned image compression in Appendix B, which were omitted in the main paper due to space limitations.
-Next, in Appendix D, we present comprehensive details about the benchmarks, including the Video Super-Resolution task, which was not discussed in the main paper.
-Finally, we provide extensive visualizations to facilitate a more detailed comparison of the results in Appendix H.
-Appendix B Preliminaries
-In this section, we briefly introduce the background of the factor fields [12], and Learned Image Compression along with their properties.
-Factor Fields.
-The concept of decomposition fields or factorized matrices in the reconstruction of 2D images or 3D scenes has shown superior rendering quality and improved efficiency recently [11, 27, 68, 26, 7, 9], while [12] first proposed a unified representation called factor fields.
-For a 1D signal $s(x)$, the factor fields formulation is:
-$\hat{s}(x)=c(x)^{T}b(\gamma(x)),$ (11)
-where $c(x)=(c_{1}(x),...,c_{N}(x))^{T}$ are spatial-varying coefficient fields, $b(x)=(b_{1}(x),...,b_{N}(x))^{T}$ are basis functions, and $\gamma(x)$ is a coordinate transformation.
-The coefficient and basis are queried by sampling with the coordinates $x$.
-This generalizes to Q-dimensional signals:
-$\hat{s}(x)=\mathcal{P}(c(x)\odot b(\gamma(x))),$ (12)
-where $\mathcal{P}$ is a projection function and $\odot$ denotes the element-wise product. In practice, for multi-dimensional coefficients:
-$\hat{s}(x)=\mathcal{P}\Bigl{(}\textbf{Concat}_{i=1}^{N}\Bigl{\{}c_{i}(x)\odot b_{i}(\gamma_{i}(x))\Bigl{\}}\Bigl{)}.\end{aligned}$
-This formulation allows the same basis to be applied at multiple spatial locations while varying the coefficients, especially when $\gamma$ is a periodic function.
-Learned Image Compression.
+Our FIPER framework explicity models multi-scale patterns and captures periodic features effectively, as visualized in fig.˜8.
+[…]
+In the following, we dive into the details.
+Specifically, we first elaborate on coordinate transformation, the background of learned image compression, and a comparison between single- and multi-image compression in appendix˜B, which were omitted in the main paper due to space limitations.
+Following this, we discuss the differences with factor fields [12] in appendix˜C.
+Next, in appendix˜D, we present comprehensive details about the experiments and benchmarks, including the Video Super-Resolution task, which was not discussed in the main paper.
+Then, we provide ablation studies regarding different components of our pipeline, validating the design in appendix˜E.
+Finally, we provide extensive visualizations to facilitate a more detailed comparison of the results in appendix˜H.
+Appendix B Method Details
+In this section, we provide the details of Sawtooth coordinate transformation and downsampling inspired by it, the background of Learned Image Compression, and comparison between single- and multi-image compression.
+B.1 Sawtooth Coordinate Transformation and Downsampling
+Figure 9: The correlation between coordinate transformation and downsampling.
+(a) The sawtooth transformation example with $k=2$.
+(b) The PixelUnShuffle downsample.
+(c) To explicitly model the information for sampling with a sawtooth, we rearrange the feature map in a dilation-like manner in the downsample layer of the Basis Swin Transformer.
+This way, the feature sampled would capture the information in the original layout correctly; in other words, $I=\text{Sawtooth-Pixel-Sampling}(\text{Sawtooth-Aware-Downsample}(I))$
+Sawtooth transformation
+Sawtooth transformation is formulated as $\gamma(x)=x\mod k,\ k\in R$.We can easily observe that such transformation implicitly captures patch-like frequency information as shown in fig.˜9a, and thus, we propose that by leveraging the inter-patch information from the sawtooth coordinate transformation, the visual correspondence between spatial locations can be effectively represented.
+Downsampling based on Sawtooth transformation
+To reconstruct a $256\times 256$ image, we formulate our Factorized Features as
+$\hat{I}(x)=\mathcal{P}\Bigl({\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl\{c_{ij}(x)\odot\psi(\alpha_{j}\cdot b_{i}(\gamma_{i}(x)))\Bigl\}\Bigl),$ (11)
+which is the same as eq.˜7 in the main paper.
+In practice, the bases are of shape $C\times H\times W$ with $N=6$.
+We set $H=W=\{256,\ 128,\ 64,\ 32,\ 16,\ 8\}$ and $C=24$ for all bases.
+Thus, we sample each basis with $\gamma_{i}(x)=x\mod k_{i}\ \text{where}\ x\in\{0,...,255\}^{2}$ and $k=\{256,128,64,32,16,8\}$.
+This formulation can be interpreted as we sample the same feature on different spatial location of a image, e.g., when using $b_{1}\in R^{24\times 128\times 128}$ with $k_{1}=128$, $x=\{0,0\}$ and $x=\{128,0\}$ will be extracting $b_{1}(0,0)$, which is also visualized in fig.˜9a.
+Therefore, we want to aggregate the information appearing in different location of the image, where these spatial locations will be sampled at the same location on the basis; that is, we want the information of $I(0,0)$ and $I(128,0)$ to be aggregated in $b_{1}(0,0)$ in previous example.
+This way, we can explicitly model repetitve structure in different scale, and such frequency modeling subsequently enhances the reconstruction.
+Based on the analysis, as visualized in fig.˜9c, we integrate Sawtooth-Aware Downsample layer into Basis Transformer.
+Such operation is essentially the inverse of Sawtooth transformation, i.e., we merge the feature together with Sawtooth-Aware Downsample, decode the feature to basis, and then sample the basis to different spatial location via Sawtooth transformation.
+B.2 Learned Image Compression.
 Following [79, 3], a learned image compression model with a channel-wise entropy model can be formulated as:
 $z$ $=h_{a}(y;\phi_{h}),\ y=g_{a}(x;\phi),$ (12)
-$\{X_{\text{mean}},X_{\text{scale}}\}$ $=h_{s}(\hat{z};\theta_{h}),\ \hat{z}=Q(z),$
-$\hat{y}$ =$=\{Q(y_{0}-\mu_{0})+\mu_{0},...,Q(y_{t}-\mu_{t})+\mu_{t}\},$
-$\hat{x}$ $=g_{s}(\overline{y};\theta),\ \overline{y}=\textbf{Refine}_{\theta_{r}}({\mu_{0},...,\mu_{t}},\hat{y}),$
-where $0<=t<l,\ \mu_{t}=e_{i}(\hat{y}_{<i},X_{\text{mean}})$.
+$\{F_{\text{mean}},F_{\text{scale}}\}$ $=h_{s}(\hat{z};\theta_{h}),\ \hat{z}=Q(z),$
+$\hat{y}$ $=\{Q(y_{0}-\mu_{0})+\mu_{0},.,Q(y_{t}-\mu_{t})+\mu_{t}\},$
+[…]
+where $0<=t<l,\ \mu_{t}=e_{i}(\hat{y}_{<i},F_{\text{mean}})$.
 The encoder $g_{a}$ transforms the raw image $x$ into a latent representation $y$.
@@
 The derivation of the equation can be found in [2].
-Next, $\hat{z}$ is decoded by $h_{s}$ to produce features $X_{\text{mean}}$ and $X_{\text{scale}}$, used to estimate the mean $\mu$ and variance $\sigma$ of $y$.
+Next, $\hat{z}$ is decoded by $h_{s}$ to produce features $F_{\text{mean}}$ and $F_{\text{scale}}$, used to estimate the mean $\mu$ and variance $\sigma$ of $y$.
 The latent $y$ is divided into $l$ slices, and each quantized around computed means $\mu_{t}$.
-These means are derived from earlier quantized slices and $X_{\text{mean}}$ by a slice network $e_{i}$.
+These $\mu_{t}$ are derived from earlier quantized slices and $F_{\text{mean}}$ by a slice network $e_{i}$.
 The quantized slices form $\hat{y}$.
@@
 The model is trained using a Lagrangian multiplier-based rate-distortion optimization:
-$L$ $=R(\hat{y})+R(\hat{z})+\lambda\cdot D(x,\hat{x})$ (15)
+[…]
 $=\mathbb{E}\left[-\log_{2}\left(p_{\hat{{y}}|\hat{{z}}}\left(\hat{{y}}\mid\hat{{z}}\right)\right)\right]+\mathbb{E}\left[-\log_{2}\left(p_{\hat{{z}}|\varphi}\left(\hat{{z}}\mid\varphi\right)\right)\right]$
-$+\lambda\cdot\mathcal{D}\left({x},\hat{{x}}\right)$
-where $R(\hat{y})$ and $R(\hat{z})$ denote bit rates, $D(x,\hat{x})$ is the distortion term (calculated by MSE), and $\lambda$ balances compression efficiency and image fidelity.
+[…]
+where R(y^)R($\hat${y}) and R(z^)R($\hat${z}) denote bit rates, D(x,x^)D($\hat{x}$) is the distortion term (calculated by MSE), and $\lambda$ balances compression efficiency and image fidelity.
 In our experiments, we follow [62], modifying only $g_{s}$ to demonstrate our representation’s effectiveness.
-Appendix C Differences from Factor Fields.
+B.3 Comparison between single- and multi-image compression
+Figure 10: The illustration of our image-compression framework (a) Traditional learning-based compression methods.
+(b) Our approach surpasses (a) by incorporating our Factorized Features and Super-Resolution (SR) Module from section˜4.2 in the main paper as information-recovery prior.
+(c) Expanding on (b), we introduces a multi-image compression strategy that utilizes both our SR Module and a Basis Merging Transformer to capture shared structure.
+We provide visual comparison of single- and multi-image compression in fig.˜10b and fig.˜10c, respectively, with traditional compression method in fig.˜10a.
+First, we have single image compression(SIC), the same as fig.˜5b in the main paper.
+The goal of the SIC pipeline is to validate two aspects: (1) that Factorized Features, which are designed to represent images, can be applied to different tasks; (2) that leveraging super‐resolution (SR) priors improves reconstruction quality.
+Note that we directly adopt the SR module’s architecture to also reduce overall complexity.
+Next, we derive multi-image compression (MIC) as in section˜4.3.1 of the main paper.
+In traditional signal processing, a shared set of bases is often used to represent diverse sources.
+We adopt this property to further reduce compression redundancy.
+Specifically, our approach transmits (1) image features, which are used to decode coefficients, and (2) the bases themselves.
+By contrast, single image compression in fig.˜10b transmits only image features.
+Intuitively, transmitting the bases provides additional information that can improve reconstruction quality.
+However, directly transmitting all bases would significantly increase the bit rate.
+To address this, we merge the bases into a single set before transmission, thereby enhancing reconstruction while minimizing extra transmission costs.
+Appendix C Differences from Factor Fields
 We explicitly highlight how Sawtooth-based sampling and Multi-frequency Modulation extend the original factor fields.
-In our 1D and 2D toy examples as shown in Figs.
-9 and 10, traditional SR/IC representations can be viewed as a “vanilla” scenario with $N=1$ and $c(x)=1$.
+In 2D toy examples as shown in fig.˜11, traditional SR/IC representations can be viewed as a “vanilla” scenario with $N=1$ and $c(x)=1$.
 factor fields already improve accuracy over such baselines, but our introduction of $\psi$ and $\alpha$ further captures both high- and low-frequency details.
 Specifically, forcing $\alpha=4$ compels the network to learn broader frequency ranges and avoid noise from the fixed modulation.
-To confirm these advantages, we trained smaller SR models and compared them with factor fields in Tab.
-5.
+To confirm these advantages, we trained smaller SR models and compared them with factor fields in table˜3.
 The results illustrate that our multi-frequency constraint yields significant gains even in simpler configurations, underscoring its importance for high-frequency detail preservation and overall reconstruction quality.
-Refer to caption Refer to caption Refer to caption Refer to caption
-(a) Vanilla fields (b) Factor fields [10] (c) Ours with $\psi$ only (d) Ours with both $\psi$ and $\alpha$
-Figure 9: 1D toy example.
 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
 Figure 11: 2D toy example.
-Table 5: Comparison with Factor Fields.
+Table 3: Comparison with Factor Fields. HAT-FA-Small denotes replace our Factorized Features with factor fields.
 Method Sawtooth $\psi$ $\alpha$ Params MACs Forward Pass Set5 Set14 B100 Urban100 Manga109
@@
 Figure 12: Detailed architecture of Super-Resolution Modules
-The architecture of our SR networks is shown in Fig.
-11.
+The architecture of our SR networks is shown in fig.˜12.
 Firstly, for Coefficient Backbones from various pre-trained models like [15, 109, 18], we replace the output upsampling layer (mostly pixel shuffle) with convolution of channel $D=256$, i.e., for input image with resolution $64\times 64$, the original pre-trained models process hidden states in $64\times 64$, and then upsample the output to $256\times 256$ for final result, where we replace the upsampling with convolution and output the final hidden state $X_{coeff}\in R^{256\times 64\times 64}$.
@@
 Then, the Sawtooth-Aware Downsample block rearranges the hidden state $h$ as discussed in the main paper, which can be viewed as:
-$h=\mathbf{rearrange}(h,\text{'b c (s h) (s w) $\rightarrow$ b (s s c) h w'}),\ s=2$ (16)
+[…]
 , and reduces the feature channel by 4. In all experiments we use $N=6$, i.e. there are 6 bases, where $Basis_{i}\in R^{24\times\frac{256}{2^{i}}\times\frac{256}{2^{i}}}\text{for i in }\{0\ ...\ 5\}$.
-Finally, we derive our Coefficient by a series of Convolution and Pixelshuffle layers and decoded the Bases and Coefficient to a predicted high-resolution image by Eq. 17:
+Finally, we derive our Coefficient by a series of Convolution and Pixelshuffle layers and decoded the Bases and Coefficient to a predicted high-resolution image by eq.˜15:
 $\hat{I}(x)=\mathcal{P}\Bigl({\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl\{c_{i}(x)\odot\psi(\alpha_{j}\cdot b_{i}(\gamma_{i}(x)))\Bigl\}\Bigl).$ (15)
@@
 Next, we apply positional embedding to the tokens, process the tokens with the transformer blocks, and finally decode the Coefficient and Basis token from the respective position with convolution layers for the output super-resolved frames.
-Tab.
-6 shows quantitative comparison with state-of-the-art methods.
+table˜4 shows quantitative comparison with state-of-the-art methods.
 All VSR experiments were conducted using bicubic 4X downsampling.
@@
 Figure 13: Detailed architecture of Compression Modules
-As discussed in the main paper, to make a fair comparison with the state-of-the-art, our entropy model and compression backbone greatly follow that of [55], as shown in Fig.
-12(more details can be bound in [55].
+As discussed in the main paper, to make a fair comparison with the state-of-the-art, our entropy model and compression backbone greatly follow that of [62], as shown in fig.˜13(more details can be bound in [62].
 What we do here is that we use the output from the last main layer and convert it to match the hidden channel of the SR Module.
@@
 Following VSR, we use transformer blocks to merge the bases.
-The difference here is that we additionally prepend a CLS token so that we can use it as the final merged basis.
-For the merged basis compression, we utilize the same architecture of the image branch, i.e., we randomly initialize a module from [55] with input and out channels change.
+The difference here is that we additionally prepend a CLS token so that we can use it as the final merged basis, i.e., if we have 16 bases of $W=H=64$ to merge into one, we rearrange the $64\times 64$ tokens as batch dimension, use the $16+1$ as sequence dim, and finally treat the additional 11(CLS) token as output.
+For the merged basis compression and transmission, we utilize the same architecture of the image branch, i.e., we randomly initialize a module from [62] with input and out channels change.
 After the Basis Merging Transformer is trained, we follow the same training setting for Basis compression while randomly sampling 1 to 24 images for a batch and setting only the Basis compression module trainable.
-Appendix E Merged-Basis Strategy & SR Priors.
+Appendix E Ablation Studies
+E.1 Merged-Basis Strategy & SR Priors.
 The SR prior accelerates detail recovery by leveraging high-resolution knowledge while basis merging exploits shared structures for efficiency.
 Since merging is introduced after, the compression model is pre-trained, and improvements stem not only from the SR prior but also from the well-initialized basis merging.
-When trained from scratch (see Tab.
-7), the merging transformer tends to find a universal solution rather than image-specific structures.
+When trained from scratch (see table˜5), the merging transformer tends to find a universal solution rather than image-specific structures.
 Additionally, merging inevitably causes information loss, prompting the coefficient branch to encode more image-specific details.
@@
 Table 5: Comparison of Merged-Basis Strategies.
-Appendix F Encoding/Decoding Times.
+E.2 Encoding/Decoding Times.
 Fig.
@@
 2, different $M$ values primarily affect the run time of the Basis Merging Transformer.
-Although merging larger $M$ increases computation, per-image overhead remains small once averaged by $M$, as shown in Tab.
-8.
+Although merging larger $M$ increases computation, per-image overhead remains small once averaged by $M$, as shown in table˜6.
 Notably, Tab.
@@
 Adjusted Dec Time (sec) by $M$ 0.1740 0.1272 0.1039 0.0922 0.0863 0.0844
+BD-Rate $\%$ 27.96 2.70 -10.11 -16.61 -19.88 -20.97
+Table 6: Comparison of Encoding/Decoding Times.
+E.3 Inference Time Comparison for SR.
+We provide the inference time in table˜7, compared with extensive SR models. The inference time of different components of FIPER is also provided.
+ATD DAT RGT HAT-S HAT HAT-L HAT-F(Ours) Coefficient Backbone (HAT-F) Basis Swin Transformer (HAT-F) Factorized Features Reconstruction (HAT-F)
+Inference Time (s) 0.0592 0.0594 0.0612 0.0387 0.0398 0.0778 0.0716 0.0400 0.0132 0.0184
+PSNR ont Set5 33.14 33.15 33.16 32.92 33.18 33.30 33.53 - - -
+Table 7: Inference time comparison for SR.
+E.4 Influence of SR Priors in Image Compression.
+We conduct experiments with various configurations to verify our image compression pipeline’s effectiveness.
+TCM-HAT refers to using the original HAT [16] instead of our SR Module in fig.˜10b.
+TCM-HAT-F represents our full pipeline, while TCM-HAT-F-Scratch initializes the SR Module randomly.
+Results demonstrate that integrating SR priors improves image compression performance, and our representation further enhances results, highlighting the robustness of Factorized Features.
+Table 8: Validation of the effectiveness of SR prior. The best PSNRs are marked in red.
+Method Kodak CLIC Tecnick
+$\lambda=0.0025$ $\lambda=0.0067$ $\lambda=0.025$ $\lambda=0.0025$ $\lambda=0.0067$ $\lambda=0.025$ $\lambda=0.0025$ $\lambda=0.0067$ $\lambda=0.025$
 […]
-Table 8: Comparison of Encoding/Decoding Times.
-Appendix G Inference Time Comparison for SR.
-We provide the inference time in Tab. 9, compared with extensive SR models. The inference time of different components of FIPER is also provided.
-HAT-S HAT HAT-L ATD DAT RGT HAT-F-S (Ours) HAT-F(Ours) HAT-L-F(Ours) Coefficient Backbone (HAT-F) Basis Swin Transformer (HAT-F) Factorized Features Reconstruction (HAT-F)
-Inference Time (s) 0.0387 0.0398 0.0778 0.0592 0.0594 0.0612 0.0671 0.0716 0.1084 0.0400 0.0132 0.0184
-Table 9: Inference time comparison for SR.
+TCM[62] 0.1533 30.0834 0.2983 32.5841 0.6253 36.1345 0.1214 31.8207 0.2235 34.2098 0.4503 37.1201 0.1268 32.0588 0.2193 34.3669 0.3981 36.9066
+TCM-HAT-F-Scratch 0.1570 30.0857 0.2976 32.5893 0.6211 36.1389 0.1214 31.9421 0.2235 34.2894 0.4503 37.1434 0.1258 32.0632 0.2189 34.3781 0.4001 36.9223
+TCM-HAT 0.1567 30.1843 0.2992 32.6454 0.6268 36.2267 0.1220 31.9737 0.2266 34.3319 0.4512 37.2486 0.1262 32.1423 0.2174 34.5124 0.3971 36.9934
+TCM-HAT-F 0.1574 30.4012 0.2998 32.8910 0.6276 36.4461 0.1229 32.1917 0.2249 34.4109 0.4512 37.3135 0.1255 32.4591 0.2186 34.7656 0.3975 37.3244
+E.5 Effectiveness of Factorized Features Design.
+We conduct experiments to verify our multi-frequency modulation.
+The quantitative performance reported on single-image regression is shown in table˜9, where each result is measured after 256 iterations.
+Compared to baseline results, our refinements in modeling pixel-level frequency have significantly improved all performance metrics.
+Additionally, our results demonstrate that the modulation function $\psi$ and the scalar $\alpha$ are interdependent, each essential to the other’s function.
+Table 9: Comparison of improvements of Factorized Features. $\psi$ and $\alpha$ are the same in eq.˜8. Note that here we test on single-image regression
+Metric PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
+Baseline [12] 22.04 0.505 0.5296
+Ours 38.44 0.999 0.0385
+No $\psi$ to control magnitude 13.46 0.147 0.766
+No $\alpha$ for pixel-wise frequency information 21.25 0.537 0.527
+E.6 Light-weight Model Comparison
+To demonstrate the effectiveness of our method even under computation-constrained scenarios, we integrate our Factorized Features framework to state-of-the-art lightweight super-resolution models and observe consistent improvement as shown in table˜10.
+Method Latency (ms) #FLOPs (G) #Params (K) Set5 PSNR / SSIM Set14 PSNR / SSIM B100 PSNR / SSIM Urban100 PSNR / SSIM Manga109 PSNR / SSIM
+SwinIR-lt 222.9 63.6 930 32.44 / 0.8976 28.77 / 0.7858 27.69 / 0.7406 26.47 / 0.7980 30.92 / 0.9151
+ELAN-lt 18.0 54.1 640 32.43 / 0.8975 28.78 / 0.7858 27.69 / 0.7406 26.54 / 0.7982 30.92 / 0.9150
+OmniSR 22.5 50.9 792 32.49 / 0.8988 28.78 / 0.7859 27.71 / 0.7415 26.64 / 0.8018 31.02 / 0.9150
+SRFormer-lt 287.2 62.8 873 32.51 / 0.8988 28.82 / 0.7872 27.73 / 0.7422 26.67 / 0.8032 31.17 / 0.9165
+ATD-lt 189.7 100.1 769 32.63 / 0.8998 28.89 / 0.7886 27.79 / 0.7440 26.97 / 0.8107 31.48 / 0.9198
+HiT-SRF 82.1 58.0 866 32.55 / 0.8999 28.87 / 0.7880 27.75 / 0.7432 26.80 / 0.8069 31.26 / 0.9171
+ASID-D8 61.8 49.6 748 32.57 / 0.8990 28.89 / 0.7898 27.78 / 0.7449 26.89 / 0.8096
+MambaIR-lt 55.8 84.6 924 32.42 / 0.8977 28.74 / 0.7847 27.68 / 0.7400 26.52 / 0.7983 30.94 / 0.9135
+MambaIRV2-lt 153.4 75.6 790 32.51 / 0.8992 28.84 / 0.7878 27.75 / 0.7426 26.82 / 0.8079 31.24 / 0.9182
+RDN 66.0 1309.2 22271 32.47 / 0.8990 28.81 / 0.7871 27.72 / 0.7419 26.61 / 0.8028 31.00 / 0.9151
+RCAN 52.2 917.6 15592 32.63 / 0.9002 28.87 / 0.7889 27.77 / 0.7436 26.82 / 0.8087 31.22 / 0.9173
+CATANet (CVPR’25) 102.4 49.3 535 32.58 / 0.8998 28.90 / 0.7880 27.75 / 0.7427 26.87 / 0.8081 31.31 / 0.9183
+ESC (ICCV’25) 21.9 149.2 968 32.68 / 0.9011 28.93 / 0.7902 27.80 / 0.7447 27.07 / 0.8144 31.54 / 0.9207
+CATANet-F (Ours) 110.3 56.7 941 32.69 / 0.9017 28.93 / 0.7906 27.82 / 0.7444 27.05 / 0.8133 31.62 / 0.9214
+ESC-F (Ours) 24.6 73.4 952 32.74 / 0.9032 28.99 / 0.7931 27.86 / 0.7483 27.11 / 0.8158 31.78 / 0.9233
+Table 10: Quantitative comparison on benchmark datasets for $\times 4$ lightweight image super-resolution.
+E.7 Task-Specific Overfitting
+To confirm that our Factorized Features are not overfitted to a single task, we further apply the same representation to other low-level vision problems such as motion deblurring in table˜11 and image dehazing in table˜12, without any task-specific modification.
+The consistent improvements across these tasks demonstrate that our features generalize well and capture transferable visual priors.
+Metric / Method DeblurGAN-v2 SRN DMPHN SDWNet MPRNet MIMO-UNet+ DeepRFT+ MAXIM-3S Stripformer MSDI-net Restormer NAFNet FFTformer GRL-B MLWNet (CVPR’24) MLWNet-F (Ours)
+PSNR 29.55 30.26 31.20 31.26 32.66 32.45 33.23 32.86 33.08 33.28 33.57 33.69 34.21 33.93 33.83 34.40
+SSIM 0.934 0.934 0.945 0.966 0.959 0.957 0.963 0.961 0.962 0.964 0.966 0.967 0.969 0.968 0.968 0.971
+Table 11: Performance comparison on the GoPro dataset for motion deblurring.
+Metric / Method DehazeNet AOD-Net MSBDN FFA-Net AECR-Net DeHamer PMNet DehazeFormer TaylorFormer LH-Net MITNet PGH2Net (AAAI’24) ConvIR (TPAMI’24) ConvIR-F (Ours)
+PSNR 13.84 13.14 15.37 14.39 15.80 16.62 16.79 16.29 16.66 18.87 16.97 17.02 17.45 18.12
+SSIM 0.43 0.41 0.49 0.45 0.47 0.56 0.51 0.51 0.56 0.561 0.606 0.61 0.802 0.823
+Table 12: Performance comparison on the Dense-Haze dataset for image dehazing.
+E.8 Analysis of Model Hyperparameters
+We further conduct a comprehensive study on three key hyperparameters of our model: the number of frequency components $N$, the frequency modulation $\alpha$, and the basis transformation $\gamma$ in table˜13 and table˜14.
+These factors collectively control the expressiveness and efficiency of the proposed Factorized Features.
+NN (Number of Frequency Components) 1 1 2 6 6 6 6 10
+$\alpha$ $\{1\}$ $\{1,4,16,64\}$ $\{1,4,16,64\}$ $\{1\}$ $\{1,4\}$ $\{1,4,16,64\}$ $\{1,4,16,64,256,1024\}$ $\{1,4,16,64\}$
+Set5 PSNR 32.62 32.64 32.70 32.62 32.71 32.74 32.69 32.73
+Set5 SSIM 0.9018 0.9020 0.9029 0.9020 0.9027 0.9032 0.9024 0.9033
+Table 13: Ablation on the number of frequency components $N$ and scaling coefficients $\alpha$.
+We select $N=6$ for balanced quality and training efficiency, as $N=10$ yields marginal gains but converges slower.
+Set5 Performance / $\gamma$ Sawtooth Sin Triangular
+PSNR 32.71 32.69 32.71
+SSIM 0.9027 0.9026 0.9023
+Table 14: Comparison of alternative transformation bases $\gamma$.
+These results show that our design choices provide a good trade-off between representation flexibility and efficiency, while maintaining stable performance across a wide range of hyperparameter configurations.
+Appendix F Failure Cases
+Failure cases typically involve extremely fine textures—details that are absent in the low-resolution (LR) input but appear in the super-resolved (SR) output.
+However, this has been a shared inherent limitation of super-resolution methods to date.
+For example, in top row of fig.˜6, some white structures are prevent in HR but completely absent in LR.
+Appendix G Future Generalization
+For image generation, for example, we can integrate our Factorized Features into the VAE decoder of a diffusion model.
+For video generation, we can employ our mergeable basis and leverage its properties to enhance structural coherence.
 Appendix H More Visualization